A Complete Bibliography of The Journal of the American Statistical Association: 1920–1929

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.04

Title word cross-reference

1921 [Ano22j]. 1923 [Ano23i, Del23]. 1924 [Ano25d]. 1925 [Ano25f, Ano25e, Ano25g]. 1926 [Ano26e, Ano26d, Ano26f, Ano27g]. 1927 [Ano27i, Ano27h, Ano27j, Ano28f]. 1928 [Ano28h, Ano28g, Ano28i, Ano29c]. 1929 [Ano29e, Ano29d, Ano29f, Ano30c].

20th [Ano25q]. 25th [Ano25q].

Abridged [Gro23]. Abstracts [Ano24r, Cha28b]. Accidents [For28a]. Accounting [Han30]. Accounts [Cop28]. Accuracy [Dow29]. Acreage [Bey29, Nor28, Smi25b]. Activities [Ano22l, Hof24, Mit28]. Activity [Sny24b]. Adapted [Wri26b]. Added [Ano26e, Ano26d, Ano26f, Ano27g, Ano27i, Ano27h, Ano27j, Ano28f, Ano28h, Ano28g, Ano28i, Ano29c, Ano29e, Ano29d, Ano29f, MP27]. Addition [Ano27a]. Address [Del23]. Adjustment [Hur23b]. Adjustments [JT27].

Adopted [Ano22a, Ano25q]. Adult [Bak23]. Advantages [Smi26]. Adventure [Ros23]. Advisory [Ano22a, Ano25p, RDG^{+}22]. Affected [Sha22]. Affecting [Bur27, Hol23c]. After [Nad30, War30]. Age [Lev28, TW26]. Ages [Har23c]. Aggregative [Ash28]. Agricultural [B.28b, Gen30]. Agriculture [Aus30, Bea30b, H.24d, Ste22c, Tru22]. Aid [Pol22, She25]. Aids [Ach29]. Alcohol [Wil26b]. Alexander [Van26]. All-Inclusive [Alt22]. Allocation [Cru26, Lor26b]. Almshouse [Cla25c]. Amentia [Har23c]. Aments [Har23c]. America [Ano23b]. American [Ano22a, Ano22j, Ano23b, Ano23i, Ano25b, Ano26l, Ano27t, Ano30c, Cha24, Kin25c, KC23, Ano24b, Ano24a, Ano24i, Ano24n, Ano26a, Ano26g, Ano26n, Ano28a, Ano28j, Ano28p, Fry27, Glo26b, K.24, Kin27b, Lot27, Van24, Wol30]. Among [B.28a]. Analyses [Ros25]. Analysis [B.28a, Cha30, Dav29, Dou23, FF29, Ing23, Kuz28c, Rid27, Sas30, Wor24]. Annals [Tho27d]. Anniversary [Lur24]. Annual [Ano22j, Ano23i, Ano26l, Ano27t, Ano30c, Cha24, Kin25c, KC23, Pag29, Van24, Ano30b]. Answer [Gil27]. Answers [Ano25a, Fry27]. Anthropometric [Boa22b]. Any [Kin28, Eze24]. Apparatus [Cro25, Jen28]. Appendix [Ano22b, Ano22a]. Applicability [HT27, HT29]. Application [She25]. Applied [Mus23, Woo22]. Appointments [Ano27z]. Appraisal [AF30, TMT30]. Approach [Ror25]. April [Ano25q]. Arc [Car28a]. Area [McM29, PR29]. Areas [Dow29]. Arithmetic [Sib28]. Arizona [Ful28]. Arrests [Wil26b]. Aspects [Edi30, JS30, Les23b]. Assembly [Ano28e]. Assist [Cro25]. Associated [Har22a]. Association [Ano22j, Ano23i, Ano24b, Ano24n, Ano26a, Ano26n, Ano28a, Ano28p, Ano30c, Van24, Ano24a, Ano24i, Ano25b, Ano26g, Ano26l, Ano27t, Ano28j, Cha24, Glo24b, Kin25c, Kin27b, KC23]. Associations [Cam28, Ano22a, Ano23b]. Assumption [Fra29]. Assumptions [Eze25]. Asymmetrical [Var28]. Attempt [Ror25, Smi29]. Attributable [Wil26b]. August [Ano22d, Ano23d, Ano24e]. Australia [Ano23c]. Automatic [Hul25]. Average [Kin28, Nor28]. Averages [JT28, McC28].

B. [Jor25]. Back [Ano22c, Ano27b]. Bank [B.29, Bur23b, MC27]. Banking [Cri22]. Bar [CS27]. Bars [Eel26, vH27b]. Base [Bac28b]. Based [Pag29, Wat23c]. Basis [Hic26, Kin27b]. Be [Ach29, Ano27c]. Become [Hol23c]. Before [Blo29a, Smi25a]. Behavior [Bac28b]. Between [Boa22a, Cru23a, Fry28, Har22a, Har22b, Har24b, Hei26]. Bews [Cla24]. Bias [Lov28]. Billings [Wil26c]. Biology [Wri26a]. Birth [Dow29]. Births [Har23c, Sha22]. Board [Dou22, Ste22a, Car24, Sno22]. Bodies [Ano27u]. Bond [Mac26, Wil25]. Book [Ach28, Ada22, Alb22a, Alb22b, And26, And25, And28, Ang27, Ang28, App25, Arn29, Ayr23, Bai23, Bar22, Bar24, Bar28, Bau29, Bax23, Bea29a, BF23, Ben29, Ber23, Ber28, Ber24, Ber25, Ber26, Ber27, Ber29, Bey27, Bez24, Bla28a, Bla28b, Bla29, Blo27, Blo29b, Bre23, Bri22, Bri29, Bro28, Bro29, Buc22, Bur23a, BB28, Bur28, Bur29a, C.23a, C.23b, Car29, Car22, Cat28, Cha28a, Cla23a, Cla23b, Cla24, Cla25b, Cla25a, Cla26, Cla28, Cla29, Coa24, Com25, Cop22, Cop29a, Cro28, Dav24a, Dav28b, Dav28a, Dav23, Dav27b, Daw24, Daw28, Dic29, Dou24a, Dow22, Dow25, Dow28, Dub24, Eme25, Eva25, Eze26a, Eze27a, Eze28, Eze29, Fal25a, Fel25, Fie23, Fis23a, Fis24b, Fis24a, Fis26b, Fis27a, Fit26, Fli24, Foe24a, Foe24b, G.25, Gad25]. Book [Gan22, Gar29, Gav27, Geh25, Glo24a, Glo26a, Glo27, Gol24, Gol27, Goo26, Goo29, Gra25, Gra28, Gri24, Hal23, Han23, Han25, Han26, Han22a, Han27a, Han28, Har27a, Har27b, Har29, Har23a, Har28, Har23b, Haw29, Hen22, Hen28, Her25a, Her27, Her28a, Her22a, Her24, Her25c, Her28b, Hew26, Hew28, Hil22a, His25, Hof22a, Hof22b, Hol22, Hol28, Hot26a, Hot26b, Hot27a, Hot28, Hur27, Hur29, Jer26, Jer29, Jon26, Kar25, Kel22, Kin22a, Kin22b, Kin25a, Kin25b, Kin27a, Kin29a, Kna23, Kni29, Kop23, Kop24a, Kop24b, Kop24c, Kop25, Kop27a, Kop27b, Kop28, Kop29, Kuc29, Kuz28a, Kuz29a, Les23a, Lev29, Lin24, Lin29, Lor24, Lot26, Lot29a, Lot29b, Lub22, Lub24, Lub27, Lyo28, May29, Mer29, Mil26a, Mil22, Mil25, Mil29, Min23, Min26, Mit22, Mit23, Mor23b]. Book [Mud22, Mud23, Mus28, Mye29, New28, Not25b, Ogb22, Ogb23, Ogb28, Orc26, Orc28, Orc29, Par26, Par25, Par29, Pat22, Pat23a, Pat23b, Pat24, Pat26, Pat27, Pat29, Pat25, Pat28, Per22, Phe22a, Pog22, Pog23, Pog26, Rau24, Ree26, Ric25a, Ric28, Rid25, Rie25a, Rie26, Rie28, Rie29a, Rie23, Rie29b, Rob26, Rob29b, Rob22, Rol28a, Rol28b, Rol29, Ros22, Ros24a, Ros27, Rub25, Rub28, Sas29, Sch28, Sec22a, Sec27, Sec28, Sha24, Shu26, Shu29, Sil29, Sim29, Sli23, Sou22, Sou23, Sou25, Squ23, Ste24a, Ste27a, Ste22b, Ste22d, Ste24b, Ste27b, Sto28, Syd24, Syd28a, Tau26, Tay29, Tho25a, Tho25b, Tho26a, Tho27a, Tho27c, Tho25e, Tho26c, Tho28, Tho29a, Try22a, Try22b, Try29, Tuc27, Ush29, Van28, Van22, Vin22, Vin24a, Vin24b, Vin27, Von27]. Book [War26, War27a, War27b, Wat22, Wat23a, Wat27a, Wey22, Wey26, Wey27, Whi23, Wil22a, Wil22b, Wil24a, Wil27b, Wil27c, Wil28c, Wil22c, Wil23b, Wil24d, Wil26d, Wil22d, Wil27d, Win25, Woo24a, Woo26, Wor28, Wor29, Wri24, Wri26a, Wri28, Wri29, Ynt28, You24, You25a, vH25a, vH25b, vH26, vH27a, vH28a, vH28b]. Books [Ano22k]. Born [Ful28]. Bring [Dav26b]. Britain [Sno22]. British [Cri22, Tho25c, Tho26b]. Broker [Sny30]. Brussels [Del23]. Budget [Ach29]. Building [Cam28, Hol30, Kin24b]. Bureau [Car24, Tob22, Ano22l, Ano26m, Ano27u, H.23a, H.23b, H.23c, H.24c, H.24b, H.24d, H.24a, H.25a, H.25b, H.25c, H.25d, H.26a, H.26b, H.26c, H.27a, H.27b, H.27c, H.28b, H.28c, H.28d, H.29b, H.29c, H.29d, H.29e, Ste27c]. Business [Ano23o, Ano28u, Ano30b, Ayr22, B.28a, Day28, Den22, Dun22, Fis23c, Fis25a, Han27b, Hun22, Ing23, Kar26, Lam22, Lan24, Mil26b, OT22, Per29, Ric25b, Rie25b, Sny24a, Sny24b, Tho26b, Tho27d, Wil24e, WH27]. Business-Cycle [Ano23o]. Buying [Han22b]. By-Laws [Ano24b, Ano26a, Ano28a].

Calculating [Hol23b, Hul25]. Calculation [Cra22, Cro25, Har27c, Jen28].

California [Bac28a]. Called [Fis25a]. Campaign [Dav26b]. Can [Ach29, Ano27c]. Canada [Ros26]. Canadian [Coa25, Dex23, Hur23a, Per29]. Capacity [MM25, You25b]. Capita [W.28]. Cars [Ros26]. Case [Mor23a, Pie26]. Cash [Bur29c]. Casualities [Roe25]. Categories [Ric24a]. Category [HT29]. Causes [Gil27, Roe26]. Census [Ano22b, Ano25p, Ano27u, Ano29p, Aus30, Bea30a, H.24c, H.24b, H.24d, H.28a, Hil22b, Hof22d, McF30, Per30, Ric24a, RDG^{+}22, Ste22c, Tru22, Tru30, Ano22l, Ano23b, Ano26m, Ano27u, H.23a, H.23b, H.23c, H.24c, H.24b, H.24d, H.24a, H.25a, H.25b, H.25c, H.25d, H.26a, H.26b, H.26c, H.27a, H.27b, H.27c, H.28b, H.28c, H.28d, H.29b, H.29c, H.29d, H.29e, Tob22]. Central [Kin29b]. Century [Lot27]. Certain [Edi30, JS30, NS24, OT22, Ror25, Wil27a]. Chaddock [Cha26a, Fis26a]. Change [Bey29]. Changes [Buc25, Bur27, Cop30, Gre25b, Man22a]. Charts [CS27]. Check [Bur29c]. Chicago [MJ28, PR29]. China [Ano27-27, Buc25, Wil28b]. Choice [Stu26]. Circle [CS27]. Circles [Eel26, vH27b]. Cities [Ano22g, Bac28a, H.24e]. City [Her25b]. Class [Kin28, Stu26]. Class-Interval [Kin28]. Classes [H.24b, Han22c]. Classification [Kin27b, Roe26]. Clearings [Sny24a]. Cleavage [WR24]. Clergymen [TW27]. Close [Kin27c]. Clothing [Wem23]. Coefficient [Bow28b, Cru23a, GM27, HT27, HT29, Hol23b]. Coefficients [Eel29, KM29]. Collecting [Ano22g]. Collection [Ano24a, Car24, Kin29b]. College [Hol23c, Tho25d]. Colleges [Glo26b]. Combination [Hol23b]. Combining [Smi26]. Come [Tho25d]. Comment [Bac28b, Lor26a, Tho27d]. Commerce [MC27, WKDR24]. Committee [Ano22a, Ano23b, Ano23n, Ano25p, Glo24b, RDG^{+}22, Van24, Ano25b]. Commodity [EH26]. Common [Bur22, Got25]. Commonwealth [Rub26]. Communication [Sec22b]. Communications [Ros24b]. Companies [Rob30]. Comparison [Car24, Man22a]. Competition [TW26]. Competitive [Ban27]. Component [Eel26]. Composite [Ano27a, Sny26]. Composition [Cha30]. Computation [Car24]. Computations [Ros25]. Computing [Cam28, Car27, KM29]. Concepts [Per24]. Concerning [Fra29, Mil26b]. Concomitance [Wat23c, Wat23b]. Conditions [OT22, TW26]. Conference [Ano25q, Car24, Dou22, Ste22a, Ano25q]. Congestion [Woo28]. Constant [Car27]. Constants [KS26, Lot27]. Constitution [Ano24b, Ano26a, Ano28a]. Construction [Fis27c, Mac26, McC28, Tho27b, Wat27b]. Consumer [Wai30]. Consumption [Try27]. Continental [Kin27c]. Contingency [HT27, HT29]. Contrasts [Har24b]. Copenhagen [Hei26]. Corporate [Eff30, Kna22]. Corporation [Dur30a, Han30, Hof24, Par30, Wat30]. Correction [vH28b]. Correlating [Smi25a, Smi26]. Correlation [Bea29b, Bow28b, Bra28, Cra22, Cru23a, Eel29, Eze24, Fal25b, GM27, Gum26, Har22a, Hol23b, Hul25, Kel25, KS26, Kuz28b, Mil24, Mus23, Nei27, Per23, Sas30, Smi29, TE23, TE27, Wat23c, Wor24]. Correspondence [Fis23b, Hil23]. Corresponding [For28b, Kin28]. Cost [Ach29, Car24, Jor25, McF24]. Cost-of-Living [Ach29]. Costs [Cru26, Lor26b]. Cotton [Ash29, J.28, KK26, Smi25b, Smi27, WKDR24]. Council [Ano27z, Ano29b, Mer26, Mit28]. Countries [Syd25]. Credit [B.29, Fal25b, Gol30, Sny30]. Cressy [Gui26]. Criminal [Fra28]. Criterion [Cru25a, Fis27c]. Criticism [Cha26a, Fis26a]. Critique [Har22c]. Crop [Bey29, J.28, Smi27]. Current [Lan24]. Curve [Dav25, Dav27a, MJ28, PR29, Ree25, Wri26b]. Curves [Dav26a, Pea24, Wat23c]. Curvilinear [Bea29b, Eze24, Eze26b]. Cycle [Ano23o, Den22, Fis23c, Fis25a, Hun22, Ing23, Lam22, OT22, WH27]. Cycles [Ber22b, Ber22a, Eic22, Kin24c, Mil26b, Tho26b]. Cyclical [Kuz29b]. Czecho [Hof22d].

D.C [Ano23i, Ano30c]. Dance [Fis23c]. Data [Ano24a, B.28a, Gre25b, Han30, Hog27, Joh29, NS24]. Date [Fis27b]. Death [Har24a, Lur24, Roe26]. Deaths [Roe26, Wil26b, Woo22]. Debits [Cop28]. Decade [Bea30b]. December [Ano22j, Ano23i, Ano30c, Ano24f, Ano25d, Ano27g, Ano28f, Ano29c]. Defining [H.27c]. Degrees [Hol23c]. Demand [Blo23, Bro25, Moo22, Pre22, Wai30, Wil29]. Department [JT27, Man22b, Mel23, WKDR24]. Departmental [Rub26]. Dependency [Ano27c, Cha23]. Dependent [Har22a]. Deposits [Bur23b]. Depreciation [Hot25, Tay23]. Depression [Hur23b]. Determination [Bey29, Car28a, Cru22, Eze26b, Gre25b, Hal26, Har22d, Hog27, Wor22]. Determined [Kel25]. Determining [Ach29, Cru23b, Har24b, KS26, Wau29]. Development [Dav26b, Dub26, Gui26, Van26, Wil26c]. Developments [Ano27-27, Gol30, Ste27c, Wai30]. Deviation [Bow28b, Cro25, Har22a, HN28, Hog27, Smi26]. Deviation-From-Trend [Smi26]. Device [Sib28]. Devices [Cam28]. Diagrams [CS27]. Difference [Bau28, Har22b, Hei26, Rob29a, Smi26]. Differences [Boa22a, Har22b]. Differential [Ach29, Har25, Hot27b]. Differentials [Whe29]. Differentiation [Mar27]. Dilemma [Ayr27]. Dinner [Ano23o]. Director [Ano23b]. Disappearing [Lun26]. Discussion [Dun22, Edi30, Her25b, Lam22, Lor26b, Sib26, Ste22c, TMT30, War30, Wil27a, vH27b]. Distribution [Cle27, Cro25, Dav25, Ell27, Hoy23, Jen28, Kin27c, McF30, Pol22, Pow28, Ric24b, Rid27, Tuc22, Zel29]. Distributions [Can26, Dav29, Har24b, Ror22, Var28]. Districts [H.27c]. Division [Tob22]. Divorce [Wil24c]. Dollar [Fis23c, Fis25a]. Doolittle [KM29, TE27]. Dreyer [Dub22]. Duration [Mil26b]. During [Han22b, Hei26]. Duties [Kin27b].

Earners [Phe22b]. Earning [TW26]. Earnings [Edi30, Eff30, JS30, Les23b, Win22]. Easter [JT27]. Economic [AF30, Ano22a, Ano23a, Ano23b, Bow28a, Bur29d, Cox30, Gre25b, Kin29c, Lev28, Sas30, TMT30, Woo24b]. Economics [Dot22, Kar24]. Educational [Glo24b]. Eell [vH27b]. Effect [Buc25, TW27]. Effects [Roe25]. Eighteenth [Lot27]. Eighth [Ano27t]. Eighty [Ano22j, Ano26l, Ano27t, Cha24, Kin25c, KC23, Ano23i]. Eighty-Eighth [Ano27t]. Eighty-Fifth [Cha24, Ano23i]. Eighty-Fourth [KC23]. Eighty-Seventh [Ano26l]. Eighty-Sixth [Kin25c]. Eighty-Third [Ano22j]. Elasticity [Moo22]. Elected [Ano22e, Ano22f, Ano22d, Ano22i, Ano23g, Ano23e, Ano23f, Ano23d, Ano24f, Ano24g, Ano24h, Ano24e, Ano25d, Ano25f, Ano25e, Ano25g]. Election [WR24]. Elementary [Mor28]. Elements [Ric25b, Rie25b]. Eliminating [Smi25a]. Empirical [Lot27]. Employees [Wem23]. Employment [Alt22, Ano23n, Ber22b, Ber22a, Ber22c, FF29, H.24e, Hur22, Hur23a, Les23b, Roe23, Wat27b]. England [Dex23]. Entry [Fry28]. Enumeration [H.27c]. Equation [Cop29b, Eze25]. Equations [Hot27b, Pag29, TE27]. Errata [Ano29a]. Erratic [TB29]. Error [Ano25k, Har22b, HN28, Hot27b, Rid24, Smi25a, Wil23c]. Errors [Eel29]. Establishments [Phe22b]. Estimate [Ano25k, Cop28, Kin30]. Estimates [Hot27b]. Estimation [Mil24]. Europe [Nad30, War30]. Evaluation [Cox30]. Events [Kuz29b]. Every [Dav26b]. Examinations [Dub22]. Exchange [Cop29b]. Expected [Woo22]. Experience [Wil24e]. Experiment [Wem23, vH27b]. Experimental [SW28]. Explain [Smi29]. Expression [Var28]. Extension [Sch25b]. Extracts [Del23].

Facilitate [Jen28]. Facilitating [Ros25]. Factor [Ban27, Kin24a]. Factoring [Bra28]. Factors [Ano27d, Bur27, Hol23c, Pow28, Woo24b]. Facts [Blo29a, Tho23]. Familial [Har25]. Families [Lot27, Sib26, Tho25d]. Family [Dou24b, Her25b, Wil30]. Famine [Roe25]. Famines [Buc25]. Farm [Ash28, BS24]. Farwell [Cla24]. February [Ano22e, Ano23e, Ano24g]. Fecundity [Har25]. Federal [Dur30a, Sno22, Van26, Wil26c]. Fellowship [Ano27z]. Fellowships [Ano29b]. Female [Sha22]. Fifteenth [Ano29p, Del23, H.28a]. Fifth [Cha24, Ano23i]. Fiftieth [Lur24]. Fifty [Sny24a]. Figures [Car24]. Final [Ano23b]. Financial [Ano22g]. Financing [Kin29d]. First [Ano28e, Bau28, Smi26, Wil23c, Ano30c]. First-Difference [Smi26]. Fisher [Hot27a]. Five [Dou24b]. Fixed [Bac28b]. Fixed-Base [Bac28b]. Flexibility [Moo22]. Fluctuation [Whi22]. Fluctuations [TW27, Wat27b]. Food [Cha30, McF24, Wil29]. Foodstuffs [Arn25]. Forces [KK26]. Forecasters [Ash29]. Forecasting [Ano25c, Ano27e, Blo23, Bos26, Bur29d, CW27, Day28, Eze27b, Hic26, Pag29, Pre22, Smi25b, Smi27]. Forecasts [AF30, Cla30, Cox30, TMT30]. Foreign [Arn25, Dur30b, Man22a, Syd25]. Form [Hol23b, Ree25]. Formation [Can26]. Formula [Dav24b, Lor26a, Mor23a, Mor28, NS24, Sch25a]. Formulae [Ros25]. Formulas [Eel29, Pag29]. Four [BS24, Win22]. Fourteenth [H.24c]. Fourth [KC23]. Free [TW26]. Freight [Bac28a, Lor26a, Sch25a]. French [Dex23]. Frequency [Can26, Cro25, Dav29, Kin28, Ric24b, Rid27, Ror22, Var28, Wri26b]. Freshmen [Pal29]. Friday [Ano30c]. Front [Ano24c, Ano26b, Ano27f, Ano28b, Ano28c, Ano28d, Ano30a]. Fuels [Try27].

Function [Cha26b]. Functions [Ror22]. Fundamental [Gre25b, Per24]. Fundamentals [Ror25]. Further [Her25b, vH27b]. Future [Dav26b, Dub26, Gui26, Osb23, Van26, Wil26c]. Futures [Ash29].

General [Ano28e, Hot25, Sny24c]. Generalized [Rid24, Sas30]. Geneva [Ano25q]. Geographical [Les23b]. Geometric [Bur29b, Sib28]. German [Car28b]. Germany [MM25, You25b, Roe25]. Government [Coa25]. Governmental [Ano25b, Van24]. Graded [Hog27]. Graduate [Hol23c]. Graphic [Bea29b, Sib28, vH27b]. Graphical [McC28]. Great [Sno22]. Group [Whi22]. Grouped [Cro25]. Groups [Kel25, Whe26b]. Growth [Blo23, Dav27a, Kin24b, Kni26, Kni27, Lot27, Pea24, Pre22, Ros23, Wri26a]. Guarantees [Dav28c].

H [Ing23]. Hampshire [R.23]. Handling [Eze24, TE23]. Harvard [Han27b, Kar26]. Hasty [Kin30]. Having [Wat23c]. Health [Ano24d, Gro23, K.28a, Syd28b]. held [Ano30c]. Higher [Hol23c, McF24]. Hog [Eze27b]. Hollins [Pal29]. Homer [Cla24]. Homogeneity [Kel25]. Horsepower [Tho29b]. Hotel [Ano30c]. Hypothesis [Mil26b].

II [Ano22b, Kni24b, Van26]. III [Gui26]. Illegal [Fry28]. Immigration [Blo29a]. Implications [Wil27a]. Implied [Eze25]. Importance [Bur29c]. Improved [Kin24a]. Improvements [Car28b]. Incidence [Bri23]. Inclusive [Alt22]. Income [Hic26, Hoy23, Kin30, Kna22, Pea25, Whi22]. Increase [DL25, Whe29, Wri26a]. Increasing [Ano26c]. Index [Alt22, Ano27a, Arn25, Ash28, BS24, Bur22, Cha25, Dav24b, Ell27, EH26, Fis23d, Fis24c, Fis25b, Fis26c, Fis27b, Fis27c, Fis28, Joh29, Kni24a, Kni24b, Lov28, Mac26, Mar27, MC27, Per29, Pet22, Sno22, Sny23, Sny24a, Sny24b, Sny24c, Sny25, Sny26, Sny28, Tho26b, Tho27b, Try27]. Indexes [B.28b, Car27, Cop29b, Sou24, Tho25c, TB29, Wat23c, Han27b, Kar26]. Indication [Bey29]. Indices [Jor25]. Individual [Cop28]. Industrial [Car24, Dou22, Han22c, Hur23b, Ste22a]. Industry [Bab28, Kin29d, Wat27b]. Inequality [Hoy23]. Infant [Ano23c, Nea25, Woo22, Woo24b, WH27]. Inference [Wil27e]. Influence [JT27, OT22, Pow28]. Influencing [Ano27d]. Influenza [Eic22]. Information [Ano22q, Ano24t, Ano25t, Ano26t, Ano27-29, Ano28v, Ano29q]. Inquiry [Ano29p]. Institute [Del23, Wil26a]. Institutional [H.24b]. Insurance [Ano28u, Dub22]. Integration [Hal25]. Interest [Bur27, K.29]. Interests [Kin27b]. Interim [Ano28e]. Interlocking [RG29]. Intermarriage [Dex23]. International [Ano25q, Ano28e, Del23, Pea28, Wil26a, Wil24b]. Interpolation [Mor23a, NS24, Sib28]. Interpretation [Edi30, Han27b, JS30, Kar26, Tho23, Wat30]. Interval [Kin28, Stu26]. Investigation [Ano28e, Boa22b]. Investment [Rob30, Ror30]. Investor [Han30]. Investors [Par30]. Invited [Ano25a]. Iron [Mye22, Wor24]. Isolation [Kin24c]. Isorropic [Wau29]. Italy [Gal27]. Items [For28b].

Iteration [KS26, KM29, TE27]. Itself [CW27]. IV [Dav26b].

J [Wri26a]. John [Wil26c]. Joint [Ano25p, Cru26, Lor26b, Ano22a, Ano23b]. June [Ano25e, Ano26d, Ano27h, Ano28g, Ano29d]. Juvenile [FF29].

Karsten [Han27b]. Kelley [KM29, TE27]. Kenneth [Cla24]. Kinetic [Pol22]. King [Van26]. Known [For28b].

L [Gui26, Ing23, Jor25]. Labor [Ano25b, Ano26c, Bri23, Bur22, Car24, Han22b, Ste27c, Van24, Wol30]. Labour [Ano25q]. Lagrange [Mor23a]. Largely [Fis23c]. Last [Hei26, Hur23b]. Law [Blo23, Bro25, Lot27, Pre22, Rid24, Wil27e]. Laws [Ano24b, Ano26a, Ano28a, Blo29a, Kni26, Kni27, Wil23c]. League [Ano24d]. Least [Cru25a]. Legislation [H.28a]. Level [Cop30, Sny24c]. Levels [Gil27]. Life [Dub22, Gro23]. Limitation [HT27]. Limitations [HT29]. Limiting [Mor23a]. Line [CW27]. Lines [Cru25a, Wau29]. Link [Bac28b]. Literature [Ano22m, Ano22n, Ano22o, Ano22p, Ano23j, Ano23k, Ano23l, Ano23m, Ano24o, Ano24p, Ano24q, Ano25l, Ano25m, Ano25n, Ano25o, Ano26o, Ano26p, Ano26q, Ano26r, Ano27v, Ano27w, Ano27x, Ano27y, Ano28q, Ano28r, Ano28s, Ano28t, Ano29m, Ano29n, Ano29o, Mer24]. Living [Ach29, Car24, Jor25]. Loan [Cam28]. Loans [Sny30]. Logarithmic [Dav25, Dav26a, Wat23c]. Logical [Ror25]. Logistic [MJ28, PR29]. London [Bow28a]. Long [Dav28c, Eic22, Fra23]. Long-Time [Dav28c, Eic22]. Losses [Wil28a]. Louis [FF29].

M [Del23]. Machine [Bra28, Hul25, Wem23]. Maintaining [She25]. Major [Wil25]. Making [KK26]. Male [Sha22]. Management [Den22, Lam22, Cla24]. Mankind [Wri26a]. Manufactured [She25]. Manufactures [Bea30a, Tho29b]. Manufacturing [MP27, Phe22b]. March [Ano25f, Ano26e, Ano27i, Ano28h, Ano29e]. Market [B.29, Fis30]. Markets [KK26, Rob30]. Massachusetts [How22, McF24, Phe22b]. Massachussetts [MP27]. Mathematical [Hot25, Ing23]. Mathematics [Ror25]. Matter [Ano22c, Ano24c, Ano26b, Ano27b, Ano27f, Ano28b, Ano28c, Ano28d, Ano30a]. May [Ano22f, Ano23f, Ano24h]. Mean [Bur29b, Har27c, Hog27]. Means [Har22d]. Measure [Wil27a]. Measured [Ano27c]. Measurement [Ano23n, Boa22a, Cle27, Fal24, Gre25a, JT28, Mil24, Wat23c, Wat23b, Wil29, Wil30]. Measurements [Pal29, Pow28]. Measures [Kel25, Ror22]. Measuring [Bau28, K.28a, Kin24a, Ror30, Wat27b]. Median [Cru23b]. Medians [Win22]. Meeting [Ano22j, Ano23i, Ano24a, Ano26l, Ano27t, Ano30c, Cha24, Del23, Kin25c, KC23, Ano30b]. Members [Ano22e, Ano22f, Ano22d, Ano22i, Ano23g, Ano23e, Ano23f, Ano23d, Ano24a, Ano24i, Ano24f, Ano24g, Ano24h, Ano24e, Ano25d, Ano25f, Ano25e, Ano25g, Ano26e, Ano26d, Ano26g, Ano26f, Ano27g, Ano27i, Ano27h, Ano27j, Ano28f, Ano28h, Ano28g, Ano28j, Ano28i, Ano29c, Ano29e, Ano29d, Ano29f, Kin27b].

Memberships [RG29]. Men [Hei26, Tuc22]. Merits [Eel26]. Method [Ano22g, Bau28, Bea29b, Eze24, Har22d, Har24b, KS26, KM29, Kin24a, Rob29a, Sch25b, TE23, TE27, Woo22]. Methods [Car27, Car24, Cha26a, Eze27b, Fis26a, Hot27a, K.28a, Smi26, Wem23]. Mexican [Blo29a, Ful28]. Mexican-Born [Ful28]. Mexico [Ful28]. Mid [Kin28]. Mid-Point [Kin28]. Migration [Bab28]. Military [Tuc22, Wil28a]. Minutes [Ano30b]. Miscellaneous [Ano22h, Ano23h, Ano24j, Ano24k, Ano24l, Ano25h, Ano25i, Ano25j, Ano26h, Ano26i, Ano26j, Ano27k, Ano27l, Ano27m, Ano27n, Ano28k, Ano28l, Ano28m, Ano28n, Ano29g, Ano29h, Ano29i, Ano29j, Cha22, EB26, HMR23, Hof22c, Not25a, PP23, Ste24c, Wes22, WV23]. Mobility [Bri23]. Moments [Har27c, Jen28, Sch25b]. Monday [Ano30c]. Mongolian [Har23c]. Monthly [For28b, Har22d, Mye22, Wor24]. Months [Bau28]. Moore [Ing23]. Mortality [Ano23c, Bak23, Hei26, Nea25, Roe25, Woo22, Woo24b, WH27]. Mothers [Har23c]. Movements [Ano27d]. Moving [Hal25, JT28, Kuz28b, McC28]. Multiple [Bra28, Eze25, KS26, KM29, Sas30, Smi29, TE23, TE27].

N.I.C.B. [Jor25]. Natality [Ano24m]. Nation [Kin29d]. National [Dou22, Hic26, Kin30, Kna22, Ste22a, Car24, MC27]. Nations [Ano24d]. Natural [DL25, Whe29, Wri26a]. Nature [Ayr22, Dun22, Kni24a, Kni24b]. Nearly [TW26]. Need [Joh29]. Needs [Par30]. Negro [Her25b]. Net [Arn25, Kin22c, Kin22d, Pea25]. Ninety [Ano30c]. Ninety-First [Ano30c]. Non [Hol23a, Pie26]. Non-Representative [Pie26]. Non-Smokers [Hol23a]. Normal [Fra29, Whe26a]. Note [Cla25c, Ric24a]. Notes [Ano22h, Ano23h, Ano24j, Ano24k, Ano24l, Ano25h, Ano25i, Ano25j, Ano26h, Ano26i, Ano26j, Ano27k, Ano27l, Ano27m, Ano27n, Ano27o, Ano27p, Ano27q, Ano27r, Ano27s, Ano28k, Ano28l, Ano28m, Ano28n, Ano29g, Ano29h, Ano29i, Ano29j, Ano29k, Ano29l, Cha22, EB26, HMR23, H.25a, H.29a, Hof22c, MH29, Not25a, PP23, Ste24c, Wes22, WV23]. November [Ano23g]. Number [Arn25, Dav24b, Eze24, Fis23d, Fis24c, Fis25b, Fis26c, Fis27b, Fis27c, Lov28, Mac26, Phe22b, Tho27b]. Numbers [BS24, Bur22, Mar27, Pet22].

Obituary [Ano27o, Ano27p]. Objects [Roe26]. Observations [Fra29]. Occupational [Bri23, Whe26b]. Occupations [Ful28]. Occurrence [Wri26b]. October [Del23]. Officers [Gro23]. Ohio [Wat27b]. Oil [Osb23]. Operating [Pea25]. Operators [Wem23]. Opinion [Ric24b]. Orders [Ano27e]. Ordinate [Cru23a]. Organization [Coa25, Gal27]. Organizations [B.28a]. Orientals [Fry28]. Origin [Har27c]. Original [Ano24a]. Oscillations [Kuz29b]. Other [Eze26b]. Our [Cop30, Fis25a]. Outlook [Kin29c]. Overlapping [TB29]. Ownership [Pow28].

PA [Ano22j]. Pandemic [Eic22]. Panic [Fis30]. Paradox [Sec23]. Part [Kni24b]. Partial [Pag29]. Particular [Pie26]. Parts [Eel26]. Past [Ano28o, Dav26b, Dub26, Gui26, Hal26, Van26, Wil26c]. Pay [MM25, You25b]. Payments [Bur29c]. Pearl [Lot27, Wri26a]. Pearl-Reed [Lot27]. Pennsylvania [Rub26]. Percentage [Cro22, Har24b, Har26, Wri26b]. Percentages [Bur29b]. Periodogram [Cru23a]. Personal [Ano27q, Ano27r, Ano27s, Ano29k, Ano29l, H.29a, MH29, Whi22]. Physical [Pal29, Per29, Ror22, Ror25]. Physically [Tuc22]. Physiology [Gum26]. Pig [Mye22, Wor24]. Pittsburgh [Ano22j]. Place [Kna22, Syd28b]. Plan [Coa25]. Plant [Wem23]. Plotting [Dav26a]. Point [Kin28]. Political [Ric24b]. Polygon [Kin28]. Poor [How22]. Population [Ano28e, Boa22b, Cla25c, Dub25, Ful28, Her25b, Hil22b, Hot27b, Kin24b, Kni26, Kni27, Lev28, Lot27, MJ28, Pea28, PR29, Ros23, Syd25, Tho23, Tru30, Wil23a, Wil28b, Wri26a]. Post [Bea30b]. Post-War [Bea30b]. Power [Ash28, Han22b, TW26, Try27]. Pre [Har22c]. Pre-War [Har22c]. Precise [Bey29]. Prediction [MJ28]. Preliminary [Ano23i, Wil23a]. Presence [Eze26b]. Present [Dub26, Hal26]. President [Mer26]. Presidential [WR24, Del23]. Presidents [Ano28o]. Price [Ano26k, Bac28b, Bos26, Buc25, Cop30, Edi30, Fra23, J.28, JS30, K.28b, KK26, Kni24a, Kni24b, MC27, Pet22, Sno22, Sny24c, Sou24, TW27]. Price-Earnings [Edi30, JS30]. Price-Fluctuations [TW27]. Price-Index [Kni24a, Kni24b]. Prices [Ano25c, Ano27d, Ash29, Bac28b, BS24, Buc25, Eff30, EH26, Eze27b, Fis23d, Han24, Hug30, K.29, Mar27, Moo22, Sni30, Wil25]. Primarily [Wat23c]. Principle [Fis27c, Roe26]. Principles [Cha26a, Fis26a, Kin24c]. Probability [Bro25]. Probable [Eel29, Har22a, Har22b, HN28, Wil27e]. Problem [Dav24b, Dub25, Got25, Mil24, Per30, Sny29, Wor24]. Problems [Ano28e, Cam23, TE23, Wat30]. Proceedings [Ano28e, Ano22j, Ano26l, Ano27t, Cha24, Kin25c, KC23]. Processes [MP27]. Producing [Pow28]. Product [She25]. Production [B.28b, Mye22, Nor28, Pol22, Sni30, Tho25c, Tho27b, Wor24]. Productivity [W.28]. Products [B.28b]. Professional [Glo24b, TW26]. Professor [Cha26a, Fis26a, Got25, Ing23, MM25, You25b]. Professors [Hol23c]. Profits [Cha25]. Program [Ano23i]. Progress [Ano22b, Ano26m, Ano27u, Dav26b, H.23a, H.23b, H.23c, H.24c, H.24b, H.24d, H.24a, H.25b, H.25c, H.25d, H.26a, H.26b, H.26c, H.27a, H.27b, H.27c, H.28b, H.28c, H.28d, H.29b, H.29c, H.29d, H.29e, K.28a, Lan24]. Progressive [Car27, Cru25b]. Proof [Mor28]. Proposal [Ano22k]. Proposed [Wat23c]. Prosperity [Dav28c, K.28b, Sny29]. Protractor [Cro22]. Public [For28a, K.28a, Lun26, Syd28b]. Publications [Ano24n, Ano26n, Ano28p, H.24c]. Puerperal [Har24a]. Purchasing [Ash28]. Purpose [Cop29b, Woo28]. Pyramiding [Sny30].

Quadrature [Kar24]. Quality [She25]. Quantitative [Gre25b]. Quantities [Boa22a]. Quantity [Kni24a, Kni24b]. Quantity-Index [Kni24a, Kni24b]. Quarterly [Ano26n, TB29]. Questions [Ano25a]. Quetelet [Lur24]. Quinquennial [H.24d]. Quota [Blo29a]. Quotas [Ano26s].

R. [Hot27a]. Railroads [Cla24]. Railway [Pea24, Pea25]. Random [Kuz29b]. Rate [DL25]. Rates [Bur27, Cam28, Cla24, Har24a, K.29, Lor26a, Sch25a, W.28, Wil24c]. Ratio [Bau28, Cra22, MP27, Sha22]. Ratios [Edi30, Hol23b, JS30]. Raymond [Wri26a]. Reconsidered [Wor22]. Record [Hur23b]. Recorded [Dow29]. Reed [Lot27]. Reference [Ano22k, Wat23c]. Refinement [Ing25]. Region [Ano23a, MJ28]. Registration [Dav26b, Dub26, McM29, Roe26, Van26]. Regression [Eze25, Eze26b, KM29, Pag29, Wau29]. Rejoinder [Ber22c, Dou22, Ric25b, You25b]. Relation [Dub22, TW26, Wil24b]. Relationship [Kin28]. Relative [Bur29c, Eel26, Hal24]. Relatives [Bac28b]. Reliability [Har26]. Relief [Ano27c, How22]. Religious [Ano27u]. Rent [Gil27]. Reply [Cha26a, Her22b, Ste22a]. Report [Ano23b, Ano23o, Ano23n, Ano25p, Ano28e, Boa22b, Dav26b, Fis27b, Fis28, Mer26, RDG^{+}22, Van24]. Reporting [For28a]. Representative [Pie26]. Representing [Eel26]. Requirements [Glo26b]. Research [Ano27z, Ano29b, Ayr22, Dun22, Hol23c, Hot27a, Mel23, Mer26, Mit28]. Resemblance [Cru23a]. Reserve [Sno22]. Resolutions [Ano22a, Ano25q]. Respect [Par30]. Restriction [Blo29a]. Results [Hil22b, Hof22d, SW28, Ste22c, Tru22]. Retraction [Whe27]. Returns [Whi22]. Review [Ach28, Ada22, Alb22a, Alb22b, And26, And25, And28, Ang27, Ang28, App25, Arn29, Ayr23, Bai23, Bar22, Bar24, Bar28, Bau29, Bax23, Bea29a, BF23, Ben29, Ber23, Ber28, Ber24, Ber25, Ber26, Ber27, Ber29, Bey27, Bez24, Bla28a, Bla28b, Bla29, Blo27, Blo29b, Bre23, Bri22, Bri29, Bro28, Bro29, Buc22, Bur23a, BB28, Bur28, Bur29a, C.23a, C.23b, Car29, Car22, Cat28, Cha28a, Cla23a, Cla23b, Cla24, Cla25b, Cla25a, Cla26, Cla28, Cla29, Coa24, Com25, Cop22, Cop29a, Cro28, Dav24a, Dav28b, Dav28a, Dav23, Dav27b, Daw24, Daw28, Dic29, Dou24a, Dow22, Dow25, Dow28, Dub24, Eme25, Eva25, Eze26a, Eze27a, Eze28, Eze29, Fal25a, Fel25, Fie23, Fis23a, Fis24b, Fis24a, Fis26b, Fis27a, Fit26, Fli24, Foe24a, Foe24b, Fry27, G.25]. Review [Gad25, Gan22, Gar29, Gav27, Geh25, Gil27, Glo24a, Glo26a, Glo27, Gol24, Gol27, Goo26, Goo29, Got25, Gra25, Gra28, Gri24, Hal23, Han23, Han25, Han26, Han22a, Han27a, Han28, Har27a, Har27b, Har29, Har23a, Har28, Har23b, Haw29, Hen22, Hen28, Her25a, Her27, Her28a, Her22a, Her24, Her25c, Her28b, Hew26, Hew28, Hil22a, His25, Hof22a, Hof22b, Hol22, Hol28, Hot26a, Hot26b, Hot27a, Hot28, Hur27, Hur29, Jer26, Jer29, Jon26, Kar25, Kel22, Kin22a, Kin22b, Kin25a, Kin25b, Kin27a, Kin29a, Kna23, Kni29, Kop23, Kop24a, Kop24b, Kop24c, Kop25, Kop27a, Kop27b, Kop28, Kop29, Kuc29, Kuz28a, Kuz29a, Les23a, Lev29, Lin24, Lin29, Lor24, Lot26, Lot29a, Lot29b, Lub22, Lub24, Lub27, Lyo28, May29, Mer29, Mil26a, Mil22, Mil25, Mil29, Min23, Min26]. Review [Mit22, Mit23, Mor23b, MM25, Mud22, Mud23, Mus28, Mye29, New28, Not25b, Ogb22, Ogb23, Ogb28, Orc26, Orc28, Orc29, Par26, Par25, Par29, Pat22, Pat23a, Pat23b, Pat24, Pat26, Pat27, Pat29, Pat25, Pat28, Per22, Phe22a, Pog22, Pog23, Pog26, Rau24, Ree26, Ric25a, Ric25b, Ric28, Rid25, Rie25a, Rie26, Rie28, Rie29a, Rie25b, Rie23, Rie29b, Rob26, Rob29b, Rob22, Rol28a, Rol28b, Rol29, Ros22, Ros24a, Ros27, Rub25, Rub28, Sas29, Sch28, Sec22a, Sec27, Sec28, Sha24, Shu26, Shu29, Sil29, Sim29, Sli23, Sou22, Sou23, Sou25, Squ23, Ste24a, Ste27a, Ste22b, Ste22d, Ste24b, Ste27b, Sto28, Syd24, Syd28a, Tau26, Tay29, Tho25a, Tho25b, Tho26a, Tho27a, Tho27c, Tho25e, Tho26c, Tho27d, Tho28, Tho29a, Try22a, Try22b, Try29, Tuc27, Ush29]. Review [Van28, Van22, Vin22, Vin24a, Vin24b, Vin27, Von27, War26, War27a, War27b, Wat22, Wat23a, Wat27a, Wey22, Wey26, Wey27, Whi23, Wil22a, Wil22b, Wil24a, Wil27b, Wil27c, Wil28c, Wil22c, Wil23b, Wil24d, Wil26d, Wil22d, Wil27d, Win25, Woo24a, Woo26, Wor28, Wor29, Wri24, Wri26a, Wri28, Wri29, Ynt28, You24, You25a, You25b, vH25a, vH25b, vH26, vH27a, vH28a]. Reviews [vH28b]. Revised [Sny25]. Revision [Fis24c, Fis25b, Fis26c, Pag29, Sny28]. Richter [Ric25b, Rie25b]. Rise [Buc25]. Role [Day28, Rob30]. Room [Woo28]. Rural [Dow29, Ric24a].

S. [Jor25]. Salaries [TW27]. Sales [Ano26s, Ano27e, JT27]. Salisbury [KM29, TE27]. Samples [SW28]. Sampling [Cam23, Pie26, Sib26]. Saturation [Ree25]. Saving [Kin22c, Kin22d]. Scale [Dav26a]. Scales [Wil30]. Schemata [Mus23]. Science [Ano24r, Ano27z, Ano29b, Cha28b, Mer26, Mit28, RG29]. Scientific [Ano28e]. Season [Gen30]. Seasonal [Bau28, Car27, Cha23, Cle27, Cru23b, Fal24, Gre25a, Hal24, Har22d, JT28, Kin24a, Rob29a, Smi25a, Whe26a]. Seasonality [Cru25b]. Second [Ano25q, HT29, Sny28, Wil23c]. Sectional [Man22b]. Secular [Cru22, For28b, Hal24, Hal25, Hal26, Smi25a, Wor22]. Securing [Fra28]. Security [Ano25c, Ano27d, B.29, Hug30, K.29, Mar27, Rob30]. Selected [Whi22]. Selection [Wem23]. September [Ano25g, Ano26f, Ano27j, Ano28i, Ano29f]. Septicemia [Har24a]. Sequences [GM27, Kuz28b]. Series [Ano22k, Ing25, Kuz28c, Per23, Ros25, Rub26, Sas30, Smi25a, Smi26, TB29, Wat23c, Whe26a]. Service [Ano24d, Cla24, Dot22, Hur22, Roe23, Tuc22, Van26]. Services [Gal27]. Seventh [Ano26l]. Sex [WR24]. Shaped [Rid27]. Shaw [Wil26c]. Shirley [Wri26a]. Short [Bur27]. Shortened [NS24]. Significance [Bac28b, Gum26, Lev28, Lot27]. Significant [Har24b]. Similarity [GM27]. Simple [Smi29]. Simplified [Bea29b, Har24b]. Since [Ano22e, Ano22f, Ano22d, Ano22i, Ano23g, Ano23e, Ano23f, Ano23d, Ano24f, Ano24g, Ano24h, Ano24e, Ano25d, Ano25f, Ano25e, Ano25g, Ano26e, Ano26d, Ano26f, Ano27g, Ano27i, Ano27h, Ano27j, Ano28f, Ano28h, Ano28g, Ano28i, Ano29c, Ano29e, Ano29d, Ano29f, Blo29a, Wil26b, Wol30].

Situation [Hol30]. Sixth [Fis28, Kin25c]. Size [Lot27, Sib26, Tho25d]. Skewness [Ban27]. Slovakia [Hof22d]. Small [SW28]. Smokers [Hol23a]. So-Called [Fis25a]. Social [Ano24r, Ano27z, Ano29b, Bow28a, Car28b, Cha28b, Joh29, Kin29b, McM29, Mer26, Mit28, OT22, RG29]. Societies [RG29]. Sociological [Lev28]. Solving [TE27]. Some [Fra29, Fra28, Hil22b, Hol23c, Per24, Wai30]. Special [Cop29b]. Spot [Ash29]. Spurious [Gum26, Mus23, Nei27]. Squares [Cru25a]. St. [FF29]. Stability [Ano26c, Bac28b, Wil27a]. Staff [Bel24]. Standard [Ano22k, Ano25k, Bow28b, Cro25, Dav24b, For28a, HN28]. Standardized [Fra28]. Standards [Glo24b, K.24]. Standpoint [Han30]. State [Dav26b]. States [Ano22g, Tob22, Win22, Ano27a, Bak23, Ber22b, Ber22a, Boa22b, Bur29c, Cop28, Cop29b, Dav26b, Dub26, Fry28, Gui26, Han22c, Hof24, Hoy23, Hur22, Kin22c, Kin22d, Kin27c, Mac26, MP27, Roe23, Ros26, Sny26, Tho23, Van26, Whe26b, Wil24b, Wil26c]. Statistical [Ano22a, Ano22j, Ano22k, Ano23b, Ano23i, Ano24b, Ano24a, Ano24i, Ano24n, Ano25b, Ano26a, Ano26g, Ano26l, Ano26n, Ano27t, Ano27-27, Ano28a, Ano28j, Ano28p, Ano30c, Cam28, Cha24, Cha30, Cru26, Del23, FF29, Gal27, Glo26b, Hol23a, How22, Kin25c, Kin27b, KC23, Lev28, Lor26b, Sas30, Sec23, Tay23, Van24, Wil26a, Wil27e, Hot27a]. Statistician [Bel24, Dub25, Syd28b]. Statisticians [Ano25q, Glo24b, Glo26b, K.24]. Statistics [Ano22g, Ano24s, Ano25b, Ano28u, Ayr27, Ber22c, Bur29d, Car24, Car28b, Cha26a, Cha26b, Coa25, Cri22, Dav26b, Day28, Dot22, Dou23, Dow29, Dub26, Fis26a, Fra28, Gui26, H.24e, Har22c, Hur22, Hur23a, Kin29b, Lan24, Les23b, McM29, Par30, Per24, Pol22, Ric25b, Rie25b, Roe23, Rub26, She25, Syd25, Tho29b, Tob22, Van26, Van24, WKDR24, Wat30, Wem23, Wil24b, Wil26c, Wil24e, Woo28, Car24, Dur30a, Ste27c, Wil26c]. Status [Ayr22, Dun22]. Steel [Hof24]. Stirling [Mor28]. Stock [Eff30, Fis30]. Store [Man22b, Mel23]. Stores [JT27]. Strains [Her25b]. Strike [Dou23]. Structure [Fal25b]. Students [Hol23c, Tho25d]. Studies [Ach29, Bac28b, Cha30, Dou22, Ste22a, vH27b]. Study [Ano24m, Cam28, Hol23a, Hol23c, How22, Lev28, Nea25, Nei27, Pea25, Woo22]. Studying [Wai30]. Subject [Hot27b]. Succession [Wil27e]. Suggestions [Fra28]. Summary [Ste22c, Tru22]. Supervising [Bel24]. Supply [Osb23]. Surfaces [Eze26b, Wau29]. Surplus [Kna22]. Survey [Ano23a, Bow28a, Wil23a, R.23]. Swedish [Pet22]. Sweeney [Wri26a]. System [Dub26, For28a].

Tables [Gro23]. Tangent [Car28a]. Tax [Got25]. Taylor [Mor23a]. Teaching [Glo26b]. Technique [Woo28]. Techniques [Wai30]. Ten [Nad30, War30, Wil24e]. Tendency [Var28]. Tennessee [Hol23a]. Term [Fra23]. Terms [Bur27, Smi29]. Texas [Ful28]. Their [Har24a, Pow28, Tho23, Gil27, Glo26b]. Theory [Bro25, Hot25, Kar24, Tay23]. Third [Ano22j, RDG^{+}22]. Thirteen [Bau28]. Thirteen-Months-Ratio-First-Difference [Bau28]. Three [Lan24]. Time [Dav28c, Eic22, Fra29, GM27, Ing25, Kuz28c, Kuz28b, Per23, Ros25, Smi25a, Smi26, Wat23c, Whe26a]. Together [Cam28]. Total [Cop28, Fis27c]. Trade [Ano24s, Arn25, Dur30b, Kin29d, Man22a, Man22b, Sny23, Sny25, Sny28]. Traffic [Bac28a, Pea24]. Training [Cha26b, Glo26b]. Trend [Ano26k, Bak23, Car28a, Cru22, Cru25a, For28b, Hal24, Hal25, Smi25a, Smi26, Whe26a, Wor22, Zel29]. Trend-Seasonal [Whe26a]. Trends [Fra23, Hal26, K.28b, Kin24c, Man22b, Wil25]. True [DL25, Whe29]. Twenty [Hei26]. Two [Eze27b, Har22b, Har24b]. Types [BS24, NS24]. Typical [Dou24b].

U [Rid27]. U-Shaped [Rid27]. U.S. [Car24, Wil24c]. Undergraduate [Cha26b]. Underlying [Kin24c]. Unemployment [Ano29p, Ber22b, Ber22a, Har22c, Hun22, Per30]. Unequal [Pow28]. Unequivocal [Kni24a, Kni24b]. Unfit [Tuc22]. Union [Ano28e, Pea28]. United [Ano27a, Bak23, Ber22b, Ber22a, Boa22b, Bur29c, Cop28, Cop29b, Dav26b, Dub26, Fry28, Gui26, Han22c, Hof24, Hoy23, Hur22, Kin22c, Kin22d, Kin27c, Mac26, MP27, Roe23, Ros26, Sny26, Tho23, Tob22, Van26, Whe26b, Wil24b, Wil26c]. Universities [Glo26b]. University [Hol23a]. Unstable [Fis25a]. untitled [Ach28, Ada22, Alb22a, Alb22b, And26, And25, And28, Ang27, Ang28, Ano25r, Ano25s, Ano27-28, App25, Arn29, Ayr23, Bai23, Bar22, Bar24, Bar28, Bau29, Bax23, Bea29a, BF23, Ben29, Ber23, Ber28, Ber24, Ber25, Ber26, Ber27, Ber29, Bey27, Bez24, Bla28a, Bla28b, Bla29, Blo27, Blo29b, Bre23, Bri22, Bri29, Bro28, Bro29, Buc22, Bur23a, BB28, Bur28, Bur29a, C.23a, C.23b, Car29, Car22, Cat28, Cha28a, Cla23a, Cla23b, Cla25b, Cla25a, Cla26, Cla28, Cla29, Coa24, Com25, Cop22, Cop29a, Cro28, Dav24a, Dav28b, Dav28a, Dav23, Dav27b, Daw24, Daw28, Dic29, Dou24a, Dow22, Dow25, Dow28, Dub24, Eme25, Eva25, Eze26a, Eze27a, Eze28, Eze29, Fal25a, Fel25, Fie23, Fis23a, Fis24b, Fis24a, Fis26b, Fis27a, Fit26, Fli24, Foe24a, Foe24b]. untitled [G.25, Gad25, Gan22, Gar29, Gav27, Geh25, Glo24a, Glo26a, Glo27, Gol24, Gol27, Goo26, Goo29, Gra25, Gra28, Gri24, Hal23, Han23, Han25, Han26, Han22a, Han27a, Han28, Har27a, Har27b, Har29, Har23a, Har28, Har23b, Haw29, Hen22, Hen28, Her25a, Her27, Her28a, Her22a, Her24, Her25c, Her28b, Hew26, Hew28, Hil22a, His25, Hof22a, Hof22b, Hol22, Hol28, Hot26a, Hot26b, Hot28, Hur27, Hur29, Jer26, Jer29, Jon26, Kar25, Kel22, Kin22a, Kin22b, Kin25a, Kin25b, Kin27a, Kin29a, Kna23, Kni29, Kop23, Kop24a, Kop24b, Kop24c, Kop25, Kop27a, Kop27b, Kop28, Kop29, Kuc29, Kuz28a, Kuz29a, Les23a, Lev29, Lin24, Lin29, Lor24, Lot26, Lot29a, Lot29b, Lub22, Lub24, Lub27, Lyo28, May29, Mer29, Mil26a, Mil22, Mil25, Mil29, Min23, Min26, Mit22, Mit23]. untitled [Mor23b, Mud22, Mud23, Mus28, Mye29, New28, Not25b, Ogb22, Ogb23, Ogb28, Orc26, Orc28, Orc29, Par26, Par25, Par29, Pat22, Pat23a, Pat23b, Pat24, Pat26, Pat27, Pat29, Pat25, Pat28, Per22, Phe22a, Pog22, Pog23, Pog26, Rau24, Ree26, Ric25a, Ric28, Rid25, Rie25a, Rie26, Rie28, Rie29a, Rie23, Rie29b, Rob26, Rob29b, Rob22, Rol28a, Rol28b, Rol29, Ros22, Ros24a, Ros27, Rub25, Rub28, Sas29, Sch28, Sec22a, Sec27, Sec28, Sha24, Shu26, Shu29, Sil29, Sim29, Sli23, Sou22, Sou23, Sou25, Squ23, Ste24a, Ste27a, Ste22b, Ste22d, Ste24b, Ste27b, Sto28, Syd24, Syd28a, Tau26, Tay29, Tho25a, Tho25b, Tho26a, Tho27a, Tho27c, Tho25e, Tho26c, Tho28, Tho29a, Try22a, Try22b, Try29, Tuc27, Ush29, Van28, Van22, Vin22, Vin24a, Vin24b, Vin27]. untitled [Von27, War26, War27a, War27b, Wat22, Wat23a, Wat27a, Wey22, Wey26, Wey27, Whi23, Wil22a, Wil22b, Wil24a, Wil27b, Wil27c, Wil28c, Wil22c, Wil23b, Wil24d, Wil26d, Wil22d, Wil27d, Win25, Woo24a, Woo26, Wor28, Wor29, Wri24, Wri28, Wri29, Ynt28, You24, You25a, vH25a, vH25b, vH26, vH27a, vH28a]. Upon [TW27, Wat23c]. Urban [Dow29, Ric24a]. Urn [Mus23]. Use [Bra28, Cru23b, JT28, TB29, Wau29, vH27b]. Used [Ach29, Car24, Wem23]. Utilization [Wat30].

V [Dub26]. Validity [GM27]. Value [Fis27c, Har22a, Man22a, MP27, Smi27]. Values [Ror30]. Vanderblue [Cla24]. Variability [Har24a, Hei26, Her25b, Wat23c]. Variable [Boa22a, Har22a]. Variables [Eze24, Eze26b]. Variate [Rob29a]. Variation [Bau28, Cha23, Cru23b, Cru25b, Fal24, Fal25b, Gre25a, Hal24, Har22d, Rob29a, Smi25a, Wat23c, Wat23b, Wri26b]. Variations [Cle27, JT28, Nor28]. Varying [Kel25]. Velocity [Bur23b]. Versus [CS27, KM29, TE27]. Villagers [Fry27]. Violent [TW27]. Vital [Dav26b, Dub26, Gui26, K.24, Tob22, Van26, Wil26c]. Volume [Ano22q, Ano24t, Ano25t, Ano26t, Ano27e, Ano27-29, Ano28v, Ano29q, Cop28, Kin22c, Kin22d, Kin24b, Man22a, Per29, Smi27, Sny23, Sny25, Sny28]. Vote [Ric24b].

Wage [Ach29, Dou22, Phe22b, Sou24, Ste22a, W.28]. Wage-Earners [Phe22b]. Wages [Ano27a, Bur22, MP27, Sny26]. War [Han22b, Roe25, Wil28a, Bea30b, Har22c, Nad30, War30, Wil23a]. Wars [Sha22]. Washington [Ano23i, Ano30c]. Water [Try27]. Wealth [Hoy23, Kin27c, Pow28]. Weekly [Ell27, EH26, Fis23d, Fis24c, Fis25b, Fis26c, Fis27b, Fis28]. Welfare [Hof24]. Westergaard [Woo22]. Wheat [Bos26]. Which [Tho25d]. White [Her25b]. Who [Hol23c]. Wholesale [Cop30, Fis23d, Fis28, Han24]. Wilbur [Gui26]. William [Van26]. Wisconsin [Got25, Alt22]. within [Fal25b]. Women [Hei26, Win22]. Work [Ano22b, Ano26m, Ano27u, Dub22, Glo24b, H.23a, H.23b, H.23c, H.24c, H.24b, H.24d, H.24a, H.25b, H.25c, H.25d, H.26a, H.26b, H.26c, H.27a, H.27b, H.27c, H.28b, H.28c, H.28d, H.29b, H.29c, H.29d, H.29e, Syd28b]. Workers [TW26, Hot27a]. Working [Fra29]. World [Wil28a, Wil23a].

Year [Lan24, Mer26, Wil24e]. Yearly [For28b]. Years [Hei26, Nad30, Sny24a, War30]. Yields [Mac26, Nor28]. York [Ano23a, Ano24a, Her25b, MC27]. Young [MM25, You25b].

Zealand [Ano24m, Nea25].

References

Achinstein:1928:BRB

[Ach28] Asher Achinstein. Book review: [untitled]. Journal of the American Statistical Association, 23(164):459–461, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277301.

Achinstein:1929:CBC

[Ach29] Asher Achinstein. Can budget and cost-of-living studies be used as aids in determining a differential wage? Journal of the American Statistical Association, 24(165):28–39, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277315.

Adams:1922:BRB

[Ada22] T. S. Adams. Book review: [untitled]. Journal of the American Statistical Association, 18(140):543–545, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276937.

Andrew:1930:AEF

[AF30] Seymour L. Andrew and Harold M. Flinn. Appraisal of economic forecasts. Journal of the American Statistical Association, 25(169): 36–41, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277194.

Albrecht:1922:BRBa

[Alb22a] Arthur E. Albrecht. Book review: [untitled]. Journal of the American Statistical Association, 18(137):136–138, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277486.

Albrecht:1922:BRBb

[Alb22b] Arthur E. Albrecht. Book review: [untitled]. Journal of the American Statistical Association, 18(138):297–299, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277551.

Altmeyer:1922:AIE

[Alt22] A. J. Altmeyer. An all-inclusive employment index for Wisconsin. Journal of the American Statistical Association, 18(138):266–268, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277536.

Andrews:1925:BRB

[And25] John B. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 20(152):591–592, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277185.

Anderson:1926:BRB

[And26] Hilding Anderson. Book review: [untitled]. Journal of the American Statistical Association, 21(154):234–237, June 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277153.

Andrews:1928:BRB

[And28] John B. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 23(162):230–231, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276913.

Angell:1927:BRB

[Ang27] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 22(158):252–255, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276791.

Angell:1928:BRB

[Ang28] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 23(163):342–343, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276753.

Anonymous:1922:ARA

[Ano22a] Anonymous. Appendix I: Resolutions adopted by the Joint Advisory Committee of the American Statistical and the Economic Associations. Journal of the American Statistical Association, 18(137):88–93, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277471.

Anonymous:1922:AIP

[Ano22b] Anonymous. Appendix II: Progress of census work in 1921. Journal of the American Statistical Association, 18(137):94–98, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277472.

Anonymous:1922:BM

[Ano22c] Anonymous. Back matter. Journal of the American Statistical Association, 18(137):??, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277496.

Anonymous:1922:MESc

[Ano22d] Anonymous. Members elected since August. Journal of the American Statistical Association, 18(140):533, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276933.

Anonymous:1922:MESa

[Ano22e] Anonymous. Members elected since February. Journal of the American Statistical Association, 18(138):275, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277539.

Anonymous:1922:MESb

[Ano22f] Anonymous. Members elected since May. Journal of the American Statistical Association, 18(139):402–403, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276963.

Anonymous:1922:MCF

[Ano22g] Anonymous. Method of collecting financial statistics of cities and states. Journal of the American Statistical Association, 18 (140):517–519, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276927.

Anonymous:1922:MN

[Ano22h] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 18(138):271–275, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277538.

Anonymous:1922:NME

[Ano22i] Anonymous. New members elected since 1921. Journal of the American Statistical Association, 18(137):126, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277482.

Anonymous:1922:PET

[Ano22j] Anonymous. Proceedings of the Eighty-Third Annual Meeting of the American Statistical Association: Pittsburgh, PA, December 27–29, 1921. Journal of the American Statistical Association, 18(137):112– 121, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277480.

Anonymous:1922:PSS

[Ano22k] Anonymous. A proposal for a standard series of statistical reference books. Journal of the American Statistical Association, 18 (140):522–525, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276929.

Anonymous:1922:RAC

[Ano22l] Anonymous. Recent activities of the Census Bureau. Journal of the American Statistical Association, 18(140):515–517, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276926.

Anonymous:1922:RLa

[Ano22m] Anonymous. Recent literature. Journal of the American Statistical Association, 18(137):150–155, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277495.

Anonymous:1922:RLb

[Ano22n] Anonymous. Recent literature. Journal of the American Statistical Association, 18(138):300–304, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277552.

Anonymous:1922:RLc

[Ano22o] Anonymous. Recent literature. Journal of the American Statistical Association, 18(139):421–424, September 1922. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276975.

Anonymous:1922:RLd

[Ano22p] Anonymous. Recent literature. Journal of the American Statistical Association, 18(140):558–560, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276944.

Anonymous:1922:VI

[Ano22q] Anonymous. Volume information. Journal of the American Statistical Association, 18(137):1–9, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277460.

Anonymous:1923:ESN

[Ano23a] Anonymous. Economic survey of the New York region. Journal of the American Statistical Association, 18(144):1042–1043, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965667.

Anonymous:1923:FRJ

[Ano23b] Anonymous. Final report of the Joint Committee of the America Statistical and the American Economic Associations to the Director of the Census, 1922. Journal of the American Statistical Association, 18(141):628–649, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277504.

Anonymous:1923:IMA

[Ano23c] Anonymous. Infant mortality in Australia. Journal of the American Statistical Association, 18(141):653–654, March 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277507.

Anonymous:1923:MESd

[Ano23d] Anonymous. Members elected since August. Journal of the American Statistical Association, 18(144):1049–1050, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965670.

Anonymous:1923:MESb

[Ano23e] Anonymous. Members elected since February. Journal of the American Statistical Association, 18(142):803, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276724.

Anonymous:1923:MESc

[Ano23f] Anonymous. Members elected since May. Journal of the American Statistical Association, 18(143):924–925, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965480.

Anonymous:1923:MESa

[Ano23g] Anonymous. Members elected since November. Journal of the American Statistical Association, 18(141):676–677, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277513.

Anonymous:1923:MN

[Ano23h] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 18(144):1045–1049, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965669.

Anonymous:1923:PPE

[Ano23i] Anonymous. Preliminary program for the Eighty-Fifth Annual Meeting of the American Statistical Association, Washington, D.C., December 27–29, 1923. Journal of the American Statistical Association, 18(142):797–802, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276723.

Anonymous:1923:RLa

[Ano23j] Anonymous. Recent literature. Journal of the American Statistical Association, 18(141):695–698, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277522.

Anonymous:1923:RLb

[Ano23k] Anonymous. Recent literature. Journal of the American Statistical Association, 18(142):831–834, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276734.

Anonymous:1923:RLc

[Ano23l] Anonymous. Recent literature. Journal of the American Statistical Association, 18(143):945–948, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965490.

Anonymous:1923:RLd

[Ano23m] Anonymous. Recent literature. Journal of the American Statistical Association, 18(144):1077–1080, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965681.

Anonymous:1923:RCM

[Ano23n] Anonymous. Report of the committee on the measurement of employment. Journal of the American Statistical Association, 18(141):654– 657, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277508.

Anonymous:1923:RBC

[Ano23o] Anonymous. Report on the business-cycle dinner. Journal of the American Statistical Association, 18(141):657–659, March 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277509.

Anonymous:1924:COS

[Ano24a] Anonymous. The collection of original statistical data: Meeting of the New York members of the American Statistical Association. Journal of the American Statistical Association, 19(146):232–234, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277237.

Anonymous:1924:CLA

[Ano24b] Anonymous. Constitution and by-laws of the American statistical association: Constitution. Journal of the American Statistical Association, 19(148):3–4, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277109.

Anonymous:1924:FM

[Ano24c] Anonymous. Front matter. Journal of the American Statistical Association, 19(148):1–2, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277108.

Anonymous:1924:LNH

[Ano24d] Anonymous. The league of nations health service. Journal of the American Statistical Association, 19(146):234–236, June 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277238.

Anonymous:1924:MESd

[Ano24e] Anonymous. Members elected since August. Journal of the American Statistical Association, 19(148):533, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281574.

Anonymous:1924:MESa

[Ano24f] Anonymous. Members elected since December. Journal of the American Statistical Association, 19(145):101–102, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277270.

Anonymous:1924:MESb

[Ano24g] Anonymous. Members elected since February. Journal of the American Statistical Association, 19(146):241, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277240.

Anonymous:1924:MESc

[Ano24h] Anonymous. Members elected since May. Journal of the American Statistical Association, 19(147):409, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277384.

Anonymous:1924:MAS

[Ano24i] Anonymous. [members of the American Statistical Association]. Journal of the American Statistical Association, 19(148):5–31, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277110.

Anonymous:1924:MNa

[Ano24j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 19(145):96–101, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277269.

Anonymous:1924:MNb

[Ano24k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 19(147):404–409, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277383.

Anonymous:1924:MNc

[Ano24l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 19(148):525–532, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281573.

Anonymous:1924:NZS

[Ano24m] Anonymous. A New Zealand study in natality. Journal of the American Statistical Association, 19(148):454–467, December 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281562.

Anonymous:1924:PAS

[Ano24n] Anonymous. Publications of the American statistical association. Journal of the American Statistical Association, 19(148):32–51, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277111.

Anonymous:1924:RLa

[Ano24o] Anonymous. Recent literature. Journal of the American Statistical Association, 19(145):131–136, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277282.

Anonymous:1924:RLb

[Ano24p] Anonymous. Recent literature. Journal of the American Statistical Association, 19(146):269–272, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277251.

Anonymous:1924:RLc

[Ano24q] Anonymous. Recent literature. Journal of the American Statistical Association, 19(148):557–562, December 1924. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281583.

Anonymous:1924:SSA

[Ano24r] Anonymous. Social science abstracts. Journal of the American Statistical Association, 19(147):390–392, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277379.

Anonymous:1924:TS

[Ano24s] Anonymous. Trade statistics. Journal of the American Statistical Association, 19(147):392–396, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277380.

Anonymous:1924:VI

[Ano24t] Anonymous. Volume information. Journal of the American Statistical Association, 19(145):1–6, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277252.

Anonymous:1925:AQI

[Ano25a] Anonymous. Answers to questions invited. Journal of the American Statistical Association, 20(151):425–426, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965531.

Anonymous:1925:CGL

[Ano25b] Anonymous. Committee on Governmental Labor Statistics of the American Statistical Association. Journal of the American Statistical Association, 20(149):93–98, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277438.

Anonymous:1925:FSP

[Ano25c] Anonymous. Forecasting security prices. Journal of the American Statistical Association, 20(150):244–249, June 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277120.

Anonymous:1925:MESa

[Ano25d] Anonymous. Members elected since December, 1924. Journal of the American Statistical Association, 20(149):127–128, March 1925.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277445.

Anonymous:1925:MESc

[Ano25e] Anonymous. Members elected since June, 1925. Journal of the American Statistical Association, 20(151):438–440, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965533.

Anonymous:1925:MESb

[Ano25f] Anonymous. Members elected since March, 1925. Journal of the American Statistical Association, 20(150):271–274, June 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277127.

Anonymous:1925:MESd

[Ano25g] Anonymous. Members elected since September, 1925. Journal of the American Statistical Association, 20(152):574–575, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277177.

Anonymous:1925:MNa

[Ano25h] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 20(150):264–270, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277126.

Anonymous:1925:MNb

[Ano25i] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 20(151):426–437, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965532.

Anonymous:1925:MNc

[Ano25j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 20(152):565–574, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277176.

Anonymous:1925:SEE

[Ano25k] Anonymous. On the standard error of estimate. Journal of the American Statistical Association, 20(150):253–255, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277122.

Anonymous:1925:RLa

[Ano25l] Anonymous. Recent literature. Journal of the American Statistical Association, 20(149):158–162, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277459.

Anonymous:1925:RLb

[Ano25m] Anonymous. Recent literature. Journal of the American Statistical Association, 20(150):300–304, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277139.

Anonymous:1925:RLc

[Ano25n] Anonymous. Recent literature. Journal of the American Statistical Association, 20(151):462–466, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965543.

Anonymous:1925:RLd

[Ano25o] Anonymous. Recent literature. Journal of the American Statistical Association, 20(152):593–598, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277186.

Anonymous:1925:RJA

[Ano25p] Anonymous. Report of the joint advisory committee on the census. Journal of the American Statistical Association, 20(149):99–104, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277439.

Anonymous:1925:SIC

[Ano25q] Anonymous. The Second International Conference of Labour Statisticians at Geneva; April 20th to 25th, 1925. Resolutions Adopted by the Conference. Journal of the American Statistical Association, 20(152):561–565, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277175.

Anonymous:1925:Ua

[Ano25r] Anonymous. [untitled]. Journal of the American Statistical Association, 20(150):280–283, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277129.

Anonymous:1925:Ub

[Ano25s] Anonymous. [untitled]. Journal of the American Statistical Association, 20(150):283–284, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277130.

Anonymous:1925:VI

[Ano25t] Anonymous. Volume information. Journal of the American Statistical Association, 20(149):1–6, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277428.

Anonymous:1926:CLA

[Ano26a] Anonymous. Constitution and by-laws of the American statistical association. Journal of the American Statistical Association, 21(156):3– 4, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276765.

Anonymous:1926:FM

[Ano26b] Anonymous. Front matter. Journal of the American Statistical Association, 21(156):1–2, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276764.

Anonymous:1926:ILS

[Ano26c] Anonymous. Increasing labor stability. Journal of the American Statistical Association, 21(156):485–489, December 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276987.

Anonymous:1926:MASb

[Ano26d] Anonymous. Members added since June, 1926. Journal of the American Statistical Association, 21(155):358–359, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277070.

Anonymous:1926:MASa

[Ano26e] Anonymous. Members added since March 1926. Journal of the American Statistical Association, 21(154):231–233, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277152.

Anonymous:1926:MASd

[Ano26f] Anonymous. Members added since September, 1926. Journal of the American Statistical Association, 21(156):491–492, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276989.

Anonymous:1926:MASc

[Ano26g] Anonymous. [members of the American Statistical Association]. Journal of the American Statistical Association, 21(156):5–42, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276766.

Anonymous:1926:MNa

[Ano26h] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 21(153):73–77, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965505.

Anonymous:1926:MNb

[Ano26i] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 21(155):354–357, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277069.

Anonymous:1926:MNc

[Ano26j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 21(156):489–491, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276988.

Anonymous:1926:PT

[Ano26k] Anonymous. The price trend. Journal of the American Statistical Association, 21(156):480–485, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276986.

Anonymous:1926:PES

[Ano26l] Anonymous. Proceedings of the Eighty-Seventh Annual Meeting of the American Statistical Association. Journal of the American Statistical Association, 21(153):78–93, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965506.

Anonymous:1926:PWC

[Ano26m] Anonymous. Progress of work in the Census Bureau. Journal of the American Statistical Association, 21(155):353–354, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277068.

Anonymous:1926:QPA

[Ano26n] Anonymous. Quarterly publications of the American statistical association. Journal of the American Statistical Association, 21(156):43– 65, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276767.

Anonymous:1926:RLa

[Ano26o] Anonymous. Recent literature. Journal of the American Statistical Association, 21(153):113–118, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965516.

Anonymous:1926:RLb

[Ano26p] Anonymous. Recent literature. Journal of the American Statistical Association, 21(154):251–255, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277161.

Anonymous:1926:RLc

[Ano26q] Anonymous. Recent literature. Journal of the American Statistical Association, 21(155):375–379, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277080.

Anonymous:1926:RLd

[Ano26r] Anonymous. Recent literature. Journal of the American Statistical Association, 21(156):513–517, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276998.

Anonymous:1926:SQ

[Ano26s] Anonymous. Sales quotas. Journal of the American Statistical Association, 21(154):220–224, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277150.

Anonymous:1926:VI

[Ano26t] Anonymous. Volume information. Journal of the American Statistical Association, 21(153):1–6, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965491.

Anonymous:1927:ANC

[Ano27a] Anonymous. An addition to the new composite index of wages in the United States. Journal of the American Statistical Association, 22(158):229, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276781.

Anonymous:1927:BM

[Ano27b] Anonymous. Back matter. Journal of the American Statistical Association, 22(157):??, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277368.

Anonymous:1927:CDR

[Ano27c] Anonymous. Can dependency and relief be measured? Journal of the American Statistical Association, 22(158):231–234, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276783.

Anonymous:1927:FIM

[Ano27d] Anonymous. Factors influencing the movements of security prices. Journal of the American Statistical Association, 22(159):381–385, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276810.

Anonymous:1927:FVO

[Ano27e] Anonymous. Forecasting the volume of orders or sales. Journal of the American Statistical Association, 22(158):235–238, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276784.

Anonymous:1927:FM

[Ano27f] Anonymous. Front matter. Journal of the American Statistical Association, 22(157):2, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277342.

Anonymous:1927:MASa

[Ano27g] Anonymous. Members added since December, 1926. Journal of the American Statistical Association, 22(157):86–88, March 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277355.

Anonymous:1927:MASc

[Ano27h] Anonymous. Members added since June, 1927. Journal of the American Statistical Association, 22(159):390–393, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276814.

Anonymous:1927:MASb

[Ano27i] Anonymous. Members added since March, 1927. Journal of the American Statistical Association, 22(158):242–246, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276788.

Anonymous:1927:MASd

[Ano27j] Anonymous. Members added since September, 1927. Journal of the American Statistical Association, 22(160):520–521, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276842.

Anonymous:1927:MNa

[Ano27k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 22(157):82–86, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277354.

Anonymous:1927:MNb

[Ano27l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 22(158):238–239, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276785.

Anonymous:1927:MNc

[Ano27m] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 22(159):386–389, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276811.

Anonymous:1927:MNd

[Ano27n] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 22(160):517–518, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276840.

Anonymous:1927:ONa

[Ano27o] Anonymous. Obituary notes. Journal of the American Statistical Association, 22(158):239–240, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276786.

Anonymous:1927:ONb

[Ano27p] Anonymous. Obituary notes. Journal of the American Statistical Association, 22(159):390, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276813.

Anonymous:1927:PNa

[Ano27q] Anonymous. Personal notes. Journal of the American Statistical Association, 22(158):240–242, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276787.

Anonymous:1927:PNb

[Ano27r] Anonymous. Personal notes. Journal of the American Statistical Association, 22(159):389–390, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276812.

Anonymous:1927:PNc

[Ano27s] Anonymous. Personal notes. Journal of the American Statistical Association, 22(160):518–520, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276841.

Anonymous:1927:PEE

[Ano27t] Anonymous. Proceedings of the Eighty-Eighth Annual Meeting of the American Statistical Association. Journal of the American Statistical Association, 22(157):89–108, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277356.

Anonymous:1927:PWC

[Ano27u] Anonymous. Progress of work in the Census Bureau: The census of religious bodies: 1926. Journal of the American Statistical Association, 22(159):380–381, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276809.

Anonymous:1927:RLa

[Ano27v] Anonymous. Recent literature. Journal of the American Statistical Association, 22(157):130–133, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277367.

Anonymous:1927:RLb

[Ano27w] Anonymous. Recent literature. Journal of the American Statistical Association, 22(158):265–270, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276798.

Anonymous:1927:RLc

[Ano27x] Anonymous. Recent literature. Journal of the American Statistical Association, 22(159):413–417, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276825.

Anonymous:1927:RLd

[Ano27y] Anonymous. Recent literature. Journal of the American Statistical Association, 22(160):546–550, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276853.

Anonymous:1927:SSR

[Ano27z] Anonymous. The social science research council fellowship appointments. Journal of the American Statistical Association, 22 (160):508–510, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276835.

Anonymous:1927:SDC

[Ano27-27] Anonymous. Statistical developments in China. Journal of the American Statistical Association, 22(160):514–515, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276838.

Anonymous:1927:U

[Ano27-28] Anonymous. [untitled]. Journal of the American Statistical Association, 22(157):126–128, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277364.

Anonymous:1927:VI

[Ano27-29] Anonymous. Volume information. Journal of the American Statistical Association, 22(157):1–6, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277341.

Anonymous:1928:CLA

[Ano28a] Anonymous. Constitution and by-laws of the American statistical association. Journal of the American Statistical Association, 23(164):4– 5, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276947.

Anonymous:1928:FMa

[Ano28b] Anonymous. Front matter. Journal of the American Statistical Association, 23(161):2, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277396.

Anonymous:1928:FMb

[Ano28c] Anonymous. Front matter. Journal of the American Statistical Association, 23(162):??, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276883.

Anonymous:1928:FMc

[Ano28d] Anonymous. Front matter. Journal of the American Statistical Association, 23(164):1–2, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276945.

Anonymous:1928:IRP

[Ano28e] Anonymous. Interim report of the proceedings of the first general assembly of the international union for the scientific investigation of population problems. Journal of the American Statistical Association, 23(163):306–317, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276744.

Anonymous:1928:MASa

[Ano28f] Anonymous. Members added since December, 1927. Journal of the American Statistical Association, 23(161):86–88, March 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277415.

Anonymous:1928:MASc

[Ano28g] Anonymous. Members added since June, 1928. Journal of the American Statistical Association, 23(163):332–334, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276749.

Anonymous:1928:MASb

[Ano28h] Anonymous. Members added since March, 1928. Journal of the American Statistical Association, 23(162):200–204, June 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276901.

Anonymous:1928:MASe

[Ano28i] Anonymous. Members added since September, 1928. Journal of the American Statistical Association, 23(164):449–451, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277295.

Anonymous:1928:MASd

[Ano28j] Anonymous. [members of the American Statistical Association]. Journal of the American Statistical Association, 23(164):6–52, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276948.

Anonymous:1928:MNa

[Ano28k] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 23(161):81–86, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277413.

Anonymous:1928:MNb

[Ano28l] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 23(162):196–200, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276900.

Anonymous:1928:MNc

[Ano28m] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 23(163):324–332, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276748.

Anonymous:1928:MNd

[Ano28n] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 23(164):445–449, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277294.

Anonymous:1928:PP

[Ano28o] Anonymous. Past presidents. Journal of the American Statistical Association, 23(164):3, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276946.

Anonymous:1928:PAS

[Ano28p] Anonymous. [publications of the American statistical association]. Journal of the American Statistical Association, 23(164):53–80, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276949.

Anonymous:1928:RLa

[Ano28q] Anonymous. Recent literature. Journal of the American Statistical Association, 23(161):110–114, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277427.

Anonymous:1928:RLb

[Ano28r] Anonymous. Recent literature. Journal of the American Statistical Association, 23(162):234–239, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276915.

Anonymous:1928:RLc

[Ano28s] Anonymous. Recent literature. Journal of the American Statistical Association, 23(163):356–360, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276763.

Anonymous:1928:RLd

[Ano28t] Anonymous. Recent literature. Journal of the American Statistical Association, 23(164):476–480, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277311.

Anonymous:1928:SIB

[Ano28u] Anonymous. Statistics of the insurance business. Journal of the American Statistical Association, 23(163):318–320, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276745.

Anonymous:1928:VI

[Ano28v] Anonymous. Volume information. Journal of the American Statistical Association, 23(161):1–8, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277395.

Anonymous:1929:E

[Ano29a] Anonymous. Errata. Journal of the American Statistical Association, 24(166):215, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276698.

Anonymous:1929:FSS

[Ano29b] Anonymous. Fellowships of the social science research council. Journal of the American Statistical Association, 24(166):187–189, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276692.

Anonymous:1929:MASa

[Ano29c] Anonymous. Members added since December, 1928. Journal of the American Statistical Association, 24(165):82–85, March 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277326.

Anonymous:1929:MASc

[Ano29d] Anonymous. Members added since June, 1929. Journal of the American Statistical Association, 24(167):327–329, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276867.

Anonymous:1929:MASb

[Ano29e] Anonymous. Members added since March, 1929. Journal of the American Statistical Association, 24(166):203–206, June 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276696.

Anonymous:1929:MASd

[Ano29f] Anonymous. Members added since September, 1929. Journal of the American Statistical Association, 24(168):435–436, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277093.

Anonymous:1929:MNa

[Ano29g] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 24(165):76–80, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277324.

Anonymous:1929:MNb

[Ano29h] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 24(166):192–199, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276694.

Anonymous:1929:MNc

[Ano29i] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 24(167):316–325, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276865.

Anonymous:1929:MNd

[Ano29j] Anonymous. Miscellaneous notes. Journal of the American Statistical Association, 24(168):431–434, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277091.

Anonymous:1929:PNa

[Ano29k] Anonymous. Personal notes. Journal of the American Statistical Association, 24(167):326–327, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276866.

Anonymous:1929:PNb

[Ano29l] Anonymous. Personal notes. Journal of the American Statistical Association, 24(168):434–435, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277092.

Anonymous:1929:RLa

[Ano29m] Anonymous. Recent literature. Journal of the American Statistical Association, 24(166):231, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276708.

Anonymous:1929:RLb

[Ano29n] Anonymous. Recent literature. Journal of the American Statistical Association, 24(167):348, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276882.

Anonymous:1929:RLc

[Ano29o] Anonymous. Recent literature. Journal of the American Statistical Association, 24(168):456, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277107.

Anonymous:1929:UIF

[Ano29p] Anonymous. Unemployment inquiry in the fifteenth census. Journal of the American Statistical Association, 24(167):314–316, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276864.

Anonymous:1929:VI

[Ano29q] Anonymous. Volume information. Journal of the American Statistical Association, 24(165):1–9, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277312.

Anonymous:1930:FM

[Ano30a] Anonymous. Front matter. Journal of the American Statistical Association, 25(169):??, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277187.

Anonymous:1930:MAB

[Ano30b] Anonymous. Minutes of the Annual Business Meeting. Journal of the American Statistical Association, 25(169):177–198, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277225.

Anonymous:1930:NFA

[Ano30c] Anonymous. Ninety-First Annual Meeting of the American Statistical Association held at Hotel Washington, Washington, D.C. Friday, December 27, to Monday, December 30, 1929. Journal of the American Statistical Association, 25(169):1–6, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277188.

Applegate:1925:BRB

[App25] Melbourne S. Applegate. Book review: [untitled]. Journal of the American Statistical Association, 20(150):295–296, June 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277137.

Arner:1925:INN

[Arn25] George B. L. Arner. An index number of net foreign trade in foodstuffs. Journal of the American Statistical Association, 20 (152):537–542, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277168.

Arner:1929:BRB

[Arn29] G. B. L. Arner. Book review: [untitled]. Journal of the American Statistical Association, 24(168):450–452, December 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277103.

Ashby:1928:AIF

[Ash28] Forrest Bee Ashby. An aggregative index of farm purchasing power. Journal of the American Statistical Association, 23(161):49– 54, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277403.

Ashby:1929:CFF

[Ash29] Forrest Bee Ashby. Cotton futures as forecasters of cotton spot prices. Journal of the American Statistical Association, 24(168):412–419, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277087.

Austin:1930:CA

[Aus30] William Lane Austin. The census of agriculture. Journal of the American Statistical Association, 25(169):130–134, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277215.

Ayres:1922:NSB

[Ayr22] Leonard P. Ayres. The nature and status of business research. Journal of the American Statistical Association, 18(137):33–39, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277465.

Ayres:1923:BRB

[Ayr23] Leonard P. Ayres. Book review: [untitled]. Journal of the American Statistical Association, 18(141):684–686, March 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277516.

Ayres:1927:DNS

[Ayr27] Leonard P. Ayres. The dilemma of the new statistics. Journal of the American Statistical Association, 22(157):1–8, March 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277343.

B:1928:ADA

[B.28a] R. W. B. The analysis of data among business organizations. Journal of the American Statistical Association, 23(162):184–187, June 1928.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276894.

B:1928:IPA

[B.28b] R. W. B. Indexes of production of agricultural products. Journal of the American Statistical Association, 23(162):193–195, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276899.

B:1929:BCS

[B.29] R. W. B. Bank credit and the security market. Journal of the American Statistical Association, 24(165):72–74, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277322.

Babb:1928:MI

[Bab28] J. J. Babb. Migration of industry. Journal of the American Statistical Association, 23(163):320–321, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276746.

Bachelor:1928:FTC

[Bac28a] Robert W. Bachelor. Freight traffic of California cities. Journal of the American Statistical Association, 23(164):440–441, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277291.

Bacon:1928:SFB

[Bac28b] Dorothy C. Bacon. The significance of fixed-base and link relatives in studies of price stability — a comment on the behavior of prices. Journal of the American Statistical Association, 23(163):274–281, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276738.

Bailey:1923:BRB

[Bai23] William B. Bailey. Book review: [untitled]. Journal of the American Statistical Association, 18(142):818–819, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276729.

Baker:1923:TAM

[Bak23] Gladden W. Baker. The trend of adult mortality in the United States. Journal of the American Statistical Association, 18(143):852–865, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965468.

Banker:1927:CFS

[Ban27] Howard J. Banker. The competitive factor in skewness. Journal of the American Statistical Association, 22(157):9–21, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277344.

Barnett:1922:BRB

[Bar22] George E. Barnett. Book review: [untitled]. Journal of the American Statistical Association, 18(139):407–409, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276965.

Barnett:1924:BRB

[Bar24] George E. Barnett. Book review: [untitled]. Journal of the American Statistical Association, 19(145):108–110, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277272.

Barnett:1928:BRB

[Bar28] George E. Barnett. Book review: [untitled]. Journal of the American Statistical Association, 23(161):100–101, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277419.

Bauman:1928:TMR

[Bau28] A. O. Bauman. Thirteen-months-ratio-first-difference method of measuring seasonal variation. Journal of the American Statistical Association, 23(163):282–290, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276739.

Baumgartner:1929:BRB

[Bau29] Rudolf Baumgartner. Book review: [untitled]. Journal of the American Statistical Association, 24(167):340–341, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276875.

Baxter:1923:BRB

[Bax23] W. J. Baxter. Book review: [untitled]. Journal of the American Statistical Association, 18(143):939–941, September 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965487.

Burgess:1928:BRBa

[BB28] Robert W. Burgess and W. Randolph Burgess. Book review: [untitled]. Journal of the American Statistical Association, 23(161):106– 107, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277423.

Bean:1929:BRB

[Bea29a] L. H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 24(165):89–90, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277328.

Bean:1929:SMG

[Bea29b] L. H. Bean. A simplified method of graphic curvilinear correlation. Journal of the American Statistical Association, 24(168):386–397, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277084.

Beales:1930:CM

[Bea30a] Le Verne Beales. The census of manufactures. Journal of the American Statistical Association, 25(169):121–124, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277213.

Bean:1930:APW

[Bea30b] Louis H. Bean. Agriculture in the post-War decade. Journal of the American Statistical Association, 25(169):155–157, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277220.

Belcher:1924:SSS

[Bel24] Donald R. Belcher. The supervising or staff statistician. Journal of the American Statistical Association, 19(145):61–65, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277259.

Bennett:1929:BRB

[Ben29] M. K. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 24(165):98–99, March 1929. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277333.

Berridge:1922:CEUb

[Ber22a] W. A. Berridge. Cycles of employment and unemployment in the United States, 1903–1914. Journal of the American Statistical Association, 18(138):227–240, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277529.

Berridge:1922:CEUa

[Ber22b] W. A. Berridge. Cycles of employment and unemployment in the United States, 1914–1921. Journal of the American Statistical Association, 18(137):42–55, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277467.

Berridge:1922:SER

[Ber22c] W. A. Berridge. Statistics on employment: a rejoinder. Journal of the American Statistical Association, 18(140):519–522, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276928.

Bernhardt:1923:BRB

[Ber23] Joshua Bernhardt. Book review: [untitled]. Journal of the American Statistical Association, 18(141):690–693, March 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277519.

Berridge:1924:BRB

[Ber24] W. A. Berridge. Book review: [untitled]. Journal of the American Statistical Association, 19(146):265–266, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277249.

Berridge:1925:BRB

[Ber25] William A. Berridge. Book review: [untitled]. Journal of the American Statistical Association, 20(151):460–461, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965542.

Berridge:1926:BRB

[Ber26] W. A. Berridge. Book review: [untitled]. Journal of the American Statistical Association, 21(156):509–510, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276996.

Berridge:1927:BRB

[Ber27] W. A. Berridge. Book review: [untitled]. Journal of the American Statistical Association, 22(159):405–406, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276820.

Bernheim:1928:BRB

[Ber28] Alfred L. Bernheim. Book review: [untitled]. Journal of the American Statistical Association, 23(164):468–470, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277306.

Berridge:1929:BRB

[Ber29] W. A. Berridge. Book review: [untitled]. Journal of the American Statistical Association, 24(166):224–225, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276704.

Beyer:1927:BRB

[Bey27] William C. Beyer. Book review: [untitled]. Journal of the American Statistical Association, 22(160):528–532, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276846.

Beyleveld:1929:DPI

[Bey29] A. J. Beyleveld. Determination of a precise indication of change in crop acreage. Journal of the American Statistical Association, 24(168):405–411, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277086.

Bezanson:1924:BRB

[Bez24] Anne Bezanson. Book review: [untitled]. Journal of the American Statistical Association, 19(147):410–411, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277385.

Belcher:1923:BRB

[BF23] Donald R. Belcher and Harold M. Flinn. Book review: [untitled]. Journal of the American Statistical Association, 18(143):928–931, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965482.

Black:1928:BRB

[Bla28a] John D. Black. Book review: [untitled]. Journal of the American Statistical Association, 23(162):211–213, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276904.

Blanchard:1928:BRB

[Bla28b] Ralph H. Blanchard. Book review: [untitled]. Journal of the American Statistical Association, 23(161):97–100, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277418.

Blanchard:1929:BRB

[Bla29] Ralph H. Blanchard. Book review: [untitled]. Journal of the American Statistical Association, 24(167):344–345, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276878.

Blood:1923:LGF

[Blo23] John Balch Blood. Law of growth in forecasting demand. Journal of the American Statistical Association, 18(141):651–653, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277506.

Bloch:1927:BRB

[Blo27] Louis Bloch. Book review: [untitled]. Journal of the American Statistical Association, 22(159):406–408, September 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276821.

Bloch:1929:FAM

[Blo29a] Louis Bloch. Facts about Mexican immigration before and since the quota restriction laws. Journal of the American Statistical Association, 24(165):50–60, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277317.

Blodgett:1929:BRB

[Blo29b] Thurston P. Blodgett. Book review: [untitled]. Journal of the American Statistical Association, 24(167):345, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276879.

Boas:1922:MDB

[Boa22a] Franz Boas. The measurement of differences between variable quantities. Journal of the American Statistical Association, 18 (140):425–445, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276916.

Boas:1922:RAI

[Boa22b] Franz Boas. Report on an anthropometric investigation of the population of the United States. Journal of the American Statistical Association, 18(138):181–209, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277526.

Bosland:1926:FPW

[Bos26] C. C. Bosland. Forecasting the price of wheat. Journal of the American Statistical Association, 21(154):149–161, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277142.

Bowley:1928:ESS

[Bow28a] A. L. Bowley. Economic and social survey of London. Journal of the American Statistical Association, 23(164):440, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277290.

Bowley:1928:SDC

[Bow28b] A. L. Bowley. The standard deviation of the correlation coefficient. Journal of the American Statistical Association, 23(161):31–34, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277400.

Brandt:1928:UMF

[Bra28] A. E. Brandt. The use of machine factoring in multiple correlation. Journal of the American Statistical Association, 23(163):291–295, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276740.

Breiby:1923:BRB

[Bre23] William Breiby. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1059–1064, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965674.

Brissenden:1922:BRB

[Bri22] Paul F. Brissenden. Book review: [untitled]. Journal of the American Statistical Association, 18(138):294–296, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277549.

Brissenden:1923:OIL

[Bri23] Paul F. Brissenden. Occupational incidence of labor mobility. Journal of the American Statistical Association, 18(144):978–992, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965659.

Brissenden:1929:BRB

[Bri29] P. F. Brissenden. Book review: [untitled]. Journal of the American Statistical Association, 24(165):90–92, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277329.

Brown:1925:LDT

[Bro25] Theodore H. Brown. The law of demand and the theory of probability. Journal of the American Statistical Association, 20(150):223–230, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277116.

Brown:1928:BRB

[Bro28] J. Douglas Brown. Book review: [untitled]. Journal of the American Statistical Association, 23(164):466–468, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277305.

Brown:1929:BRB

[Bro29] Harry Gunnison Brown. Book review: [untitled]. Journal of the American Statistical Association, 24(165):92–93, March 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277330.

Bean:1924:FTI

[BS24] L. H. Bean and O. C. Stine. Four types of index numbers of farm prices. Journal of the American Statistical Association, 19(145):30– 35, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277256.

Buck:1922:BRB

[Buc22] George B. Buck. Book review: [untitled]. Journal of the American Statistical Association, 18(139):412–414, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276968.

Buck:1925:PCC

[Buc25] J. Lossing Buck. Price changes in China: The effect of famines and the recent rise in prices. Journal of the American Statistical Association, 20(150):238–241, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277118.

Burgess:1922:INW

[Bur22] W. Randolph Burgess. Index numbers for the wages of common labor. Journal of the American Statistical Association, 18(137):101– 103, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277474.

Burgess:1923:BRB

[Bur23a] W. Randolph Burgess. Book review: [untitled]. Journal of the American Statistical Association, 18(141):688–689, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277518.

Burgess:1923:VBD

[Bur23b] W. Randolph Burgess. Velocity of bank deposits. Journal of the American Statistical Association, 18(142):727–740, June 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276711.

Burgess:1927:FAC

[Bur27] W. Randolph Burgess. Factors affecting changes in short terms interest rates. Journal of the American Statistical Association, 22(158):195–201, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276772.

Burgess:1928:BRBb

[Bur28] Robert W. Burgess. Book review: [untitled]. Journal of the American Statistical Association, 23(163):335–336, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276750.

Burgess:1929:BRB

[Bur29a] W. Randolph Burgess. Book review: [untitled]. Journal of the American Statistical Association, 24(165):100, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277334.

Burns:1929:GMP

[Bur29b] Arthur F. Burns. The geometric mean of percentages. Journal of the American Statistical Association, 24(167):296–300, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276859.

Burns:1929:RIC

[Bur29c] Arthur F. Burns. The relative importance of check and cash payments in the United States, 1919–1928. Journal of the American Statistical Association, 24(168):420–424, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277088.

Burns:1929:SEF

[Bur29d] E. M. Burns. Statistics and economic forecasting. Journal of the American Statistical Association, 24(166):152–163, June 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276686.

C:1923:BRBa

[C.23a] W. L. C. Book review: [untitled]. Journal of the American Statistical Association, 18(141):678–682, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277514.

C:1923:BRBb

[C.23b] W. L. C. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1057–1059, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965673.

Camp:1923:PS

[Cam23] Burton H. Camp. Problems in sampling. Journal of the American Statistical Association, 18(144):964–977, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965658.

Camp:1928:DCR

[Cam28] C. C. Camp. Devices for computing the rates together with a statistical study of building and loan associations. Journal of the American Statistical Association, 23(164):411–416, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277286.

Canning:1926:FFD

[Can26] John B. Canning. The formation of frequency distributions. Journal of the American Statistical Association, 21(154):133–148, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277141.

Carver:1922:BRB

[Car22] Harry C. Carver. Book review: [untitled]. Journal of the American Statistical Association, 18(137):148, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277493.

Carr:1924:CLS

[Car24] Elma B. Carr. Cost of living statistics of U.S. Bureau of Labor Statistics and the National Industrial Conference Board: Comparison of methods used in the collection and computation of cost of living figures. Journal of the American Statistical Association, 19 (148):484–507, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281565.

Carmichael:1927:MCS

[Car27] F. L. Carmichael. Methods of computing seasonal indexes: Constant and progressive. Journal of the American Statistical Association, 22(159):339–354, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276803.

Carmichael:1928:ATT

[Car28a] F. L. Carmichael. The arc tangent in trend determination. Journal of the American Statistical Association, 23(163):253–262, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276736.

Carroll:1928:RIG

[Car28b] Mollie Ray Carroll. Recent improvements in German social statistics. Journal of the American Statistical Association, 23(164):386–397, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277284.

Carpenter:1929:BRB

[Car29] Niles Carpenter. Book review: [untitled]. Journal of the American Statistical Association, 24(168):446–447, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277099.

Catlin:1928:BRB

[Cat28] Warren B. Catlin. Book review: [untitled]. Journal of the American Statistical Association, 23(164):464–466, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277304.

Chaddock:1922:MN

[Cha22] Robert E. Chaddock. Miscellaneous notes. Journal of the American Statistical Association, 18(140):528–533, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276932.

Chapin:1923:SVD

[Cha23] F. Stuart Chapin. Seasonal variation of dependency. Journal of the American Statistical Association, 18(143):911–914, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965475.

Chaddock:1924:PEF

[Cha24] Robert E. Chaddock. Proceedings of the Eighty-Fifth Annual Meeting of the American Statistical Association. Journal of the American Statistical Association, 19(145):86–95, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277267.

Chambers:1925:IP

[Cha25] C. R. Chambers. An index of profits. Journal of the American Statistical Association, 20(149):104–107, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277440.

Chaddock:1926:CPC

[Cha26a] Robert E. Chaddock. [A criticism of Professor Chaddock’s “Principles and Methods of Statistics”: a reply]. Journal of the American Statistical Association, 21(155):350, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277065.

Chaddock:1926:FSU

[Cha26b] Robert E. Chaddock. The function of statistics in undergraduate training. Journal of the American Statistical Association, 21(153): 1–8, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965492.

Chambers:1928:BRB

[Cha28a] C. R. Chambers. Book review: [untitled]. Journal of the American Statistical Association, 23(163):351, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276759.

Chapin:1928:SSA

[Cha28b] F. Stuart Chapin. Social science abstracts. Journal of the American Statistical Association, 23(162):187–188, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276895.

Chatfield:1930:SAF

[Cha30] Charlotte Chatfield. Statistical analysis in food composition studies. Journal of the American Statistical Association, 25(169):153–154, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277219.

Claghorn:1923:BRB

[Cla23a] Kate Holladay Claghorn. Book review: [untitled]. Journal of the American Statistical Association, 18(142):829–830, June 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276733.

Clark:1923:BRB

[Cla23b] W. C. Clark. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1053–1056, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965672.

Clark:1924:BRB

[Cla24] J. M. Clark. Book review: Railroads; Rates–Service–Management, by Homer Bews Vanderblue; Kenneth Farwell. Journal of the American Statistical Association, 19(146):259–262, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277246.

Clark:1925:BRBb

[Cla25a] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 20(151):458–460, September 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965540.

Clark:1925:BRBa

[Cla25b] W. C. Clark. Book review: [untitled]. Journal of the American Statistical Association, 20(151):456–458, September 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965539.

Clark:1925:NAP

[Cla25c] Percy L. Clark, Jr. A note on almshouse population. Journal of the American Statistical Association, 20(149):107–109, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277441.

Clark:1926:BRB

[Cla26] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 21(156):506–507, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276994.

Clark:1928:BRB

[Cla28] J. M. Clark. Book review: [untitled]. Journal of the American Statistical Association, 23(162):220–224, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276909.

Clark:1929:BRB

[Cla29] Fred E. Clark. Book review: [untitled]. Journal of the American Statistical Association, 24(168):454–455, December 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277106.

Clay:1930:F

[Cla30] Paul Clay. Forecasts for 1930. Journal of the American Statistical Association, 25(169):97–99, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277207.

Clendenin:1927:MVS

[Cle27] J. C. Clendenin. Measurement of variations in seasonal distribution. Journal of the American Statistical Association, 22(158):213– 220, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276775.

Coats:1924:BRB

[Coa24] R. H. Coats. Book review: [untitled]. Journal of the American Statistical Association, 19(148):534–539, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281575.

Coats:1925:CPO

[Coa25] R. H. Coats. The Canadian plan of organization of government statistics. Journal of the American Statistical Association, 20(149):90–93, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277437.

Commons:1925:BRB

[Com25] John R. Commons. Book review: [untitled]. Journal of the American Statistical Association, 20(149):138–139, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277449.

Copeland:1922:BRB

[Cop22] Melvin T. Copeland. Book review: [untitled]. Journal of the American Statistical Association, 18(138):296–297, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277550.

Copeland:1928:ETV

[Cop28] Morris A. Copeland. An estimate of total volume of debits to individual accounts in the United States. Journal of the American Statistical Association, 23(163):301–303, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276742.

Copeland:1929:BRB

[Cop29a] Melvin T. Copeland. Book review: [untitled]. Journal of the American Statistical Association, 24(168):447–449, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277100.

Copeland:1929:SPI

[Cop29b] Morris A. Copeland. Special purpose indexes for the equation of exchange for the United States, 1919–1927. Journal of the American Statistical Association, 24(166):109–122, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276682.

Copeland:1930:RCO

[Cop30] Morris A. Copeland. Recent changes in our wholesale price level. Journal of the American Statistical Association, 25(169):164–169, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277222.

Cox:1930:EEF

[Cox30] Garfield V. Cox. Evaluation of economic forecasts. Journal of the American Statistical Association, 25(169):31–35, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277193.

Crathorne:1922:CCR

[Cra22] A. R. Crathorne. Calculation of the correlation ratio. Journal of the American Statistical Association, 18(139):394–396, September 1922.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276960.

Crick:1922:BBS

[Cri22] W. F. Crick. British banking statistics. Journal of the American Statistical Association, 18(140):506–514, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276924.

Croxton:1922:PP

[Cro22] Frederick E. Croxton. A percentage protractor. Journal of the American Statistical Association, 18(137):108–109, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277477.

Croxton:1925:AAC

[Cro25] Frederick E. Croxton. An apparatus to assist in the calculation of the standard deviation of a grouped frequency distribution. Journal of the American Statistical Association, 20(152):532–536, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277167.

Croxton:1928:BRB

[Cro28] Frederick E. Croxton. Book review: [untitled]. Journal of the American Statistical Association, 23(164):455–456, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277298.

Crum:1922:DST

[Cru22] W. L. Crum. The determination of secular trend. Journal of the American Statistical Association, 18(138):210–215, June 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277527.

Crum:1923:RBO

[Cru23a] W. L. Crum. The resemblance between the ordinate of the periodogram and the correlation coefficient. Journal of the American Statistical Association, 18(143):889–899, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965471.

Crum:1923:UMD

[Cru23b] W. L. Crum. The use of the median in determining seasonal variation. Journal of the American Statistical Association, 18(141):607– 614, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277501.

Crum:1925:LSC

[Cru25a] W. L. Crum. The least squares criterion for trend lines. Journal of the American Statistical Association, 20(150):211–222, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277115.

Crum:1925:PVS

[Cru25b] W. L. Crum. Progressive variation in seasonality. Journal of the American Statistical Association, 20(149):48–64, March 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277433.

Crum:1926:SAJ

[Cru26] W. L. Crum. The statistical allocation of joint costs. Journal of the American Statistical Association, 21(153):9–24, March 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965493.

Croxton:1927:BCV

[CS27] Frederick E. Croxton and Roy E. Stryker. Bar charts versus circle diagrams. Journal of the American Statistical Association, 22 (160):473–482, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276829.

Comer:1927:FLI

[CW27] H. D. Comer and Ralph J. Watkins. Forecasting a line by itself. Journal of the American Statistical Association, 22(160):505–507, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276834.

Davis:1923:BRB

[Dav23] Joseph S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 18(143):936–937, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965485.

Davenport:1924:BRB

[Dav24a] Charles B. Davenport. Book review: [untitled]. Journal of the American Statistical Association, 19(146):264, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277248.

Davies:1924:PSI

[Dav24b] George R. Davies. The problem of a standard index number formula. Journal of the American Statistical Association, 19(146):180– 188, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277229.

Davies:1925:LCD

[Dav25] George R. Davies. The logarithmic curve of distribution. Journal of the American Statistical Association, 20(152):467–480, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277162.

Davis:1926:SPL

[Dav26a] J. S. Davis. A scale for plotting logarithmic curves. Journal of the American Statistical Association, 21(155):353, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277067.

Davis:1926:PFD

[Dav26b] William H. Davis. The past and future development of vital statistics in the United States: IV, 1926 progress report in the campaign to bring every state into the registration. Journal of the American Statistical Association, 21(155):274–279, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277054.

Davies:1927:GC

[Dav27a] George R. Davies. The growth curve. Journal of the American Statistical Association, 22(159):370–374, September 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276807.

Davis:1927:BRB

[Dav27b] Joseph S. Davis. Book review: [untitled]. Journal of the American Statistical Association, 22(157):109–115, March 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277357.

Davenport:1928:BRBb

[Dav28a] C. B. Davenport. Book review: [untitled]. Journal of the American Statistical Association, 23(163):344, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276754.

Davenport:1928:BRBa

[Dav28b] Charles B. Davenport. Book review: [untitled]. Journal of the American Statistical Association, 23(161):101–102, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277420.

Davis:1928:LTG

[Dav28c] Robert M. Davis. Long-time guarantees of prosperity. Journal of the American Statistical Association, 23(162):137–143, June 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276886.

Davies:1929:AFD

[Dav29] G. R. Davies. The analysis of frequency distributions. Journal of the American Statistical Association, 24(168):349–366, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277081.

Dawson:1924:BRB

[Daw24] Miles M. Dawson. Book review: [untitled]. Journal of the American Statistical Association, 19(148):550–553, December 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281579.

Dawson:1928:BRB

[Daw28] Owen L. Dawson. Book review: [untitled]. Journal of the American Statistical Association, 23(164):474, December 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277309.

Day:1928:RSB

[Day28] Edmund E. Day. The role of statistics in business forecasting. Journal of the American Statistical Association, 23(161):1–9, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277397.

Delatour:1923:EPA

[Del23] M. Delatour. Extracts from Presidential Address of M. Delatour at the Fifteenth Meeting of the International Statistical Institute at Brussels, October 1, 1923. Journal of the American Statistical Association, 18(144):1038–1039, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965665.

Dennison:1922:MBC

[Den22] Henry S. Dennison. Management and the business cycle. Journal of the American Statistical Association, 18(137):20–31, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277463.

Dexter:1923:FCI

[Dex23] Robert C. Dexter. French–Canadian intermarriage in New England. Journal of the American Statistical Association, 18(143):914–915, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965476.

Dickinson:1929:BRB

[Dic29] Z. Clark Dickinson. Book review: [untitled]. Journal of the American Statistical Association, 24(166):221–223, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276703.

Dublin:1925:TRN

[DL25] Louis I. Dublin and Alfred J. Lotka. On the true rate of natural increase. Journal of the American Statistical Association, 20 (151):305–339, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965517.

Doten:1922:SSE

[Dot22] Carroll W. Doten. Statistics in the service of economics. Journal of the American Statistical Association, 18(137):1–7, March 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277461.

Douglas:1922:WSN

[Dou22] Paul H. Douglas. Wage studies of the national industrial conference board: a rejoinder. Journal of the American Statistical Association, 18(138):263–266, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277535.

Douglas:1923:ASS

[Dou23] Paul H. Douglas. An analysis of strike statistics, 1881–1921. Journal of the American Statistical Association, 18(143):866–877, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965469.

Douglas:1924:BRB

[Dou24a] Paul H. Douglas. Book review: [untitled]. Journal of the American Statistical Association, 19(147):424–425, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277393.

Douglas:1924:FFT

[Dou24b] Paul H. Douglas. Is the family of five typical? Journal of the American Statistical Association, 19(147):314–328, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277371.

Downey:1922:BRB

[Dow22] E. H. Downey. Book review: [untitled]. Journal of the American Statistical Association, 18(138):283–284, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277543.

Dowrie:1925:BRB

[Dow25] George W. Dowrie. Book review: [untitled]. Journal of the American Statistical Association, 20(152):590–591, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277184.

Dowrie:1928:BRB

[Dow28] George W. Dowrie. Book review: [untitled]. Journal of the American Statistical Association, 23(164):470–472, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277307.

Downes:1929:ARB

[Dow29] Jean Downes. The accuracy of the recorded birth statistics in urban and rural areas. Journal of the American Statistical Association, 24(165):15–27, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277314.

Dublin:1922:WDR

[Dub22] Louis I. Dublin. The work of dreyer in relation to life insurance examinations. Journal of the American Statistical Association, 18(138):216–226, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277528.

Dublin:1924:BRB

[Dub24] Louis I. Dublin. Book review: [untitled]. Journal of the American Statistical Association, 19(146):249–253, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277243.

Dublin:1925:SPP

[Dub25] Louis I. Dublin. The statistician and the population problem. Journal of the American Statistical Association, 20(149):1–12, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277429.

Dublin:1926:PFD

[Dub26] Louis I. Dublin. The past and future development of vital statistics in the United States: V, the present registration system. Journal of the American Statistical Association, 21(155):279–281, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277055.

Duncan:1922:NSB

[Dun22] C. S. Duncan. The nature and status of business research: Discussion. Journal of the American Statistical Association, 18(137):39–41, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277466.

Durand:1930:FCS

[Dur30a] E. Dana Durand. Federal Corporation Statistics. Journal of the American Statistical Association, 25(169):23–30, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277192.

Durand:1930:FT

[Dur30b] E. Dana Durand. The foreign trade of 1929. Journal of the American Statistical Association, 25(169):51–55, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277196.

Edgeworth:1926:MN

[EB26] F. Y. Edgeworth and A. L. Bowley. Miscellaneous notes. Journal of the American Statistical Association, 21(154):224–230, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277151.

Edie:1930:CAI

[Edi30] Lionel D. Edie. Certain aspects of the interpretation of price-earnings ratios: Discussion. Journal of the American Statistical Association, 25(169):110–112, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277210.

Eells:1926:RMC

[Eel26] Walter Crosby Eells. The relative merits of circles and bars for representing component parts. Journal of the American Statistical Association, 21(154):119–132, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277140.

Eells:1929:FPE

[Eel29] Walter Crosby Eells. Formulas for probable errors of coefficients of correlation. Journal of the American Statistical Association, 24(166):170–173, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276688.

Effinger:1930:CES

[Eff30] Robert C. Effinger. Corporate earnings and stock prices. Journal of the American Statistical Association, 25(169):80–82, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277203.

Ellsworth:1926:NWI

[EH26] D. W. Ellsworth and Ruth Houghton. A new weekly index of commodity prices. Journal of the American Statistical Association, 21 (156):471–475, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276982.

Eichel:1922:LTC

[Eic22] Otto R. Eichel. The long-time cycles of pandemic influenza. Journal of the American Statistical Association, 18(140):446–454, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276917.

Elliott:1927:WID

[Ell27] D. C. Elliott. A weekly index of distribution. Journal of the American Statistical Association, 22(159):375–379, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276808.

Emerson:1925:BRB

[Eme25] Haven Emerson. Book review: [untitled]. Journal of the American Statistical Association, 20(149):129–133, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277446.

Evans:1925:BRB

[Eva25] Percy H. Evans. Book review: [untitled]. Journal of the American Statistical Association, 20(150):285–286, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277132.

Ezekiel:1924:MHC

[Eze24] Mordecai Ezekiel. A method of handling curvilinear correlation for any number of variables. Journal of the American Statistical Association, 19(148):431–453, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281561.

Ezekiel:1925:AIM

[Eze25] Mordecai Ezekiel. The assumptions implied in the multiple regression equation. Journal of the American Statistical Association, 20 (151):405–408, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965524.

Ezekiel:1926:BRB

[Eze26a] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 21(153):99–101, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965509.

Ezekiel:1926:DCR

[Eze26b] Mordecai Ezekiel. The determination of curvilinear regression “surfaces” in the presence of other variables. Journal of the American Statistical Association, 21(155):310–320, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277058.

Ezekiel:1927:BRB

[Eze27a] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 22(159):397–400, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276816.

Ezekiel:1927:TMF

[Eze27b] Mordecai Ezekiel. Two methods of forecasting hog prices. Journal of the American Statistical Association, 22(157):22–30, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277345.

Ezekiel:1928:BRB

[Eze28] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 23(162):205–208, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276902.

Ezekiel:1929:BRB

[Eze29] Mordecai Ezekiel. Book review: [untitled]. Journal of the American Statistical Association, 24(166):215, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276699.

Falkner:1924:MSV

[Fal24] Helen D. Falkner. The measurement of seasonal variation. Journal of the American Statistical Association, 19(146):167–179, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277228.

Falkner:1925:BRB

[Fal25a] Helen D. Falkner. Book review: [untitled]. Journal of the American Statistical Association, 20(149):140–141, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277450.

Falkner:1925:VCW

[Fal25b] Helen D. Falkner. Variation and correlation within the credit structure. Journal of the American Statistical Association, 20 (152):496–511, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277164.

Feldman:1925:BRB

[Fel25] H. Feldman. Book review: [untitled]. Journal of the American Statistical Association, 20(149):144–146, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277452.

Fletcher:1929:SAJ

[FF29] Ralph Fletcher and Mildred Fletcher. A statistical analysis of juvenile employment in St. Louis. Journal of the American Statistical Association, 24(166):174–177, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276689.

Field:1923:BRB

[Fie23] James A. Field. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1051–1053, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965671.

Fisher:1923:BRB

[Fis23a] Arne Fisher. Book review: [untitled]. Journal of the American Statistical Association, 18(142):816–818, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276728.

Fisher:1923:C

[Fis23b] Arne Fisher. Correspondence. Journal of the American Statistical Association, 18(142):793–795, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276721.

Fisher:1923:BCL

[Fis23c] Irving Fisher. The business cycle largely a “dance of the dollar”. Journal of the American Statistical Association, 18(144):1024–1028, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965663.

Fisher:1923:WIN

[Fis23d] Irving Fisher. A weekly index number of wholesale prices. Journal of the American Statistical Association, 18(143):835–840, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965466.

Fisher:1924:BRBb

[Fis24a] Arne Fisher. Book review: [untitled]. Journal of the American Statistical Association, 19(147):413–417, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277387.

Fisher:1924:BRBa

[Fis24b] Irving Fisher. Book review: [untitled]. Journal of the American Statistical Association, 19(145):123–125, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277278.

Fisher:1924:RWI

[Fis24c] Irving Fisher. Revision of the weekly index number. Journal of the American Statistical Association, 19(147):336–347, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277373.

Fisher:1925:OUD

[Fis25a] Irving Fisher. Our unstable dollar and the so-called business cycle. Journal of the American Statistical Association, 20(150):179–202, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277113.

Fisher:1925:RWI

[Fis25b] Irving Fisher. Revision of the weekly index number. Journal of the American Statistical Association, 20(151):409–413, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965525.

Fisher:1926:CPC

[Fis26a] Arne Fisher. A criticism of Professor Chaddock’s “Principles and Methods of Statistics”. Journal of the American Statistical Association, 21(155):347–350, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277064.

Fisher:1926:BRB

[Fis26b] Irving Fisher. Book review: [untitled]. Journal of the American Statistical Association, 21(153):110–112, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965514.

Fisher:1926:RWI

[Fis26c] Irving Fisher. Revision of the weekly index number. Journal of the American Statistical Association, 21(155):344–346, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277063.

Fisher:1927:BRB

[Fis27a] Arne Fisher. Book review: [untitled]. Journal of the American Statistical Association, 22(159):402–404, September 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276818.

Fisher:1927:RDW

[Fis27b] Irving Fisher. Report to date of the weekly index number. Journal of the American Statistical Association, 22(158):221–223, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276776.

Fisher:1927:TVC

[Fis27c] Irving Fisher. The “total value criterion”: a new principle in index number construction. Journal of the American Statistical Association, 22(160):419–441, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276826.

Fisher:1928:WWI

[Fis28] Irving Fisher. Weekly wholesale index: Sixth report. Journal of the American Statistical Association, 23(164):429–434, December 1928.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277288.

Fisher:1930:SMP

[Fis30] Irving Fisher. The stock market panic in 1929. Journal of the American Statistical Association, 25(169):93–96, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277206.

Fitch:1926:BRB

[Fit26] John A. Fitch. Book review: [untitled]. Journal of the American Statistical Association, 21(153):108–110, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965513.

Flinn:1924:BRB

[Fli24] Harold M. Flinn. Book review: [untitled]. Journal of the American Statistical Association, 19(145):111–113, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277273.

Foerster:1924:BRBa

[Foe24a] Robert F. Foerster. Book review: [untitled]. Journal of the American Statistical Association, 19(147):419–420, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277389.

Foerster:1924:BRBb

[Foe24b] Robert F. Foerster. Book review: [untitled]. Journal of the American Statistical Association, 19(148):553, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281580.

Forney:1928:SRS

[For28a] R. L. Forney. A standard reporting system for public accidents. Journal of the American Statistical Association, 23(161):79–80, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277412.

Forsyth:1928:MIS

[For28b] C. H. Forsyth. Monthly items of secular trend corresponding to known yearly items. Journal of the American Statistical Association, 23(163):296–300, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276741.

Frank:1923:LTP

[Fra23] Lawrence K. Frank. Long term price trends. Journal of the American Statistical Association, 18(143):904–908, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965473.

Frankel:1928:SSS

[Fra28] Emil Frankel. Some suggestions for securing standardized criminal statistics. Journal of the American Statistical Association, 23(161): 77–78, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277411.

Frain:1929:SOC

[Fra29] H. LaRue Frain. Some observations concerning the assumption of normal working time. Journal of the American Statistical Association, 24(167):276–287, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276857.

Fry:1927:ARA

[Fry27] C. Luther Fry. Answers to the review of “American Villagers”. Journal of the American Statistical Association, 22(158):227, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276778.

Fry:1928:IEO

[Fry28] C. Luther Fry. Illegal entry of orientals into the United States between 1910 and 1920. Journal of the American Statistical Association, 23(162):173–177, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276890.

Fuller:1928:OMB

[Ful28] Roden Fuller. Occupations of the Mexican-born population of Texas, New Mexico and Arizona, 1900–1920. Journal of the American Statistical Association, 23(161):64–67, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277407.

G:1925:BRB

[G.25] E. A. G. Book review: [untitled]. Journal of the American Statistical Association, 20(151):460, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965541.

Gadsby:1925:BRB

[Gad25] Margaret Gadsby. Book review: [untitled]. Journal of the American Statistical Association, 20(152):579–584, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277180.

Galvani:1927:NOS

[Gal27] L. Galvani. The new organization of the statistical services in Italy. Journal of the American Statistical Association, 22(159):355–358, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276804.

Gantt:1922:BRB

[Gan22] Margaret E. Gantt. Book review: [untitled]. Journal of the American Statistical Association, 18(138):290–292, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277546.

Gardner:1929:BRB

[Gar29] Walter R. Gardner. Book review: [untitled]. Journal of the American Statistical Association, 24(165):94–95, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277331.

Gavett:1927:BRB

[Gav27] G. Irving Gavett. Book review: [untitled]. Journal of the American Statistical Association, 22(160):523–527, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276844.

Gehlke:1925:BRB

[Geh25] C. E. Gehlke. Book review: [untitled]. Journal of the American Statistical Association, 20(149):150–152, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277455.

Genung:1930:AS

[Gen30] A. B. Genung. The 1929 agricultural season. Journal of the American Statistical Association, 25(169):58–63, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277198.

Gillman:1927:ARR

[Gil27] J. M. Gillman. Answer to the review of “rent levels and their causes”. Journal of the American Statistical Association, 22(160):515–516, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276839.

Glover:1924:BRB

[Glo24a] James W. Glover. Book review: [untitled]. Journal of the American Statistical Association, 19(145):125–126, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277279.

Glover:1924:WAC

[Glo24b] James W. Glover. Work of the Association’s committee on professional and educational standards for statisticians. Journal of the American Statistical Association, 19(147):399–404, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277382.

Glover:1926:BRB

[Glo26a] James W. Glover. Book review: [untitled]. Journal of the American Statistical Association, 21(155):371–372, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277077.

Glover:1926:RST

[Glo26b] James W. Glover. Requirements for statisticians and their training: Statistical teaching in American colleges and universities. Journal of the American Statistical Association, 21(156):419–446, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276978.

Glover:1927:BRB

[Glo27] James W. Glover. Book review: [untitled]. Journal of the American Statistical Association, 22(160):522–523, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276843.

Gressens:1927:VCT

[GM27] O. Gressens and E. D. Mouzon, Jr. The validity of correlation in time sequences and a new coefficient of similarity. Journal of the American Statistical Association, 22(160):483–492, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276830.

Goldenweiser:1924:BRB

[Gol24] E. A. Goldenweiser. Book review: [untitled]. Journal of the American Statistical Association, 19(145):120–123, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277277.

Goldenweiser:1927:BRB

[Gol27] E. A. Goldenweiser. Book review: [untitled]. Journal of the American Statistical Association, 22(158):249–252, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276790.

Goldenweiser:1930:CD

[Gol30] E. A. Goldenweiser. Credit developments in 1929. Journal of the American Statistical Association, 25(169):69–72, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277200.

Goodrich:1926:BRB

[Goo26] Carter Goodrich. Book review: [untitled]. Journal of the American Statistical Association, 21(155):372–373, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277078.

Goodrich:1929:BRB

[Goo29] Carter Goodrich. Book review: [untitled]. Journal of the American Statistical Association, 24(167):346–347, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276881.

Gottlieb:1925:PCR

[Got25] L. R. Gottlieb. Professor Common’s review of “The Tax Problem in Wisconsin”. Journal of the American Statistical Association, 20 (151):419–421, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965528.

Graham:1925:BRB

[Gra25] Frank D. Graham. Book review: [untitled]. Journal of the American Statistical Association, 20(150):294–295, June 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277136.

Graham:1928:BRB

[Gra28] Frank D. Graham. Book review: [untitled]. Journal of the American Statistical Association, 23(162):232–233, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276914.

Gressens:1925:MSV

[Gre25a] O. Gressens. On the measurement of seasonal variation. Journal of the American Statistical Association, 20(150):203–210, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277114.

Gressens:1925:QDF

[Gre25b] O. Gressens. The quantitative determination of fundamental changes in economic data. Journal of the American Statistical Association, 20(152):549–555, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277171.

Gries:1924:BRB

[Gri24] John M. Gries. Book review: [untitled]. Journal of the American Statistical Association, 19(147):423–424, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277392.

Grove:1923:ALT

[Gro23] Charles C. Grove. Abridged life tables for health officers. Journal of the American Statistical Association, 18(144):1028–1038, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965664.

Guilfoy:1926:PFD

[Gui26] William H. Guilfoy. The past and future development of vital statistics in the United States: III, Cressy L. Wilbur. Journal of the American Statistical Association, 21(155):272–274, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277053.

Gumbel:1926:SCS

[Gum26] E. J. Gumbel. Spurious correlation and its significance to physiology. Journal of the American Statistical Association, 21(154):179– 194, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277144.

H:1923:PWCa

[H.23a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 18(142):788–791, June 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276719.

H:1923:PWCb

[H.23b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 18(143):915–918, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965477.

H:1923:PWCc

[H.23c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 18(144):1040–1042, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965666.

H:1924:PWCd

[H.24a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 19(148):524–525, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281572.

H:1924:PWCb

[H.24b] J. A. H. Progress of work in the Census Bureau: Census of institutional classes. Journal of the American Statistical Association, 19(146):231–232, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277236.

H:1924:PWCa

[H.24c] J. A. H. Progress of work in the Census Bureau: Fourteenth census publications. Journal of the American Statistical Association, 19(145):82–84, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277265.

H:1924:PWCc

[H.24d] J. A. H. Progress of work in the Census Bureau: Quinquennial census of agriculture. Journal of the American Statistical Association, 19(147):396–399, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277381.

H:1924:ESC

[H.24e] R. H. Employment statistics for cities. Journal of the American Statistical Association, 19(145):84–85, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277266.

H:1925:NCB

[H.25a] J. A. H. Notes from the Census Bureau. Journal of the American Statistical Association, 20(151):423–425, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965530.

H:1925:PWCa

[H.25b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 20(149):120–122, March 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277443.

H:1925:PWCb

[H.25c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 20(150):261–263, June 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277125.

H:1925:PWCc

[H.25d] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 20(152):559–561, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277174.

H:1926:PWCa

[H.26a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 21(153):66–68, March 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965503.

H:1926:PWCb

[H.26b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 21(154):218–220, June 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277149.

H:1926:PWCc

[H.26c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 21(156):479–480, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276985.

H:1927:PWCa

[H.27a] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 22(157):81–82, March 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277353.

H:1927:PWCb

[H.27b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 22(158):230–231, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276782.

H:1927:PWCc

[H.27c] J. A. H. Progress of work in the Census Bureau: Defining enumeration districts. Journal of the American Statistical Association, 22(160):510–513, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276836.

H:1928:LFC

[H.28a] J. A. H. Legislation for the fifteenth census. Journal of the American Statistical Association, 23(163):321–323, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276747.

H:1928:PWCa

[H.28b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 23(161):76–77, March 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277410.

H:1928:PWCb

[H.28c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 23(162):188–190, June 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276896.

H:1928:PWCc

[H.28d] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 23(164):442–444, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277293.

H:1929:PN

[H.29a] J. A. H. Personal notes. Journal of the American Statistical Association, 24(165):80–82, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277325.

H:1929:PWCa

[H.29b] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 24(165):74–75, March 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277323.

H:1929:PWCb

[H.29c] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 24(166):189–191, June 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276693.

H:1929:PWCc

[H.29d] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 24(167):311–314, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276863.

H:1929:PWCd

[H.29e] J. A. H. Progress of work in the Census Bureau. Journal of the American Statistical Association, 24(168):427–431, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277090.

Halsey:1923:BRB

[Hal23] Olga S. Halsey. Book review: [untitled]. Journal of the American Statistical Association, 18(141):686–688, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277517.

Hall:1924:SVR

[Hal24] Lincoln W. Hall. Seasonal variation as a relative of secular trend. Journal of the American Statistical Association, 19(146):156–166, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277227.

Hall:1925:MST

[Hal25] Lincoln W. Hall. A moving secular trend and moving integration. Journal of the American Statistical Association, 20(149):13–24, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277430.

Hall:1926:DPP

[Hal26] Lincoln W. Hall. The determination of past and present secular trends. Journal of the American Statistical Association, 21(154):206– 212, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277146.

Hansen:1922:BRB

[Han22a] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 18(140):538–540, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276935.

Hansen:1922:BPL

[Han22b] Alvin H. Hansen. The buying power of labor during the war. Journal of the American Statistical Association, 18(137):56–66, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277468.

Hansen:1922:ICU

[Han22c] Alvin H. Hansen. Industrial classes in the United States in 1920. Journal of the American Statistical Association, 18(140):503–506, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276923.

Hankins:1923:BRB

[Han23] F. H. Hankins. Book review: [untitled]. Journal of the American Statistical Association, 18(143):932–936, September 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965484.

Hansen:1924:WP

[Han24] Alvin H. Hansen. Wholesale prices, 1801–1840. Journal of the American Statistical Association, 19(147):377–381, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277376.

Hankins:1925:BRB

[Han25] F. H. Hankins. Book review: [untitled]. Journal of the American Statistical Association, 20(150):288–290, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277134.

Hankins:1926:BRB

[Han26] F. H. Hankins. Book review: [untitled]. Journal of the American Statistical Association, 21(155):365–366, September 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277073.

Hansen:1927:BRB

[Han27a] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 22(158):261–262, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276795.

Hansen:1927:KIH

[Han27b] Alvin H. Hansen. Karsten’s interpretation of the Harvard Business Indexes. Journal of the American Statistical Association, 22 (159):367–369, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276806.

Hansen:1928:BRB

[Han28] Alvin H. Hansen. Book review: [untitled]. Journal of the American Statistical Association, 23(162):213–214, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276905.

Haney:1930:CAD

[Han30] Lewis H. Haney. Corporation accounting data from the standpoint of the investor. Journal of the American Statistical Association, 25(169): 7–12, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277189.

Harris:1922:CBV

[Har22a] J. Arthur Harris. On the correlation between a variable and the deviation of an associated but not dependent variable from its probable value. Journal of the American Statistical Association, 18 (139):393–394, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276959.

Harris:1922:PED

[Har22b] J. Arthur Harris. The probable error of the difference between two differences. Journal of the American Statistical Association, 18(140):514, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276925.

Hart:1922:PWU

[Har22c] Hornell Hart. Pre-War unemployment statistics: a critique. Journal of the American Statistical Association, 18(139):385–391, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276957.

Hart:1922:MMM

[Har22d] William L. Hart. The method of monthly means for determination of a seasonal variation. Journal of the American Statistical Association, 18(139):341–349, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276952.

Harris:1923:BRB

[Har23a] Henry J. Harris. Book review: [untitled]. Journal of the American Statistical Association, 18(143):931–932, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965483.

Hart:1923:BRB

[Har23b] Hornell Hart. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1064–1067, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965675.

Hart:1923:MAA

[Har23c] Hornell Hart. Mongolian amentia and the ages of the mothers at the births of the aments. Journal of the American Statistical Association, 18(143):900–903, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965472.

Harmon:1924:DRP

[Har24a] G. E. Harmon. Death rates from puerperal septicemia and their variability. Journal of the American Statistical Association, 19(146):221– 226, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277234.

Hart:1924:SMD

[Har24b] Hornell Hart. A simplified method of determining what contrasts between two percentage distributions are significant. Journal of the American Statistical Association, 19(145):65–70, March 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277260.

Hart:1925:FDF

[Har25] Hornell Hart. Familial differential fecundity. Journal of the American Statistical Association, 20(149):25–30, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277431.

Hart:1926:RP

[Har26] Hornell Hart. The reliability of a percentage. Journal of the American Statistical Association, 21(153):40–46, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965496.

Hardy:1927:BRBa

[Har27a] Charles O. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 22(157):120–122, March 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277360.

Hardy:1927:BRBb

[Har27b] Charles O. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 22(160):536–537, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276849.

Harris:1927:CMA

[Har27c] J. Arthur Harris. The calculation of moments about the mean as origin. Journal of the American Statistical Association, 22(157):75– 78, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277351.

Harrison:1928:BRB

[Har28] Shelby M. Harrison. Book review: [untitled]. Journal of the American Statistical Association, 23(164):472–473, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277308.

Hardy:1929:BRB

[Har29] Charles O. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 24(166):227–229, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276706.

Hawtrey:1929:BRB

[Haw29] R. G. Hawtrey. Book review: [untitled]. Journal of the American Statistical Association, 24(166):216–219, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276700.

Heiberg:1926:VDB

[Hei26] Poul Heiberg. The variability in the difference between the mortality of men and women in Copenhagen during the last twenty years. Journal of the American Statistical Association, 21(154):213–215, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277147.

Henderson:1922:BRB

[Hen22] Robert Henderson. Book review: [untitled]. Journal of the American Statistical Association, 18(140):548–552, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276939.

Henderson:1928:BRB

[Hen28] Robert Henderson. Book review: [untitled]. Journal of the American Statistical Association, 23(161):109, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277426.

Herwitz:1922:BRB

[Her22a] H. K. Herwitz. Book review: [untitled]. Journal of the American Statistical Association, 18(139):404–407, September 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276964.

Herwitz:1922:R

[Her22b] H. K. Herwitz. A reply. Journal of the American Statistical Association, 18(140):528, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276931.

Herwitz:1924:BRB

[Her24] H. K. Herwitz. Book review: [untitled]. Journal of the American Statistical Association, 19(146):253–257, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277244.

Herskovits:1925:BRB

[Her25a] Melville J. Herskovits. Book review: [untitled]. Journal of the American Statistical Association, 20(152):588–590, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277183.

Herskovits:1925:FDV

[Her25b] Melville J. Herskovits. A further discussion of the variability of family strains in the negro–white population of New York city. Journal of the American Statistical Association, 20(151):380–389, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965521.

Herwitz:1925:BRB

[Her25c] H. K. Herwitz. Book review: [untitled]. Journal of the American Statistical Association, 20(152):576–578, December 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277178.

Herskovits:1927:BRB

[Her27] Melville J. Herskovits. Book review: [untitled]. Journal of the American Statistical Association, 22(160):542–544, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276851.

Herskovits:1928:BRB

[Her28a] Melville J. Herskovits. Book review: [untitled]. Journal of the American Statistical Association, 23(163):352–354, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276760.

Herwitz:1928:BRB

[Her28b] H. K. Herwitz. Book review: [untitled]. Journal of the American Statistical Association, 23(161):108–109, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277425.

Hewes:1926:BRB

[Hew26] Amy Hewes. Book review: [untitled]. Journal of the American Statistical Association, 21(155):367–369, September 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277075.

Hewes:1928:BRB

[Hew28] Amy Hewes. Book review: [untitled]. Journal of the American Statistical Association, 23(163):350–351, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276758.

Hickernell:1926:BFN

[Hic26] Warren F. Hickernell. A basis of forecasting national income. Journal of the American Statistical Association, 21(155):350–353, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277066.

Hill:1922:BRB

[Hil22a] J. A. Hill. Book review: [untitled]. Journal of the American Statistical Association, 18(139):417–418, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276971.

Hill:1922:SRP

[Hil22b] Joseph A. Hill. Some results of the 1920 population census. Journal of the American Statistical Association, 18(139):350–358, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276953.

Hill:1923:C

[Hil23] Joseph A. Hill. Correspondence. Journal of the American Statistical Association, 18(141):668–669, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277511.

Hiscock:1925:BRB

[His25] I. V. Hiscock. Book review: [untitled]. Journal of the American Statistical Association, 20(149):149–150, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277454.

H:1923:MN

[HMR23] J. A. H., L. M., and W. S. R. Miscellaneous notes. Journal of the American Statistical Association, 18(141):669–676, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277512.

Harris:1928:PES

[HN28] J. Arthur Harris and Marie M. Ness. The probable error of the standard deviation. Journal of the American Statistical Association, 23(162):178–179, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276891.

Hoffman:1922:BRBa

[Hof22a] F. L. Hoffman. Book review: [untitled]. Journal of the American Statistical Association, 18(140):545–548, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276938.

Hoffman:1922:BRBb

[Hof22b] F. L. Hoffman. Book review: [untitled]. Journal of the American Statistical Association, 18(140):555–556, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276942.

Hoffman:1922:MN

[Hof22c] F. L. Hoffman. Miscellaneous notes. Journal of the American Statistical Association, 18(137):121–125, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277481.

Hoffman:1922:RCC

[Hof22d] F. L. Hoffman. The results of the census of Czecho–Slovakia for 1921. Journal of the American Statistical Association, 18(137):109– 110, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277478.

Hoffman:1924:WAU

[Hof24] F. L. Hoffman. Welfare activities of the United States steel corporation. Journal of the American Statistical Association, 19 (148):522–524, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281571.

Hogg:1927:DMD

[Hog27] Margaret H. Hogg. The determination of the mean deviation for graded data. Journal of the American Statistical Association, 22 (160):501–504, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276833.

Hollander:1922:BRB

[Hol22] Jacob H. Hollander. Book review: [untitled]. Journal of the American Statistical Association, 18(139):411, September 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276967.

Holt:1923:SSS

[Hol23a] William L. Holt. A statistical study of smokers and non-smokers at the University of Tennessee. Journal of the American Statistical Association, 18(142):766–772, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276714.

Holzinger:1923:CFC

[Hol23b] Karl J. Holzinger. A combination form for calculating the correlation coefficient and ratios. Journal of the American Statistical Association, 18(141):623–627, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277503.

Holzinger:1923:HDC

[Hol23c] Karl J. Holzinger. Higher degrees of college professors: a study of some factors affecting the research of graduate students who become college professors. Journal of the American Statistical Association, 18(143):878–888, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965470.

Hollander:1928:BRB

[Hol28] Jacob H. Hollander. Book review: [untitled]. Journal of the American Statistical Association, 23(161):104–106, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277422.

Holden:1930:BS

[Hol30] Thomas S. Holden. The building situation. Journal of the American Statistical Association, 25(169):64–68, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277199.

Hotelling:1925:GMT

[Hot25] Harold Hotelling. A general mathematical theory of depreciation. Journal of the American Statistical Association, 20(151):340–353, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2965518.

Hotelling:1926:BRBa

[Hot26a] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 21(155):360–363, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277071.

Hotelling:1926:BRBb

[Hot26b] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 21(156):503–505, December 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276993.

Hotelling:1927:BRR

[Hot27a] Harold Hotelling. Book review: R. A. Fisher, Statistical Methods for Research Workers. Journal of the American Statistical Association, 22(159):411–412, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276824.

Hotelling:1927:DES

[Hot27b] Harold Hotelling. Differential equations subject to error, and population estimates. Journal of the American Statistical Association, 22(159):283–314, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276800.

Hotelling:1928:BRB

[Hot28] Harold Hotelling. Book review: [untitled]. Journal of the American Statistical Association, 23(163):346, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276756.

Howland:1922:SSP

[How22] Katherine E. Howland. A statistical study of poor relief in Massachusetts. Journal of the American Statistical Association, 18 (140):480–489, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276920.

Hoyt:1923:IDW

[Hoy23] Homer Hoyt. The inequality in the distribution of wealth and income in the United States. Journal of the American Statistical Association, 18(141):650–651, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277505.

Harris:1927:LAC

[HT27] J. Arthur Harris and Alan E. Treloar. On a limitation in the applicability of the contingency coefficient. Journal of the American Statistical Association, 22(160):460–472, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276828.

Harris:1929:SCL

[HT29] J. Arthur Harris and Chi Tu. A second category of limitations in the applicability of the contingency coefficient. Journal of the American Statistical Association, 24(168):367–375, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277082.

Hughes:1930:SP

[Hug30] James F. Hughes. Security prices. Journal of the American Statistical Association, 25(169):76–79, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277202.

Hull:1925:ACC

[Hul25] Clark L. Hull. An automatic correlation calculating machine. Journal of the American Statistical Association, 20(152):522–531, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277166.

Hunt:1922:UBC

[Hun22] Edward Eyre Hunt. Unemployment and the business cycle. Journal of the American Statistical Association, 18(137):110–112, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277479.

Hurlin:1922:ESU

[Hur22] Ralph G. Hurlin. The employment statistics of the United States employment service. Journal of the American Statistical Association, 18(140):490–497, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276921.

Hurlin:1923:CES

[Hur23a] Ralph G. Hurlin. Canadian employment statistics. Journal of the American Statistical Association, 18(141):615–622, March 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277502.

Hurlin:1923:RIA

[Hur23b] Ralph G. Hurlin. A record of industrial adjustment to the last depression. Journal of the American Statistical Association, 18(142):785– 788, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276718.

Hurlin:1927:BRB

[Hur27] Ralph G. Hurlin. Book review: [untitled]. Journal of the American Statistical Association, 22(157):122–123, March 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277361.

Hurlin:1929:BRB

[Hur29] Ralph G. Hurlin. Book review: [untitled]. Journal of the American Statistical Association, 24(168):442–443, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277095.

Ingraham:1923:PHM

[Ing23] Mark H. Ingraham. On Professor H. L. Moore’s mathematical analysis of the business cycle. Journal of the American Statistical Association, 18(142):759–765, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276713.

Ingraham:1925:RTS

[Ing25] Olin Ingraham. The refinement of time series. Journal of the American Statistical Association, 20(150):231–237, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277117.

J:1928:CCP

[J.28] K. W. J. The cotton crop and its price. Journal of the American Statistical Association, 23(162):181–184, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276893.

Jenkins:1928:AFC

[Jen28] Thomas N. Jenkins. Apparatus to facilitate the calculation of the moments of a distribution. Journal of the American Statistical Association, 23(161):58–60, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277405.

Jerome:1926:BRB

[Jer26] Harry Jerome. Book review: [untitled]. Journal of the American Statistical Association, 21(154):249–250, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277160.

Jerome:1929:BRB

[Jer29] Harry Jerome. Book review: [untitled]. Journal of the American Statistical Association, 24(167):343–344, September 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276877.

Johnston:1929:NIS

[Joh29] Mary Johnston. The need for an index for social data. Journal of the American Statistical Association, 24(168):398–404, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277085.

Jones:1926:BRB

[Jon26] Grosvenor Jones. Book review: [untitled]. Journal of the American Statistical Association, 21(155):370–371, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277076.

Jordan:1925:CLI

[Jor25] Virgil Jordan. The cost of living indices of the N.I.C.B. and of the B. of L. S. Journal of the American Statistical Association, 20(150):249– 253, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277121.

Johnson:1930:CAI

[JS30] Norris O. Johnson and John T. Shirer. Certain aspects of the interpretation of price-earnings ratios. Journal of the American Statistical Association, 25(169):106–110, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277209.

Joy:1927:AIE

[JT27] Aryness Joy and Woodlief Thomas. Adjustments for the influence of Easter in department stores sales. Journal of the American Statistical Association, 22(160):493–496, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276831.

Joy:1928:UMA

[JT28] Aryness Joy and Woodlief Thomas. The use of moving averages in the measurement of seasonal variations. Journal of the American Statistical Association, 23(163):241–252, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276735.

K:1924:SAV

[K.24] E. W. K. Standards for American vital statisticians. Journal of the American Statistical Association, 19(145):81–82, March 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277264.

K:1928:NMM

[K.28a] W. I. K. New methods of measuring public health progress. Journal of the American Statistical Association, 23(161):72–76, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277409.

K:1928:PTP

[K.28b] W. I. K. Price trends and prosperity. Journal of the American Statistical Association, 23(161):68–72, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277408.

K:1929:IRS

[K.29] W. I. K. Interest rates and security prices. Journal of the American Statistical Association, 24(165):69–72, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277321.

Karsten:1924:TQE

[Kar24] Karl G. Karsten. The theory of quadrature in economics. Journal of the American Statistical Association, 19(145):14–29, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277255.

Karsten:1925:BRB

[Kar25] Karl G. Karsten. Book review: [untitled]. Journal of the American Statistical Association, 20(150):284–285, June 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277131.

Karsten:1926:HBI

[Kar26] Karl G. Karsten. The Harvard Business Indexes — a new interpretation. Journal of the American Statistical Association, 21 (156):399–418, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276977.

Kopf:1923:PEF

[KC23] Edwin W. Kopf and Robert E. Chaddock. Proceedings of the Eighty-Fourth Annual Meeting of the American Statistical Association. Journal of the American Statistical Association, 18(141):659–667, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277510.

Kelley:1922:BRB

[Kel22] Truman L. Kelley. Book review: [untitled]. Journal of the American Statistical Association, 18(137):132–134, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277484.

Kelley:1925:MCD

[Kel25] Truman L. Kelley. Measures of correlation determined from groups of varying homogeneity. Journal of the American Statistical Association, 20(152):512–521, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277165.

King:1922:BRBa

[Kin22a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 18(137):148–149, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277494.

King:1922:BRBb

[Kin22b] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 18(139):419–420, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276974.

King:1922:NVSa

[Kin22c] Willford I. King. The net volume of saving in the United States. Journal of the American Statistical Association, 18(139):305–323, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276950.

King:1922:NVSb

[Kin22d] Willford I. King. The net volume of saving in the United States. Journal of the American Statistical Association, 18(140):455–470, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276918.

King:1924:IMM

[Kin24a] W. Willford I. King. An improved method for measuring the seasonal factor. Journal of the American Statistical Association, 19 (147):301–313, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277370.

King:1924:PGV

[Kin24b] Willford I. King. Population growth and the volume of building. Journal of the American Statistical Association, 19(145):9–13, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277254.

King:1924:PUI

[Kin24c] Willford I. King. Principles underlying the isolation of cycles and trends. Journal of the American Statistical Association, 19 (148):468–475, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281563.

King:1925:BRBa

[Kin25a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 20(149):152–153, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277456.

King:1925:BRBb

[Kin25b] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 20(152):584–586, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277181.

King:1925:PES

[Kin25c] Willford I. King. Proceedings of the Eighty-Sixth Annual Meeting of the American Statistical Association. Journal of the American Statistical Association, 20(149):109–119, March 1925. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277442.

King:1927:BRB

[Kin27a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 22(159):408–409, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276822.

King:1927:CMA

[Kin27b] Willford I. King. Classification of members of American Statistical Association on basis of duties and interests. Journal of the American Statistical Association, 22(158):224–226, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276777.

King:1927:WDC

[Kin27c] Willford I. King. Wealth distribution in the continental United States at the close of 1921. Journal of the American Statistical Association, 22(158):135–153, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276768.

King:1928:RFP

[Kin28] Willford I. King. The relationship in a frequency polygon of the average of any class to the mid-point of the corresponding class-interval. Journal of the American Statistical Association, 23(164):435–439, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277289.

King:1929:BRB

[Kin29a] Willford I. King. Book review: [untitled]. Journal of the American Statistical Association, 24(166):220–221, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276701.

King:1929:CCS

[Kin29b] Willford I. King. Central collection of social statistics. Journal of the American Statistical Association, 24(166):178–181, June 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276690.

King:1929:EO

[Kin29c] Willford I. King. Economic outlook for 1929. Journal of the American Statistical Association, 24(166):182–187, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276691.

King:1929:FNT

[Kin29d] Willford I. King. Financing the Nation’s trade and industry. Journal of the American Statistical Association, 24(167):307–311, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276862.

King:1930:HEN

[Kin30] Willford I. King. A hasty estimate of the national income in 1929. Journal of the American Statistical Association, 25(169):73–75, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277201.

Killough:1926:PMF

[KK26] Hugh B. Killough and Lucy W. Killough. Price making forces in cotton markets. Journal of the American Statistical Association, 21(153):47–54, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965497.

Kelley:1929:DVK

[KM29] Truman L. Kelley and Quinn McNemar. Doolittle versus the Kelley– Salisbury iteration method for computing multiple regression coefficients. Journal of the American Statistical Association, 24(166):164– 169, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276687.

Knauth:1922:PCS

[Kna22] Oswald W. Knauth. The place of corporate surplus in the national income. Journal of the American Statistical Association, 18(138):157– 166, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277523.

Knauth:1923:BRB

[Kna23] O. W. Knauth. Book review: [untitled]. Journal of the American Statistical Association, 18(141):694, March 1923. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277521.

Knibbs:1924:NUPa

[Kni24a] George H. Knibbs. The nature of an unequivocal price-index and quantity-index. Journal of the American Statistical Association, 19(145):42–60, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277258.

Knibbs:1924:NUPb

[Kni24b] George H. Knibbs. The nature of an unequivocal price-index and quantity-index: Part II. Journal of the American Statistical Association, 19(146):196–205, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277231.

Knibbs:1926:LGP

[Kni26] George H. Knibbs. The laws of growth of a population. Journal of the American Statistical Association, 21(156):381–398, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276976.

Knibbs:1927:LGP

[Kni27] George H. Knibbs. The laws of growth of a population. Journal of the American Statistical Association, 22(157):49–59, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277348.

Knight:1929:BRB

[Kni29] Frank H. Knight. Book review: [untitled]. Journal of the American Statistical Association, 24(167):345–346, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276880.

Kopf:1923:BRB

[Kop23] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 18(142):828, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276732.

Kopf:1924:BRBa

[Kop24a] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 19(145):114–116, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277274.

Kopf:1924:BRBb

[Kop24b] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 19(146):266–268, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277250.

Kopf:1924:BRBc

[Kop24c] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 19(148):554–555, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281581.

Kopf:1925:BRB

[Kop25] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 20(151):452–456, September 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965538.

Kopf:1927:BRBa

[Kop27a] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 22(157):123–125, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277362.

Kopf:1927:BRBb

[Kop27b] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 22(158):263–264, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276797.

Kopf:1928:BRB

[Kop28] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 23(162):228–230, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276912.

Kopf:1929:BRB

[Kop29] E. W. Kopf. Book review: [untitled]. Journal of the American Statistical Association, 24(168):452–453, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277104.

Kelley:1926:IMD

[KS26] Truman L. Kelley and Frank S. Salisbury. An iteration method for determining multiple correlation constants. Journal of the American Statistical Association, 21(155):282–292, September 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277056.

Kuczynski:1929:BRB

[Kuc29] Jurgen Kuczynski. Book review: [untitled]. Journal of the American Statistical Association, 24(165):106, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277340.

Kuznets:1928:BRB

[Kuz28a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 23(162):218–220, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276908.

Kuznets:1928:MCT

[Kuz28b] Simon Kuznets. On moving correlation of time sequences. Journal of the American Statistical Association, 23(162):121–136, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276885.

Kuznets:1928:ATS

[Kuz28c] Simon Kuznets. On the analysis of time series. Journal of the American Statistical Association, 23(164):398–410, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277285.

Kuznets:1929:BRB

[Kuz29a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 24(168):445–446, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277098.

Kuznets:1929:REC

[Kuz29b] Simon Kuznets. Random events and cyclical oscillations. Journal of the American Statistical Association, 24(167):258–275, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276856.

Lamson:1922:MBC

[Lam22] F. L. Lamson. Management and the business cycle: Discussion. Journal of the American Statistical Association, 18(137):31–32, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277464.

Lane:1924:TYP

[Lan24] Mortimer B. Lane. Three years’ progress in current business statistics. Journal of the American Statistical Association, 19(148):508–515, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281566.

Lescohier:1923:BRB

[Les23a] D. D. Lescohier. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1069–1071, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965677.

Lescohier:1923:GAS

[Les23b] Don D. Lescohier. Geographical aspects of statistics of employment and earnings. Journal of the American Statistical Association, 18(142):772–776, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276715.

Levin:1928:SSE

[Lev28] Nathan P. Levin. A statistical study of the economic and sociological significance of population age. Journal of the American Statistical Association, 23(161):40–48, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277402.

Leven:1929:BRB

[Lev29] Maurice Leven. Book review: [untitled]. Journal of the American Statistical Association, 24(167):337–338, September 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276873.

Lincoln:1924:BRB

[Lin24] Edmond E. Lincoln. Book review: [untitled]. Journal of the American Statistical Association, 19(145):129–130, March 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277281.

Lincoln:1929:BRB

[Lin29] Edmond E. Lincoln. Book review: [untitled]. Journal of the American Statistical Association, 24(168):444–445, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277097.

Lorenz:1924:BRB

[Lor24] M. O. Lorenz. Book review: [untitled]. Journal of the American Statistical Association, 19(147):412–413, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277386.

Lorenz:1926:CFF

[Lor26a] M. O. Lorenz. Comment on “A Formula for Freight Rates”. Journal of the American Statistical Association, 21(153):66, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965502.

Lorenz:1926:SAJ

[Lor26b] M. O. Lorenz. The statistical allocation of joint costs: Discussion. Journal of the American Statistical Association, 21(153):25–26, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2965494.

Lotka:1926:BRB

[Lot26] Alfred J. Lotka. Book review: [untitled]. Journal of the American Statistical Association, 21(155):374, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277079.

Lotka:1927:SAF

[Lot27] Alfred J. Lotka. The size of American families in the eighteenth century: And the significance of the empirical constants in the pearlreed law of population growth. Journal of the American Statistical Association, 22(158):154–170, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276769.

Lotka:1929:BRBa

[Lot29a] Alfred J. Lotka. Book review: [untitled]. Journal of the American Statistical Association, 24(165):103–104, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277337.

Lotka:1929:BRBb

[Lot29b] Alfred J. Lotka. Book review: [untitled]. Journal of the American Statistical Association, 24(167):332–333, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276869.

Lovitt:1928:INB

[Lov28] W. V. Lovitt. Index number bias. Journal of the American Statistical Association, 23(161):10–17, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277398.

Lubin:1922:BRB

[Lub22] Isador Lubin. Book review: [untitled]. Journal of the American Statistical Association, 18(137):138–140, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277487.

Lubin:1924:BRB

[Lub24] Isador Lubin. Book review: [untitled]. Journal of the American Statistical Association, 19(148):547–550, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281578.

Lubin:1927:BRB

[Lub27] Isador Lubin. Book review: [untitled]. Journal of the American Statistical Association, 22(160):544–545, December 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276852.

Lundberg:1926:DP

[Lun26] George A. Lundberg. The disappearing “public”. Journal of the American Statistical Association, 21(156):476–477, December 1926.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276983.

Lurguin:1924:FAQ

[Lur24] Constant Lurguin. The fiftieth anniversary of Quetelet’s death. Journal of the American Statistical Association, 19(148):520–522, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281570.

Lyon:1928:BRB

[Lyo28] Leverett S. Lyon. Book review: [untitled]. Journal of the American Statistical Association, 23(162):208–211, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276903.

Macaulay:1926:CIN

[Mac26] Fred R. Macaulay. The construction of an index number of bond yields in the United States, 1859 to 1926. Journal of the American Statistical Association, 21(153):27–39, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965495.

Mann:1922:CCV

[Man22a] Lawrence B. Mann. Comparison of changes in the value and the volume of foreign trade. Journal of the American Statistical Association, 18(137):99–101, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277473.

Mann:1922:STD

[Man22b] Lawrence B. Mann. Sectional trends in department store trade, 1919– 22. Journal of the American Statistical Association, 18(138):255–258, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277533.

Marshall:1927:DIN

[Mar27] Herbert Marshall. Differentiation in index numbers of security prices. Journal of the American Statistical Association, 22(157):60– 65, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277349.

May:1929:BRB

[May29] George O. May. Book review: [untitled]. Journal of the American Statistical Association, 24(168):443–444, December 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277096.

Miller:1927:PIN

[MC27] E. M. Miller and F. J. Cavanaugh. Price index of the National Bank of Commerce in New York. Journal of the American Statistical Association, 22(158):171–183, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276770.

McChesney:1928:GCM

[McC28] Ruth McChesney. The graphical construction of moving averages. Journal of the American Statistical Association, 23(162):164–172, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276889.

McFall:1924:HCF

[McF24] R. J. McFall. The higher cost of food in Massachusetts. Journal of the American Statistical Association, 19(147):362–377, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277375.

McFall:1930:CD

[McF30] Robert J. McFall. The census of distribution. Journal of the American Statistical Association, 25(169):125–129, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277214.

McMillen:1929:RAS

[McM29] A. W. McMillen. A registration area for social statistics. Journal of the American Statistical Association, 24(165):68–69, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277320.

Melville:1923:DSR

[Mel23] Andrew H. Melville. Department store research. Journal of the American Statistical Association, 18(143):918–919, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965478.

Merriam:1924:RL

[Mer24] Lewis Merriam. Recent literature. Journal of the American Statistical Association, 19(147):426–430, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277394.

Merriam:1926:RPS

[Mer26] Charles E. Merriam. Report of the president of the social science research council for the year 1925. Journal of the American Statistical Association, 21(153):68–72, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2965504.

Meriam:1929:BRB

[Mer29] R. S. Meriam. Book review: [untitled]. Journal of the American Statistical Association, 24(165):104–105, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277338.

M:1929:PN

[MH29] W. C. M. and J. A. H. Personal notes. Journal of the American Statistical Association, 24(166):200–203, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276695.

Mills:1922:BRB

[Mil22] Frederick C. Mills. Book review: [untitled]. Journal of the American Statistical Association, 18(140):552–554, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276940.

Mills:1924:MCP

[Mil24] Frederick C. Mills. The measurement of correlation and the problem of estimation. Journal of the American Statistical Association, 19(147):273–300, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277369.

Mills:1925:BRB

[Mil25] Frederick C. Mills. Book review: [untitled]. Journal of the American Statistical Association, 20(152):586–588, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277182.

Millis:1926:BRB

[Mil26a] H. A. Millis. Book review: [untitled]. Journal of the American Statistical Association, 21(156):508–509, December 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276995.

Mills:1926:HCD

[Mil26b] Frederick C. Mills. An hypothesis concerning the duration of business cycles. Journal of the American Statistical Association, 21 (156):447–457, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276979.

Mills:1929:BRB

[Mil29] Frederick C. Mills. Book review: [untitled]. Journal of the American Statistical Association, 24(168):453–454, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277105.

Miner:1923:BRB

[Min23] John Rice Miner. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1073–1074, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965679.

Miner:1926:BRB

[Min26] John Rice Miner. Book review: [untitled]. Journal of the American Statistical Association, 21(153):101–102, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965510.

Mitchell:1922:BRB

[Mit22] Wesley C. Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 18(138):281–283, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277542.

Mitchell:1923:BRB

[Mit23] Wesley C. Mitchell. Book review: [untitled]. Journal of the American Statistical Association, 18(142):820–822, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276730.

Mitchell:1928:ASS

[Mit28] Wesley C. Mitchell. Activities of the social science research council. Journal of the American Statistical Association, 23(162):192–193, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276898.

Monk:1928:LCP

[MJ28] Ardis T. Monk and Helen R. Jeter. The logistic curve and the prediction of the population of the Chicago region. Journal of the American Statistical Association, 23(164):361–385, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277283.

Moulton:1925:PYR

[MM25] H. G. Moulton and C. E. McGuire. Professor Young’s review of “Germany’s Capacity to Pay”. Journal of the American Statistical Association, 20(150):256–260, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277123.

Moore:1922:EDF

[Moo22] Henry Ludwell Moore. Elasticity of demand and flexibility of prices. Journal of the American Statistical Association, 18(137):8–19, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277462.

Moritz:1923:TIF

[Mor23a] Robert E. Moritz. On Taylor’s interpolation formula as a limiting case of the interpolation formula of Lagrange. Journal of the American Statistical Association, 18(142):781–784, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276717.

Morley:1923:BRB

[Mor23b] Felix Morley. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1068–1069, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965676.

Moritz:1928:EPS

[Mor28] Robert E. Moritz. An elementary proof of Stirling’s formula. Journal of the American Statistical Association, 23(161):55–57, March 1928.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277404.

Moulton:1927:RWV

[MP27] M. Estelle Moulton and Raymond E. Pannier. Ratio of wages to value added by manufacturing processes in massachussetts and the United States. Journal of the American Statistical Association, 22 (159):359–366, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276805.

Mudgett:1922:BRB

[Mud22] Bruce D. Mudgett. Book review: [untitled]. Journal of the American Statistical Association, 18(137):146–147, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277492.

Mudgett:1923:BRB

[Mud23] Bruce D. Mudgett. Book review: [untitled]. Journal of the American Statistical Association, 18(143):941–942, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965488.

Musselman:1923:SCA

[Mus23] John Rogers Musselman. Spurious correlation applied to urn schemata. Journal of the American Statistical Association, 18 (143):908–911, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965474.

Musselman:1928:BRB

[Mus28] J. R. Musselman. Book review: [untitled]. Journal of the American Statistical Association, 23(164):457–459, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277300.

Myers:1922:MPP

[Mye22] Margaret G. Myers. Monthly production of pig iron, 1884 to 1903. Journal of the American Statistical Association, 18(138):247–249, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277531.

Myers:1929:BRB

[Mye29] Margaret G. Myers. Book review: [untitled]. Journal of the American Statistical Association, 24(167):341–343, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276876.

Nadler:1930:ETY

[Nad30] Marcus Nadler. Europe ten years after the War. Journal of the American Statistical Association, 25(169):170–174, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277223.

Neale:1925:NZS

[Nea25] Edward P. Neale. A New Zealand study in infant mortality. Journal of the American Statistical Association, 20(151):354–367, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965519.

Neifeld:1927:SSC

[Nei27] M. R. Neifeld. A study of spurious correlation. Journal of the American Statistical Association, 22(159):331–338, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276802.

Newsholme:1928:BRB

[New28] Arthur Newsholme. Book review: [untitled]. Journal of the American Statistical Association, 23(164):452–453, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277296.

Norton:1928:AVY

[Nor28] L. J. Norton. Average variations in yields, acreage and production. Journal of the American Statistical Association, 23(161):61–63, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277406.

Notestein:1925:MN

[Not25a] F. W. Notestein. Miscellaneous notes. Journal of the American Statistical Association, 20(149):122–127, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277444.

Notestein:1925:BRB

[Not25b] Frank W. Notestein. Book review: [untitled]. Journal of the American Statistical Association, 20(151):449–452, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965537.

Neale:1924:SIF

[NS24] E. P. Neale and D. M. Y. Sommerville. A shortened interpolation formula for certain types of data. Journal of the American Statistical Association, 19(148):515–517, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281567.

Ogburn:1922:BRB

[Ogb22] William F. Ogburn. Book review: [untitled]. Journal of the American Statistical Association, 18(138):284–286, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277544.

Ogburn:1923:BRB

[Ogb23] William F. Ogburn. Book review: [untitled]. Journal of the American Statistical Association, 18(142):808–809, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276726.

Ogburn:1928:BRB

[Ogb28] William F. Ogburn. Book review: [untitled]. Journal of the American Statistical Association, 23(164):474–475, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277310.

Orchard:1926:BRB

[Orc26] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 21(154):239–242, June 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277155.

Orchard:1928:BRB

[Orc28] Dorothy J. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 23(162):224–227, June 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276910.

Orchard:1929:BRB

[Orc29] John E. Orchard. Book review: [untitled]. Journal of the American Statistical Association, 24(167):334–335, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276870.

Osborn:1923:FOS

[Osb23] C. C. Osborn. The future oil supply. Journal of the American Statistical Association, 18(144):1004–1009, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965661.

Ogburn:1922:IBC

[OT22] William F. Ogburn and Dorothy S. Thomas. The influence of the business cycle on certain social conditions. Journal of the American Statistical Association, 18(139):324–340, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276951.

Page:1929:ARF

[Pag29] A. J. Page. On the annual revision of forecasting formulas based on partial regression equations. Journal of the American Statistical Association, 24(166):123–126, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276683.

Palmer:1929:PMH

[Pal29] Gladys L. Palmer. The physical measurements of hollins freshmen, 1920–1927. Journal of the American Statistical Association, 24(165): 40–49, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277316.

Parry:1925:BRB

[Par25] Carl E. Parry. Book review: [untitled]. Journal of the American Statistical Association, 20(152):578–579, December 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277179.

Parmelee:1926:BRB

[Par26] Julius H. Parmelee. Book review: [untitled]. Journal of the American Statistical Association, 21(154):242–243, June 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277156.

Parry:1929:BRB

[Par29] Carl E. Parry. Book review: [untitled]. Journal of the American Statistical Association, 24(166):225–227, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276705.

Parker:1930:NIR

[Par30] Harold G. Parker. Needs of investors with respect to corporation statistics. Journal of the American Statistical Association, 25(169): 13–17, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277190.

Paton:1922:BRB

[Pat22] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 18(138):286–289, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277545.

Paton:1923:BRBa

[Pat23a] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 18(142):804–808, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276725.

Paton:1923:BRBb

[Pat23b] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1071–1073, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965678.

Paton:1924:BRB

[Pat24] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 19(146):257–259, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277245.

Patterson:1925:BRB

[Pat25] Ernest Minor Patterson. Book review: [untitled]. Journal of the American Statistical Association, 20(151):446–448, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965536.

Paton:1926:BRB

[Pat26] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 21(153):102–105, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965511.

Paton:1927:BRB

[Pat27] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 22(157):116–118, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277358.

Patterson:1928:BRB

[Pat28] Ernest Minor Patterson. Book review: [untitled]. Journal of the American Statistical Association, 23(163):344–346, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276755.

Paton:1929:BRB

[Pat29] W. A. Paton. Book review: [untitled]. Journal of the American Statistical Association, 24(165):96–97, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277332.

Peabody:1924:GCR

[Pea24] Leroy E. Peabody. Growth curves and railway traffic. Journal of the American Statistical Association, 19(148):476–483, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281564.

Peabody:1925:SNR

[Pea25] Leroy E. Peabody. A study of net railway operating income. Journal of the American Statistical Association, 20(151):368–379, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2965520.

Pearl:1928:IPU

[Pea28] Raymond Pearl. International population union. Journal of the American Statistical Association, 23(162):190–192, June 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276897.

Person:1922:BRB

[Per22] H. S. Person. Book review: [untitled]. Journal of the American Statistical Association, 18(138):278–280, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277541.

Persons:1923:CTS

[Per23] Warren M. Persons. Correlation of time series. Journal of the American Statistical Association, 18(142):713–726, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276710.

Persons:1924:SFC

[Per24] Warren M. Persons. Some fundamental concepts of statistics. Journal of the American Statistical Association, 19(145):1–8, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277253.

Perry:1929:NIP

[Per29] F. Gardiner Perry. A new index of the physical volume of Canadian business. Journal of the American Statistical Association, 24(166):127–143, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276684.

Persons:1930:UCP

[Per30] Charles E. Persons. Unemployment as a census problem. Journal of the American Statistical Association, 25(169):117–120, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277212.

Peterson:1922:NSP

[Pet22] Ruth E. K. Peterson. New Swedish price index numbers. Journal of the American Statistical Association, 18(139):396–397, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276961.

Phelps:1922:BRB

[Phe22a] Roswell F. Phelps. Book review: [untitled]. Journal of the American Statistical Association, 18(140):540–543, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276936.

Phelps:1922:NWE

[Phe22b] Roswell F. Phelps. Number of wage-earners in manufacturing establishments in Massachusetts, 1914–21. Journal of the American Statistical Association, 18(137):104–106, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277475.

Pietra:1926:PCN

[Pie26] G. Pietra. A particular case of non-representative sampling. Journal of the American Statistical Association, 21(155):330–332, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277060.

Pogue:1922:BRB

[Pog22] Joseph E. Pogue. Book review: [untitled]. Journal of the American Statistical Association, 18(137):134–136, March 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277485.

Pogue:1923:BRB

[Pog23] Joseph E. Pogue. Book review: [untitled]. Journal of the American Statistical Association, 18(144):1075–1076, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965680.

Pogue:1926:BRB

[Pog26] Joseph E. Pogue. Book review: [untitled]. Journal of the American Statistical Association, 21(156):511–512, December 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276997.

Polakov:1922:KSA

[Pol22] Walter N. Polakov. Kinetic statistics as an aid to production and distribution. Journal of the American Statistical Association, 18 (139):359–365, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276954.

Powers:1928:FPU

[Pow28] Le Grand Powers. Factors producing unequal wealth distribution and ownership and measurements of their influence. Journal of the American Statistical Association, 23(164):417–428, December 1928.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277287.

Pidgin:1923:MN

[PP23] Charles Felton Pidgin and R. F. P. Miscellaneous notes. Journal of the American Statistical Association, 18(143):919–924, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965479.

Pearl:1929:PAA

[PR29] Raymond Pearl and Lowell J. Reed. The population of an area around Chicago and the logistic curve. Journal of the American Statistical Association, 24(165):66–67, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277319.

Prescott:1922:LGF

[Pre22] Raymond B. Prescott. Law of growth in forecasting demand. Journal of the American Statistical Association, 18(140):471–479, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276919.

R:1923:NHS

[R.23] W. S. R. The New Hampshire Survey. Journal of the American Statistical Association, 18(144):1043–1045, December 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965668.

Rauber:1924:BRB

[Rau24] H. W. Rauber. Book review: [untitled]. Journal of the American Statistical Association, 19(148):555–556, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281582.

Rossiter:1922:TRA

[RDG^{+}22] William S. Rossiter, Carroll W. Doten, Edwin F. Gay, Wesley C. Mitchell, Edwin R. A. Seligman, Walter F. Willcox, and Allyn A. Young. Third report of the advisory committee of the census. Journal of the American Statistical Association, 18(137):82–87, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277470.

Reed:1925:FSC

[Ree25] Lowell J. Reed. A form of saturation curve. Journal of the American Statistical Association, 20(151):390–396, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965522.

Reed:1926:BRB

[Ree26] Lowell J. Reed. Book review: [untitled]. Journal of the American Statistical Association, 21(155):367, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277074.

Rice:1929:IMS

[RG29] Stuart A. Rice and Morris Green. Interlocking memberships of social science societies. Journal of the American Statistical Association, 24(167):303–306, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276861.

Rice:1924:NCC

[Ric24a] Stuart A. Rice. A note on the census categories, “urban” and “rural”. Journal of the American Statistical Association, 19(145):79–81, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277263.

Rice:1924:PVF

[Ric24b] Stuart A. Rice. The political vote as a frequency distribution of opinion. Journal of the American Statistical Association, 19(145): 70–75, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277261.

Richter:1925:BRB

[Ric25a] F. E. Richter. Book review: [untitled]. Journal of the American Statistical Association, 20(150):287–288, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277133.

Richter:1925:RRE

[Ric25b] F. E. Richter. Richter’s review of “elements of business statistics”: a rejoinder. Journal of the American Statistical Association, 20(152):557–559, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277173.

Richter:1928:BRB

[Ric28] F. E. Richter. Book review: [untitled]. Journal of the American Statistical Association, 23(163):339–342, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276752.

Rider:1924:GLE

[Rid24] Paul R. Rider. A generalized law of error. Journal of the American Statistical Association, 19(146):217–220, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277233.

Rider:1925:BRB

[Rid25] Paul R. Rider. Book review: [untitled]. Journal of the American Statistical Association, 20(150):297–299, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277138.

Rider:1927:AUS

[Rid27] Paul R. Rider. Analysis of a U-shaped frequency distribution. Journal of the American Statistical Association, 22(158):202–208, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276773.

Rietz:1923:BRB

[Rie23] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 18(143):942–944, September 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965489.

Riefler:1925:BRB

[Rie25a] Winfield W. Riefler. Book review: [untitled]. Journal of the American Statistical Association, 20(149):156–157, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277458.

Riegel:1925:RRE

[Rie25b] Robert Riegel. Richter’s review of “elements of business statistics”. Journal of the American Statistical Association, 20(152):556–557, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277172.

Riefler:1926:BRB

[Rie26] Winfield W. Riefler. Book review: [untitled]. Journal of the American Statistical Association, 21(156):499–503, December 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276992.

Riefler:1928:BRB

[Rie28] Winfield W. Riefler. Book review: [untitled]. Journal of the American Statistical Association, 23(163):354–355, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276762.

Riefler:1929:BRB

[Rie29a] Winfield W. Riefler. Book review: [untitled]. Journal of the American Statistical Association, 24(167):335–337, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276871.

Rietz:1929:BRB

[Rie29b] H. L. Rietz. Book review: [untitled]. Journal of the American Statistical Association, 24(166):221, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2276702.

Robinson:1922:BRB

[Rob22] Louis N. Robinson. Book review: [untitled]. Journal of the American Statistical Association, 18(138):293–294, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277548.

Robbins:1926:BRB

[Rob26] Rainard B. Robbins. Book review: [untitled]. Journal of the American Statistical Association, 21(154):237–239, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277154.

Robb:1929:VDM

[Rob29a] Richard A. Robb. The variate difference method of seasonal variation. Journal of the American Statistical Association, 24(167):250–257, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276855.

Robbins:1929:BRB

[Rob29b] Rainard B. Robbins. Book review: [untitled]. Journal of the American Statistical Association, 24(165):105, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277339.

Robinson:1930:RIC

[Rob30] Leland Rex Robinson. The role of investment companies in the security markets of 1929. Journal of the American Statistical Association, 25(169):83–87, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277204.

Roelse:1923:ESU

[Roe23] Harold V. Roelse. The employment statistics of the United States employment service. Journal of the American Statistical Association, 18(142):791–793, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276720.

Roesle:1925:MGE

[Roe25] E. Roesle. The mortality in Germany, 1913–1921: The effects of war casualities and famine on mortality. Journal of the American Statistical Association, 20(150):163–178, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277112.

Roesle:1926:CCD

[Roe26] E. Roesle. Classification of causes of deaths and death registration: Principle and objects of the classification of the causes of death. Journal of the American Statistical Association, 21(154):195–205, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277145.

Rolph:1928:BRBa

[Rol28a] John G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 23(161):107–108, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277424.

Rolph:1928:BRBb

[Rol28b] John G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 23(162):227–228, June 1928. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276911.

Rolph:1929:BRB

[Rol29] John G. Rolph. Book review: [untitled]. Journal of the American Statistical Association, 24(168):449, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277101.

Rorty:1922:PMF

[Ror22] M. C. Rorty. Physical measures as functions of frequency distributions. Journal of the American Statistical Association, 18 (139):391–393, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276958.

Rorty:1925:ALA

[Ror25] M. C. Rorty. An attempt at a logical approach to certain fundamentals in physical mathematics. Journal of the American Statistical Association, 20(151):421–423, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965529.

Rorty:1930:MIV

[Ror30] Malcolm C. Rorty. Measuring investment values. Journal of the American Statistical Association, 25(169):100–105, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277208.

Ross:1922:BRB

[Ros22] Frank A. Ross. Book review: [untitled]. Journal of the American Statistical Association, 18(137):143–144, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277490.

Rossiter:1923:APG

[Ros23] William S. Rossiter. The adventure of population growth. Journal of the American Statistical Association, 18(141):561–574, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277497.

Ross:1924:BRB

[Ros24a] Frank A. Ross. Book review: [untitled]. Journal of the American Statistical Association, 19(147):421–423, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277391.

Rossiter:1924:C

[Ros24b] W. S. Rossiter. Communications. Journal of the American Statistical Association, 19(145):95–96, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2277268.

Ross:1925:FFC

[Ros25] Frank Alexander Ross. Formulae for facilitating computations in time series analyses. Journal of the American Statistical Association, 20(149):75–79, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277435.

Rossiter:1926:USC

[Ros26] W. S. Rossiter. United States cars in Canada. Journal of the American Statistical Association, 21(156):478–479, December 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276984.

Roswell:1927:BRB

[Ros27] David S. Roswell. Book review: [untitled]. Journal of the American Statistical Association, 22(158):258–261, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276794.

Rubinow:1925:BRB

[Rub25] I. M. Rubinow. Book review: [untitled]. Journal of the American Statistical Association, 20(150):275–280, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277128.

Rubinow:1926:NSC

[Rub26] I. M. Rubinow. The new series on Commonwealth of Pennsylvania departmental statistics. Journal of the American Statistical Association, 21(154):215–217, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277148.

Rubinow:1928:BRB

[Rub28] I. M. Rubinow. Book review: [untitled]. Journal of the American Statistical Association, 23(161):102–104, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277421.

Sasuly:1929:BRB

[Sas29] Max Sasuly. Book review: [untitled]. Journal of the American Statistical Association, 24(167):337, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276872.

Sasuly:1930:GMC

[Sas30] Max Sasuly. Generalized multiple correlation analysis of economic statistical series. Journal of the American Statistical Association, 25(169):146–152, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277218.

Schaffer:1925:FFR

[Sch25a] John G. Schaffer. A formula for freight rates. Journal of the American Statistical Association, 20(151):416–419, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965527.

Schultz:1925:EMM

[Sch25b] Henry Schultz. An extension of the method of moments. Journal of the American Statistical Association, 20(150):242–244, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277119.

Schultz:1928:BRB

[Sch28] Henry Schultz. Book review: [untitled]. Journal of the American Statistical Association, 23(162):215–216, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276906.

Secrist:1922:BRB

[Sec22a] Horace Secrist. Book review: [untitled]. Journal of the American Statistical Association, 18(137):140–142, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277488.

Secrist:1922:C

[Sec22b] Horace Secrist. A communication. Journal of the American Statistical Association, 18(140):525–528, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276930.

Secrist:1923:SP

[Sec23] Horace Secrist. A “statistical paradox”. Journal of the American Statistical Association, 18(142):776–780, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276716.

Secrist:1927:BRB

[Sec27] Horace Secrist. Book review: [untitled]. Journal of the American Statistical Association, 22(157):129, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277366.

Secrist:1928:BRB

[Sec28] Horace Secrist. Book review: [untitled]. Journal of the American Statistical Association, 23(164):463–464, December 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277303.

Shaw:1922:RMF

[Sha22] E. R. Shaw. The ratio of male to female births as affected by wars. Journal of the American Statistical Association, 18(138):250–255, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2277532.

Sharfman:1924:BRB

[Sha24] I. L. Sharfman. Book review: [untitled]. Journal of the American Statistical Association, 19(145):116–118, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277275.

Shewhart:1925:ASA

[She25] W. A. Shewhart. The application of statistics as an aid in maintaining quality of a manufactured product. Journal of the American Statistical Association, 20(152):546–548, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277170.

Shugrue:1926:BRB

[Shu26] Martin J. Shugrue. Book review: [untitled]. Journal of the American Statistical Association, 21(153):112, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2965515.

Shultz:1929:BRB

[Shu29] William J. Shultz. Book review: [untitled]. Journal of the American Statistical Association, 24(165):101–102, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277335.

Sibley:1926:SFD

[Sib26] Elbridge Sibley. The size of families: a discussion of sampling. Journal of the American Statistical Association, 21(155):333–334, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277061.

Sibley:1928:GDA

[Sib28] Elbridge Sibley. A graphic device for arithmetic or geometric interpolation. Journal of the American Statistical Association, 23 (164):441–442, December 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277292.

Silberling:1929:BRB

[Sil29] Norman J. Silberling. Book review: [untitled]. Journal of the American Statistical Association, 24(167):330–331, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276868.

Simpson:1929:BRB

[Sim29] Kemper Simpson. Book review: [untitled]. Journal of the American Statistical Association, 24(168):449–450, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277102.

Slichter:1923:BRB

[Sli23] Sumner H. Slichter. Book review: [untitled]. Journal of the American Statistical Association, 18(142):809–816, June 1923. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276727.

Smith:1925:EES

[Smi25a] Bradford B. Smith. The error in eliminating secular trend and seasonal variation before correlating time series. Journal of the American Statistical Association, 20(152):543–545, December 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277169.

Smith:1925:FAC

[Smi25b] Bradford B. Smith. Forecasting the acreage of cotton. Journal of the American Statistical Association, 20(149):31–47, March 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277432.

Smith:1926:CAF

[Smi26] Bradford B. Smith. Combining the advantages of first-difference and deviation-from-trend methods of correlating time series. Journal of the American Statistical Association, 21(153):55–59, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965498.

Smith:1927:FVV

[Smi27] Bradford Bixby Smith. Forecasting the volume and value of the cotton crop. Journal of the American Statistical Association, 22 (160):442–459, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276827.

Smith:1929:AAE

[Smi29] Bradford Bixby Smith. Another attempt to explain multiple correlation in simple terms. Journal of the American Statistical Association, 24(165):61–65, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277318.

Snider:1930:PP

[Sni30] Joseph L. Snider. Production and prices in 1929. Journal of the American Statistical Association, 25(169):56–57, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277197.

Snodgrass:1922:FRB

[Sno22] Katharine Snodgrass. The Federal Reserve Board’s price index for Great Britain. Journal of the American Statistical Association, 18(138):241–247, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277530.

Snyder:1923:NIV

[Sny23] Carl Snyder. A new index of the volume of trade. Journal of the American Statistical Association, 18(144):949–963, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965657.

Snyder:1924:NCI

[Sny24a] Carl Snyder. A new clearings index of business for fifty years. Journal of the American Statistical Association, 19(147):329–335, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277372.

Snyder:1924:NIB

[Sny24b] Carl Snyder. A new index of business activity. Journal of the American Statistical Association, 19(145):36–41, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277257.

Snyder:1924:NIG

[Sny24c] Carl Snyder. A new index of the general price level from 1875. Journal of the American Statistical Association, 19(146):189–195, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277230.

Snyder:1925:RIV

[Sny25] Carl Snyder. The revised index of the volume of trade. Journal of the American Statistical Association, 20(151):397–404, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965523.

Snyder:1926:NCI

[Sny26] Carl Snyder. A new composite index of wages in the United States. Journal of the American Statistical Association, 21(156):466–470, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276981.

Snyder:1928:IVT

[Sny28] Carl Snyder. The index of volume of trade: Second revision. Journal of the American Statistical Association, 23(162):154–163, June 1928.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276888.

Snyder:1929:PP

[Sny29] Carl Snyder. The problem of prosperity. Journal of the American Statistical Association, 24(165):1–14, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277313.

Snyder:1930:BLP

[Sny30] Carl Snyder. Brokers’ loans and the pyramiding of credit. Journal of the American Statistical Association, 25(169):88–92, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277205.

Soule:1922:BRB

[Sou22] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 18(140):554–555, December 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276941.

Soule:1923:BRB

[Sou23] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 18(142):822–827, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276731.

Soule:1924:WIP

[Sou24] George Soule. Wage indexes and price indexes. Journal of the American Statistical Association, 19(148):517–519, December 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281568.

Soule:1925:BRB

[Sou25] George Soule. Book review: [untitled]. Journal of the American Statistical Association, 20(149):141–143, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277451.

Squires:1923:BRB

[Squ23] B. M. Squires. Book review: [untitled]. Journal of the American Statistical Association, 18(141):682–684, March 1923. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277515.

Stecker:1922:WSN

[Ste22a] Margaret Loomis Stecker. Wage studies of the national industrial conference board: a reply. Journal of the American Statistical Association, 18(138):258–263, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277534.

Stephens:1922:BRB

[Ste22b] Margaret Gantt Stephens. Book review: [untitled]. Journal of the American Statistical Association, 18(139):419, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276973.

Stewart:1922:SRC

[Ste22c] C. L. Stewart. A summary of the results of the 1920 census of agriculture: Discussion. Journal of the American Statistical Association, 18(138):177–180, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277525.

Stewart:1922:BRB

[Ste22d] W. W. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 18(138):292–293, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277547.

Stecker:1924:BRB

[Ste24a] Margaret Loomis Stecker. Book review: [untitled]. Journal of the American Statistical Association, 19(145):126–129, March 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277280.

Stewart:1924:BRB

[Ste24b] Bryce M. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 19(148):540–543, December 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281576.

Stewart:1924:MN

[Ste24c] Ethelbert Stewart. Miscellaneous notes. Journal of the American Statistical Association, 19(146):236–240, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277239.

Stecker:1927:BRB

[Ste27a] Margaret Loomis Stecker. Book review: [untitled]. Journal of the American Statistical Association, 22(158):255–257, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276792.

Stewart:1927:BRB

[Ste27b] Bryce M. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 22(159):400–401, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276817.

Stewart:1927:RDB

[Ste27c] Ethelbert Stewart. Recent developments in the Bureau of Labor Statistics. Journal of the American Statistical Association, 22 (160):513–514, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276837.

Stocking:1928:BRB

[Sto28] George Ward Stocking. Book review: [untitled]. Journal of the American Statistical Association, 23(162):216–218, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276907.

Sturges:1926:CCI

[Stu26] Herbert A. Sturges. The choice of a class interval. Journal of the American Statistical Association, 21(153):65–66, March 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965501.

Shewhart:1928:SSN

[SW28] W. A. Shewhart and F. W. Winters. Small samples–new experimental results. Journal of the American Statistical Association, 23(162):144– 153, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276887.

Sydenstricker:1924:BRB

[Syd24] Edgar Sydenstricker. Book review: [untitled]. Journal of the American Statistical Association, 19(147):420–421, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277390.

Sydenstricker:1925:PSF

[Syd25] Edgar Sydenstricker. Population statistics of foreign countries. Journal of the American Statistical Association, 20(149):80–89, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277436.

Sydenstricker:1928:BRB

[Syd28a] Edgar Sydenstricker. Book review: [untitled]. Journal of the American Statistical Association, 23(164):462–463, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277302.

Sydenstricker:1928:SPP

[Syd28b] Edgar Sydenstricker. The Statistician’s place in public health work. Journal of the American Statistical Association, 23(162):115–120, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276884.

Taussig:1926:BRB

[Tau26] F. W. Taussig. Book review: [untitled]. Journal of the American Statistical Association, 21(156):497–499, December 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276991.

Taylor:1923:STD

[Tay23] J. S. Taylor. A statistical theory of depreciation. Journal of the American Statistical Association, 18(144):1010–1023, December 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965662.

Taylor:1929:BRB

[Tay29] Maurice Taylor. Book review: [untitled]. Journal of the American Statistical Association, 24(165):86–89, March 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277327.

Thompson:1929:UOQ

[TB29] Donald S. Thompson and Robert W. Bachelor. The use of overlapping quarterly indexes in erratic series. Journal of the American Statistical Association, 24(167):301–302, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276860.

Tolley:1923:MHM

[TE23] H. R. Tolley and M. J. B. Ezekiel. A method of handling multiple correlation problems. Journal of the American Statistical Association, 18(144):993–1003, December 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2965660.

Tolley:1927:DMS

[TE27] H. R. Tolley and Mordecai Ezekiel. The Doolittle method for solving multiple correlation equations versus the Kelley–Salisbury “iteration” method. Journal of the American Statistical Association, 22(160):497–500, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276832.

Thompson:1923:PFU

[Tho23] Warren S. Thompson. Population facts for the United States and their interpretation. Journal of the American Statistical Association, 18(141):575–587, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277498.

Thomas:1925:BRBa

[Tho25a] Dorothy Swaine Thomas. Book review: [untitled]. Journal of the American Statistical Association, 20(151):441–443, September 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965534.

Thomas:1925:BRBb

[Tho25b] Woodlief Thomas. Book review: [untitled]. Journal of the American Statistical Association, 20(151):443–446, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965535.

Thomas:1925:BPI

[Tho25c] Woodlief Thomas. British production indexes. Journal of the American Statistical Association, 20(151):413–416, September 1925. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965526.

Thompson:1925:SFW

[Tho25d] Warren S. Thompson. Size of families from which college students come. Journal of the American Statistical Association, 20 (152):481–495, December 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277163.

Thorp:1925:BRB

[Tho25e] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 20(149):135–138, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277448.

Thomas:1926:BRB

[Tho26a] Dorothy Swaine Thomas. Book review: [untitled]. Journal of the American Statistical Association, 21(155):363–365, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277072.

Thomas:1926:IBB

[Tho26b] Dorothy Swaine Thomas. An index of British business cycles. Journal of the American Statistical Association, 21(153):60–63, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965499.

Thorp:1926:BRB

[Tho26c] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 21(153):96–99, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965508.

Thomas:1927:BRB

[Tho27a] Dorothy Swaine Thomas. Book review: [untitled]. Journal of the American Statistical Association, 22(158):262–263, June 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276796.

Thomas:1927:CIN

[Tho27b] Woodlief Thomas. Construction of an index number of production. Journal of the American Statistical Association, 22(159):315–330, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276801.

Thompson:1927:BRB

[Tho27c] Warren S. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 22(157):119–120, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277359.

Thorp:1927:CRB

[Tho27d] Willard L. Thorp. Comment on the review of “Business Annals”. Journal of the American Statistical Association, 22(158):228, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276779.

Thorp:1928:BRB

[Tho28] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 23(161):94–97, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277417.

Thorp:1929:BRB

[Tho29a] Willard L. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 24(167):339–340, September 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276874.

Thorp:1929:HSM

[Tho29b] Willard L. Thorp. Horsepower statistics for manufactures. Journal of the American Statistical Association, 24(168):376–385, December 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277083.

Tucker:1930:AEF

[TMT30] Donald Tucker, Arthur W. Marget, and John G. Thompson. Appraisal of economic forecasts: Discussion. Journal of the American Statistical Association, 25(169):41–50, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277195.

Tobey:1922:DVS

[Tob22] James A. Tobey. The division of vital statistics of the United States Bureau of the Census. Journal of the American Statistical Association, 18(138):268–271, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277537.

Truesdell:1922:SRC

[Tru22] Leon E. Truesdell. A summary of the results of the 1920 census of agriculture. Journal of the American Statistical Association, 18(138):167– 177, June 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277524.

Truesdell:1930:CP

[Tru30] Leon E. Truesdell. The census of population. Journal of the American Statistical Association, 25(169):113–116, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277211.

Tryon:1922:BRBa

[Try22a] F. G. Tryon. Book review: [untitled]. Journal of the American Statistical Association, 18(137):144–146, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277491.

Tryon:1922:BRBb

[Try22b] F. G. Tryon. Book review: [untitled]. Journal of the American Statistical Association, 18(139):418–419, September 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276972.

Tryon:1927:ICF

[Try27] F. G. Tryon. An index of consumption of fuels and water power. Journal of the American Statistical Association, 22(159):271–282, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276799.

Tryon:1929:BRB

[Try29] F. G. Tryon. Book review: [untitled]. Journal of the American Statistical Association, 24(166):229–230, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276707.

Tucker:1922:DMP

[Tuc22] Rufus S. Tucker. Distribution of men physically unfit for military service. Journal of the American Statistical Association, 18 (139):377–384, September 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276956.

Tucker:1927:BRB

[Tuc27] Rufus S. Tucker. Book review: [untitled]. Journal of the American Statistical Association, 22(158):247–249, June 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276789.

Thorndike:1926:REP

[TW26] E. L. Thorndike and Ella Woodyard. The relation of earning power to age in professional workers under conditions of nearly free competition. Journal of the American Statistical Association, 21 (155):293–309, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277057.

Thorndike:1927:EVP

[TW27] Edward L. Thorndike and Ella Woodyard. The effect of violent pricefluctuations upon the salaries of clergymen. Journal of the American Statistical Association, 22(157):66–74, March 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277350.

Usher:1929:BRB

[Ush29] Abbott Payson Usher. Book review: [untitled]. Journal of the American Statistical Association, 24(165):102, March 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277336.

Vanderblue:1922:BRB

[Van22] Homer B. Vanderblue. Book review: [untitled]. Journal of the American Statistical Association, 18(137):142–143, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277489.

VanKleeck:1924:ARC

[Van24] Mary Van Kleeck. Annual report of the committee of the American statistical association on governmental labor statistics. Journal of the American Statistical Association, 19(145):76–79, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277262.

VanBuren:1926:PFD

[Van26] George H. Van Buren. The past and future development of vital statistics in the United States: II, William Alexander King and the Federal Registration Service, 1900–1906. Journal of the American Statistical Association, 21(155):267–272, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277052.

VanSickle:1928:BRB

[Van28] John V. Van Sickle. Book review: [untitled]. Journal of the American Statistical Association, 23(164):456–457, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277299.

Varga:1928:EAT

[Var28] Stephen Varga. An expression for the asymmetrical tendency of frequency distributions. Journal of the American Statistical Association, 23(161):35–39, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277401.

vonHuhn:1925:BRBa

[vH25a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 20(149):153–156, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277457.

vonHuhn:1925:BRBb

[vH25b] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 20(150):290–293, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277135.

vonHuhn:1926:BRB

[vH26] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 21(154):245–247, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277158.

vonHuhn:1927:BRBb

[vH27a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 22(159):409–411, September 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276823.

vonHuhn:1927:FSG

[vH27b] R. von Huhn. Further studies in the graphic use of circles and bars: a discussion of the Eell’s experiment. Journal of the American Statistical Association, 22(157):31–39, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277346.

vonHuhn:1928:BRB

[vH28a] R. von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 23(163):354, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276761.

vonHuhn:1928:CBR

[vH28b] R. von Huhn. A correction: Book reviews. Journal of the American Statistical Association, 23(161):85, March 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277414.

Viner:1922:BRB

[Vin22] Jacob Viner. Book review: [untitled]. Journal of the American Statistical Association, 18(137):127–132, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277483.

Viner:1924:BRBa

[Vin24a] Jacob Viner. Book review: [untitled]. Journal of the American Statistical Association, 19(146):262–264, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277247.

Viner:1924:BRBb

[Vin24b] Jacob Viner. Book review: [untitled]. Journal of the American Statistical Association, 19(148):544–546, December 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2281577.

Viner:1927:BRB

[Vin27] Jacob Viner. Book review: [untitled]. Journal of the American Statistical Association, 22(157):125–126, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277363.

VonHuhn:1927:BRBa

[Von27] R. Von Huhn. Book review: [untitled]. Journal of the American Statistical Association, 22(157):128–129, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277365.

W:1928:WRP

[W.28] L. W. Wage rates and per capita productivity. Journal of the American Statistical Association, 23(162):180–181, June 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276892.

Waite:1930:SDT

[Wai30] Warren C. Waite. Some developments in the techniques of studying consumer demand. Journal of the American Statistical Association, 25(169):140–145, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277217.

Warren:1926:BRB

[War26] Robert B. Warren. Book review: [untitled]. Journal of the American Statistical Association, 21(153):94–96, March 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2965507.

Warren:1927:BRBa

[War27a] R. B. Warren. Book review: [untitled]. Journal of the American Statistical Association, 22(158):257–258, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276793.

Warren:1927:BRBb

[War27b] Robert Warren. Book review: [untitled]. Journal of the American Statistical Association, 22(160):527–528, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276845.

Warren:1930:ETY

[War30] Robert B. Warren. Europe ten years after the War: Discussion. Journal of the American Statistical Association, 25(169):174–176, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277224.

Watkins:1922:BRB

[Wat22] G. P. Watkins. Book review: [untitled]. Journal of the American Statistical Association, 18(139):409–410, September 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276966.

Watkins:1923:BRB

[Wat23a] G. P. Watkins. Book review: [untitled]. Journal of the American Statistical Association, 18(141):694, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2277520.

Watkins:1923:MCVb

[Wat23b] G. P. Watkins. The measurement of concomitance of variation. Journal of the American Statistical Association, 18(142):741–758, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276712.

Watkins:1923:MCVa

[Wat23c] G. P. Watkins. The measurement of concomitance of variation: Proposed indexes of variability and correlation based upon logarithmic curves and having reference primarily to time series. Journal of the American Statistical Association, 18(141):588–599, March 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277499.

Watkins:1927:BRB

[Wat27a] G. P. Watkins. Book review: [untitled]. Journal of the American Statistical Association, 22(159):404–405, September 1927. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276819.

Watkins:1927:MFE

[Wat27b] Ralph J. Watkins. Measuring the fluctuations of employment in the Ohio construction industry. Journal of the American Statistical Association, 22(157):40–48, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277347.

Watkins:1930:PUI

[Wat30] George P. Watkins. Problems in the utilization and interpretation of corporation statistics. Journal of the American Statistical Association, 25(169):18–22, March 1930. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277191.

Waugh:1929:UIL

[Wau29] Frederick V. Waugh. The use of isorropic lines in determining regression surfaces. Journal of the American Statistical Association, 24(166):144–151, June 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276685.

Wembridge:1923:ESS

[Wem23] Harry A. Wembridge. Experiment and statistics in the selection of employees: Methods used in the selection of machine operators in a clothing plant. Journal of the American Statistical Association, 18(141):600–606, March 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277500.

Westergaard:1922:MN

[Wes22] Harald Westergaard. Miscellaneous notes. Journal of the American Statistical Association, 18(139):398–402, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276962.

Weyforth:1922:BRB

[Wey22] William O. Weyforth. Book review: [untitled]. Journal of the American Statistical Association, 18(139):415–417, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276970.

Weyforth:1926:BRB

[Wey26] William O. Weyforth. Book review: [untitled]. Journal of the American Statistical Association, 21(153):105–107, March 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965512.

Weyforth:1927:BRB

[Wey27] William O. Weyforth. Book review: [untitled]. Journal of the American Statistical Association, 22(160):538–542, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276850.

Woodbury:1927:IMB

[WH27] Robert M. Woodbury and F. H. Hankins. Infant mortality and the business cycle. Journal of the American Statistical Association, 22(157):79–81, March 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277352.

Whelden:1926:TSN

[Whe26a] C. H. Whelden, Jr. The trend-seasonal normal in time series. Journal of the American Statistical Association, 21(155):321–329, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277059.

Whelpton:1926:OGU

[Whe26b] P. K. Whelpton. Occupational groups in the United States, 1820–1920. Journal of the American Statistical Association, 21 (155):335–343, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277062.

Whelden:1927:R

[Whe27] C. H. Whelden, Jr. A retraction. Journal of the American Statistical Association, 22(158):229, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276780.

Whelpton:1929:DTN

[Whe29] P. K. Whelpton. Differentials in true natural increase. Journal of the American Statistical Association, 24(167):233–249, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276854.

White:1922:IFS

[Whi22] Edward White. Income fluctuation of a selected group of personal returns. Journal of the American Statistical Association, 18(137):67– 81, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2277469.

Whipple:1923:BRB

[Whi23] George C. Whipple. Book review: [untitled]. Journal of the American Statistical Association, 18(143):926–928, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965481.

Willcox:1922:BRBa

[Wil22a] Walter F. Willcox. Book review: [untitled]. Journal of the American Statistical Association, 18(138):276–278, June 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277540.

Willcox:1922:BRBb

[Wil22b] Walter F. Willcox. Book review: [untitled]. Journal of the American Statistical Association, 18(140):534–538, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276934.

Willis:1922:BRB

[Wil22c] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 18(139):414–415, September 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276969.

Wilson:1922:BRB

[Wil22d] Edwin Bidwell Wilson. Book review: [untitled]. Journal of the American Statistical Association, 18(140):556–557, December 1922. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276943.

Willcox:1923:PWW

[Wil23a] Walter F. Willcox. Population and the World War: a preliminary survey. Journal of the American Statistical Association, 18(142):699– 712, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2276709.

Willis:1923:BRB

[Wil23b] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 18(143):938–939, September 1923. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965486.

Wilson:1923:FSL

[Wil23c] Edwin Bidwell Wilson. First and second laws of error. Journal of the American Statistical Association, 18(143):841–851, September 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2965467.

Willcox:1924:BRB

[Wil24a] Walter F. Willcox. Book review: [untitled]. Journal of the American Statistical Association, 19(146):246–249, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277242.

Willcox:1924:RUS

[Wil24b] Walter F. Willcox. The relation of the United States to international statistics. Journal of the American Statistical Association, 19(147):348–361, September 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277374.

Willcox:1924:UDR

[Wil24c] Walter F. Willcox. U.S. divorce rates, 1922. Journal of the American Statistical Association, 19(147):387–389, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277378.

Willis:1924:BRB

[Wil24d] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 19(147):417–419, September 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277388.

Willis:1924:TYE

[Wil24e] H. Parker Willis. Ten years’ experience in business statistics. Journal of the American Statistical Association, 19(146):206–217, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277232.

Wilson:1925:MTB

[Wil25] McDonald H. Wilson. The major trends of bond prices. Journal of the American Statistical Association, 20(149):65–74, March 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2277434.

Wilcox:1926:ISI

[Wil26a] Walter F. Wilcox. International statistical institute. Journal of the American Statistical Association, 21(153):64–65, March 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2965500.

Willcox:1926:ADS

[Wil26b] Walter F. Willcox. Arrests and deaths since 1920 attributable to alcohol. Journal of the American Statistical Association, 21 (156):458–465, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276980.

Willcox:1926:PFD

[Wil26c] Walter F. Willcox. The past and future development of vital statistics in the United States: I, John Shaw Billings and Federal Vital Statistics. Journal of the American Statistical Association, 21 (155):257–266, September 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277051.

Willis:1926:BRB

[Wil26d] H. Parker Willis. Book review: [untitled]. Journal of the American Statistical Association, 21(154):243–245, June 1926. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277157.

Will:1927:MSD

[Wil27a] Harry Sylvester Will. A measure of stability with a discussion of certain of its implications. Journal of the American Statistical Association, 22(158):184–194, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276771.

Williams:1927:BRBa

[Wil27b] Faith M. Williams. Book review: [untitled]. Journal of the American Statistical Association, 22(159):394–397, September 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276815.

Williams:1927:BRBb

[Wil27c] Faith M. Williams. Book review: [untitled]. Journal of the American Statistical Association, 22(160):532–536, December 1927. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276848.

Wilson:1927:BRB

[Wil27d] E. B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 22(160):532, December 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2276847.

Wilson:1927:PIL

[Wil27e] Edwin B. Wilson. Probable inference, the law of succession, and statistical inference. Journal of the American Statistical Association, 22(158):209–212, June 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276774.

Willcox:1928:MLW

[Wil28a] Walter F. Willcox. Military losses in the world war. Journal of the American Statistical Association, 23(163):304–305, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2276743.

Willcox:1928:PC

[Wil28b] Walter F. Willcox. The population of China in 1910. Journal of the American Statistical Association, 23(161):18–30, March 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277399.

Williamson:1928:BRB

[Wil28c] K. M. Williamson. Book review: [untitled]. Journal of the American Statistical Association, 23(163):347–349, September 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2276757.

Williams:1929:MDF

[Wil29] Faith M. Williams. The measurement of the demand for food. Journal of the American Statistical Association, 24(167):288–295, September 1929. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2276858.

Williams:1930:SFM

[Wil30] Faith M. Williams. Scales for family measurement. Journal of the American Statistical Association, 25(169):135–139, March 1930. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277216.

Winslow:1922:MWE

[Win22] Mary N. Winslow. Medians of Women’s earnings in four states. Journal of the American Statistical Association, 18(137):106–108, March 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277476.

Wing:1925:BRB

[Win25] David L. Wing. Book review: [untitled]. Journal of the American Statistical Association, 20(149):147–149, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277453.

W:1924:CSD

[WKDR24] W. F. W., B. W. Kilgore, L. I. Dublin, and W. S. Rossiter. Cotton statistics from the Department of Commerce. Journal of the American Statistical Association, 19(146):226–230, June 1924. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277235.

Wolman:1930:ALS

[Wol30] Leo Wolman. American labor since 1920. Journal of the American Statistical Association, 25(169):158–163, March 1930. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277221.

Woodbury:1922:WME

[Woo22] Robert M. Woodbury. Westergaard’s method of expected deaths as applied to the study of infant mortality. Journal of the American Statistical Association, 18(139):366–376, September 1922. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276955.

Woodbury:1924:BRB

[Woo24a] Robert M. Woodbury. Book review: [untitled]. Journal of the American Statistical Association, 19(145):119–120, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277276.

Woodbury:1924:EFI

[Woo24b] Robert Morse Woodbury. Economic factors in infant mortality. Journal of the American Statistical Association, 19(146):137–155, June 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2277226.

Woodbury:1926:BRB

[Woo26] Helen Sumner Woodbury. Book review: [untitled]. Journal of the American Statistical Association, 21(154):247–249, June 1926. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277159.

Wood:1928:SRC

[Woo28] Edith Elmer Wood. The statistics of Room congestion: Purpose and technique. Journal of the American Statistical Association, 23 (163):263–273, September 1928. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276737.

Working:1922:DST

[Wor22] Holbrook Working. The determination of secular trend reconsidered. Journal of the American Statistical Association, 18(140):497–502, December 1922. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276922.

Working:1924:MPI

[Wor24] Holbrook Working. Monthly pig iron production, 1884–1902: An analysis of a problem in correlation. Journal of the American Statistical Association, 19(147):381–386, September 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277377.

Working:1928:BRB

[Wor28] Holbrook Working. Book review: [untitled]. Journal of the American Statistical Association, 23(161):89–94, March 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277416.

Working:1929:BRB

[Wor29] Holbrook Working. Book review: [untitled]. Journal of the American Statistical Association, 24(168):437–441, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277094.

Willey:1924:SCP

[WR24] Malcolm M. Willey and Stuart A. Rice. A sex cleavage in the presidential election of 1920. Journal of the American Statistical Association, 19(148):519–520, December 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281569.

Wright:1924:BRB

[Wri24] Philip Green Wright. Book review: [untitled]. Journal of the American Statistical Association, 19(145):103–108, March 1924. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277271.

Wright:1926:BRB

[Wri26a] Sewall Wright. Book review: The Biology of Population Growth, by Raymond Pearl; The Natural Increase of Mankind, by J. Shirley Sweeney. Journal of the American Statistical Association, 21 (156):493–497, December 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2276990.

Wright:1926:FCA

[Wri26b] Sewall Wright. A frequency curve adapted to variation in percentage occurrence. Journal of the American Statistical Association, 21(154):162–178, June 1926. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277143.

Wright:1928:BRB

[Wri28] Sewall Wright. Book review: [untitled]. Journal of the American Statistical Association, 23(163):336–339, September 1928. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276751.

Wright:1929:BRB

[Wri29] Philip G. Wright. Book review: [untitled]. Journal of the American Statistical Association, 24(166):207–215, June 1929. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2276697.

Westergaard:1923:MN

[WV23] Harald Westergaard and George H. Van Buren. Miscellaneous notes. Journal of the American Statistical Association, 18(142):795–797, June 1923. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2276722.

Yntema:1928:BRB

[Ynt28] Theodore O. Yntema. Book review: [untitled]. Journal of the American Statistical Association, 23(164):453–455, December 1928. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277297.

Young:1924:BRB

[You24] Allyn A. Young. Book review: [untitled]. Journal of the American Statistical Association, 19(146):242–246, June 1924. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277241.

Young:1925:BRB

[You25a] Allyn A. Young. Book review: [untitled]. Journal of the American Statistical Association, 20(149):133–135, March 1925. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/2277447.

Young:1925:PYR

[You25b] Allyn A. Young. Professor Young’s review of “Germany’s Capacity to Pay”: a rejoinder. Journal of the American Statistical Association, 20(150):260–261, June 1925. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2277124.

Zelomek:1929:NTD

[Zel29] A. W. Zelomek. The new trend in distribution. Journal of the American Statistical Association, 24(168):425–427, December 1929. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2277089.